### v2

# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

Abstract
The deployment of extremely large antenna arrays (ELAAs) and operation at higher frequency bands in wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) systems introduce significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties. Combined with dual-wideband impacts, these effects fundamentally reshape the sparsity patterns of wideband XL-MIMO channels in the angular-delay domain, making existing sparsity-based channel estimation methods inadequate. To address these challenges, this paper revisits the channel estimation problem for wideband XL-MIMO systems, considering dual-wideband effects, spherical wavefront, and SnS properties. By leveraging the spatial-chirp property of near-field array responses, we quantitatively characterize the sparsity patterns of wideband XL-MIMO channels in the angular-delay domain, revealing global block sparsity and local common-delay sparsity. Building on this structured sparsity, we formulate the wideband XL-MIMO channel estimation problem as a multiple measurement vector (MMV)-based Bayesian inference task and propose a novel column-wise hierarchical prior model to effectively capture the sparsity characteristics. To enable efficient channel reconstruction, we develop an MMV-based variational message passing (MMV-VMP) algorithm, tailored to the complex factor graph induced by the hierarchical prior. Simulation results validate the proposed algorithm, demonstrating its convergence and superior performance compared to existing methods, thus establishing its effectiveness in addressing the challenges of wideband XL-MIMO channel estimation under complex near-field conditions.
Index Terms:
Wideband XL-MIMO systems, dual-wideband effects, near-field effects, channel estimation, Bayesian inference.
I Introduction
Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2]. By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[3]. It also delivers substantial beamforming gains, effectively mitigating the severe path loss characteristic of millimeter-wave (mmWave) frequencies. Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [4]. However, realizing these performance benefits critically depends on obtaining accurate channel state information (CSI), underscoring the importance of precise channel knowledge.
Typically, wideband XL-MIMO systems adopt orthogonal frequency division multiplexing (OFDM) to combat the frequency-selective fading of wireless channels [5]. However, due to the pronounced impact of dual-wideband effects, it is challenging to acquire accurate CSI. Specifically, the large system bandwidth shortens the OFDM symbol duration, potentially causing the propagation delay difference across the antenna array to exceed the symbol duration, leading to the spatial wideband effect [6]. Concurrently, the varying central frequencies of subcarriers introduce the frequency-wideband effect, where phase shifts differ across subcarrier channels, further complicating the channel estimation process.
Moreover, with the deployment of ELAAs and operation at higher frequency bands, near-field effects, such as the spherical wavefront effect [7, 8, 9, 10] and spatially non-stationary (SnS) properties [11, 12], become prominent, exacerbating the dual-wideband effects. For instance, the sparsity structure of SnS dual-wideband XL-MIMO channels in the angular-delay domain undergoes significant changes. Specifically, the curvature of the spherical wavefront renders a single spatial frequency insufficient to characterize a near-field propagation path, requiring multiple spatial frequencies instead. This spatial frequency spread, combined with the frequency-wideband effect, amplifies angular spread compared to conventional massive MIMO systems [6, 13]. Additionally, the spherical wavefront effect disrupts the linear path delay variations across adjacent antennas, while SnS properties influence both spatial frequency and delay spreads. These combined factors reshape the sparse patterns of SnS dual-wideband XL-MIMO channels in the angular-delay domain, significantly complicating channel estimation in wideband XL-MIMO systems.
I-A Related Works
Due to the limited scattering resulting from the highly directional propagation behavior of mmWave, mmWave channels typically exhibit significant sparsity. Consequently, most existing channel estimation methods are designed to exploit this inherent sparsity structure to simplify and improve the estimation process. For instance, leveraging the sparsity in the polar domain, various channel estimation and beam training schemes have been proposed to address the spherical effects in XL-MIMO channels [14, 15, 16]. Considering the inherent spatial-domain sparsity induced by SnS properties, several Bayesian inference-based methods have been developed [17, 18] to mitigate the SnS effects. Additionally, exploiting the joint sparsity of XL-MIMO channels in both the spatial and angular domains, methods such as [8, 19] have proposed two-stage and joint VR detection and channel estimation algorithms, considering both spherical wavefront and SnS properties. However, despite these advancements, these methods predominantly focus on narrowband systems without the considerations of dual-wideband effects.
While some studies have explored wideband channels, such works remain limited in scope. For instance, [20] leverages the common polar-domain sparsity across subcarrier channels, employing the simultaneous orthogonal matching pursuit (SOMP) method for channel estimation. Nevertheless, as the bandwidth and array aperture increase in wideband XL-MIMO systems, the dual-wideband effects become more pronounced, challenging the assumption of common sparsity across subcarriers and diminishing the applicability of such methods.
Considering the sparsity differences among subcarriers, [13, 21] proposed methods that exploit subcarrier-dependent variations in sparsity patterns. Specifically, [13] recognized the sparsity-dependent pattern across subcarriers in dual-wideband channels and proposed jointly utilizing information from multiple subcarriers to enhance dual-wideband channel estimation performance. [21] developed a deep learning (DL)-based approach by unfolding the sparse Bayesian learning (SBL) algorithm into a deep neural network (DNN), where each SBL layer is carefully designed to capture subcarrier-dependent sparsity through a tailored variance parameter update mechanism. In addition, [22, 23, 24] developed the super-resolution-based estimation techniques to enhance performance by exploiting angular-delay domain sparsity. Unfortunately, these methods are tailored for far-field channels and cannot be readily extended to near-field scenarios due to the differences of sparsity patterns.
More recently, methods addressing both the spherical wavefront effect and dual-wideband effects have been proposed. For example, [25] introduced a bilinear pattern detection (BPD)-based approach to recover wideband XL-MIMO channels, while [26] developed a hybrid message passing algorithm based on constrained Bethe free energy minimization. However, despite these advancements, none of these approaches explicitly incorporate SnS properties, highlighting a significant gap in wideband XL-MIMO channel estimation.
I-B Motivations and Contributions
Due to the randomness of the environment and user locations, spherical wavefront effects and SnS properties are inevitable in wideband XL-MIMO systems. However, the channel estimation problem that jointly considers the spherical wavefront effect, SnS properties, and dual-wideband effects has not been well addressed to date. Specifically, a well-established channel model for wideband XL-MIMO systems is lacking, and the sparsity patterns of these channels in the angular-delay domain have yet to be fully explored. Additionally, with the increasing number of antennas and subcarriers in XL-MIMO systems, there is a pressing need for computationally efficient channel estimation algorithms to handle the growing complexity of these systems. To address these issues, this paper investigates channel estimation for XL-MIMO systems by incorporating the spherical wavefront effect, SnS properties, and dual-wideband effects. The main contributions of this paper are summarized as follows:
  • To accurately characterize the SnS dual-wideband XL-MIMO channels, we derive a novel spatial-frequency domain model based on the measurement findings reported in [11], which comprehensively accounts for the spherical wavefront effect, SnS characteristics, and dual-wideband effects. By leveraging the spatial-chirp property of the array response, we quantitatively investigate the impact of the spherical wavefront effect, SnS properties, and dual-wideband effects on the angular and delay spreads. Furthermore, we reveal the sparsity patterns of SnS dual-wideband XL-MIMO channels in the angular-delay domain, incorporating both global block sparsity and local common-delay sparsity.
  • Exploiting the sparsity of the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem. Recognizing the critical role of an accurate prior model in Bayesian inference, we propose a novel column-wise hierarchical prior model, where Markov random field (MRF) and common-delay sparsity mechanism is introduced to capture both the global block sparsity and the local common-delay sparsity, ensuring a more accurate and structured representation of the channel characteristics.
  • To enable efficient Bayesian inference, we propose a MMV-based variational message passing (MMV-VMP) algorithm tailored for the column-wise hierarchical prior model. Unlike the variance states propagation algorithm in [27], which relies on first-moment matching methods under the mean-field approximation to update prior model parameters, the MMV-VMP algorithm employs structured variational inference combined with belief propagation. This approach effectively handles the higher-order dependencies and loops in the complex factor graph, ensuring precise and reliable message updates.
  • Compared to existing methods, the proposed MMV-VMP algorithm achieves superior performance by jointly exploiting the global and common-delay sparsity of SnS dual-wideband channels in the angular-delay domain. Additionally, as the algorithm avoids matrix inversion operations, it maintains moderate computational complexity, making it well-suited for practical deployment in XL-MIMO systems.
I-C Organization and Notations
The rest of the paper is organized as follows: Section II introduces the mmWave XL-MIMO system model. Section III develops a spatial-frequency channel model, incorporating spherical wavefront effects, SnS properties, and dual-wideband effects, and analyzes the angular-delay characteristics of SnS dual-wideband XL-MIMO channels. Section IV formulates the channel estimation problem as a sparse signal recovery task and proposes a column-wise hierarchical sparse prior model. Section V presents the MMV-VMP algorithm. Simulation results and conclusions are provided in Sections VI and VII, respectively.
Notations: lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\odot$ denotes the Hadamard product; $\mathbf{I}_{n}$ denotes a $n\times n$ identity matrix; $\mathbb{C}^{m\times n}$ denotes a $m\times n$ complex matrix. $\mathbf{1}_{n}$ denotes a $n\times 1$ all-one vector. In addition, a random variable $x\in\mathbb{C}$ drawn from the complex Gaussian distribution with mean $x_{0}$ and variance $v$ is characterized by the probability density function $\mathcal{CN}(x;m,v)=\exp\{-{\left|x-m\right|^{2}}/{v}\}/{\pi v}$; a random variable $\gamma\in\mathbb{R}$ from Gamma distribution with mean $a/b$ and variance $a/b^{2}$ is characterized by the probability density function $\mathcal{G}a(\gamma;a,b)\propto\gamma^{a-1}\exp(-\gamma b)$.
II System Model
Figure 1: Illustration of the XL-MIMO system model.
Consider a mmWave XL-MIMO system, as shown in Fig. 1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration to serve a single-antenna user. To reduce hardware cost and energy consumption, the BS adopts a hybrid analog-digital architecture, where $N_{\mathrm{RF}}$ radio frequency (RF) chains are connected to the ELAA through phase shifters. The spacing between adjacent antennas is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$, with $c$ and $f_{c}$ indicating the speed of light and the central carrier frequency, respectively. In addition, to address frequency-selective fading, OFDM is utilized. Without loss of generality, assume that $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
In this paper, we focus on the uplink channel estimation, where the BS combines the pilot signal using all RF chains associated with different beams. Let the total number of received beams be denoted by $M$, with $M$ being an integer multiple of $N_{\mathrm{RF}}$. Consequently, the BS requires $P=M/N_{\mathrm{RF}}$ channel uses to cycle through all beams for a given pilot symbol. Denote $\iota_{k,p}$ as the pilot symbol of the $k$-th subcarrier in the $p$-th channel use. Then, the received signal for the $k$-th subcarrier in the $p$-th channel use can be expressed as
$\mathbf{y}_{k,p}=\mathbf{W}_{p}\mathbf{h}_{k}\iota_{k,p}+\mathbf{n}_{k,p},$ (1)
where $\mathbf{W}_{p}\in\mathbb{C}^{N_{\mathrm{RF}}\times N}$ denotes the combining matrix in the $p$-th channel use, with each entry adhering to a constant modulus constraint and independently generated from the set $\{-1/\sqrt{N},1/\sqrt{N}\}$; $\mathbf{h}_{k}\in\mathbb{C}^{\mathrm{N}\times 1}$ denotes the spatial-domain channel vector for the $k$-th subcarrier, while $\mathbf{n}_{k,p}\sim\mathcal{CN}\left(\mathbf{n}_{k,p};\mathbf{0},\beta^{-1}\mathbf{I}_{N_{\mathrm{RF}}}\right)$ represents the noise vector associated with the $k$-th subcarrier in the $p$-th channel use.
Assuming that all-one pilot symbols are adopted, i.e., $\iota_{k,p}=1$ for all $k$ and $p$, and collecting the pilot symbols across all $P$ channel uses, the received signal for the $k$-th subcarrier can be written in a compact form as
$\mathbf{y}_{k}=\mathbf{W}\mathbf{h}_{k}+{\mathbf{n}}_{k},$ (2)
where $\mathbf{W}=[\mathbf{W}_{1}^{\mathrm{T}},\mathbf{W}_{2}^{\mathrm{T}},\cdots,\mathbf{W}_{P}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{C}^{M\times N}$ and ${\mathbf{n}}_{k}=[\mathbf{n}_{k,1}^{\mathrm{T}},\mathbf{n}_{k,2}^{\mathrm{T}},\cdots,\mathbf{n}_{k,P}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{C}^{M\times 1}$ are the collective receive combining matrix and the effective noise vector, respectively. Furthermore, concatenating the received signals corresponding to different pilot subcarriers, the overall received signal $\mathbf{Y}=\left[\mathbf{y}_{1},\mathbf{y}_{2},\cdots,\mathbf{y}_{K}\right]\in\mathbb{C}^{M\times K}$ is given by
$\mathbf{Y}=\mathbf{W}\mathbf{H}+{\mathbf{N}},$ (3)
where $\mathbf{H}=\left[\mathbf{h}_{1},\cdots,\mathbf{h}_{K}\right]\in\mathbb{C}^{N\times K}$ and ${\mathbf{N}}=\left[{\mathbf{n}}_{1},\cdots,{\mathbf{n}}_{K}\right]\in\mathbb{C}^{M\times K}$ indicating the spatial-frequency channel matrix and effective noise matrix, respectively.
III SnS Dual-Wideband XL-MIMO Channel Model
In this section, we first derive a spatial-frequency channel model that incorporates the effects of spherical wavefront effect, SnS properties, and dual-wideband effects. Then, we quantitatively analyze the angular-delay characteristics of SnS dual-wideband XL-MIMO channels.
III-A Channel Modeling for SnS XL-MIMO Systems
In XL-MIMO systems, the use of ELAA and high-frequency bands facilitates near-field communications over hundreds of meters, invalidating the far-field plane wavefront assumption and necessitating the consideration of spherical wavefront effects. Additionally, SnS properties arise as different array portions observe varying propagation conditions, causing power variations across array elements. Furthermore, dual-wideband effects in wideband XL-MIMO systems influence channel characteristics, requiring careful modeling.
Assume that there are $L$ propagation paths between the BS and user. Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l>1$ indicates reflection or diffraction path, as illustrated in Fig. 1. Denote $\tau_{l,n}$ as the time delay of the $l$-th path for the $n$-th antenna. Then the baseband channel impulse response of the $n$-th received antenna is given by
$h_{n}(t)=\sum_{l=1}^{L}\tilde{\alpha}_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l,n}}\delta(t-\tau_{l,n}),$ (4)
where $\tilde{\alpha}_{l}\in\mathbb{C}$ denotes the complex path gain. $\kappa_{l,n}$ is introduced to characterize the SnS properties from the perspective of multipath propagation mechanisms, and it is given by [11]
$\kappa_{l,n}\begin{cases}=0,\,n\notin\boldsymbol{\phi}_{l},\\ =1,\,n\in\boldsymbol{\phi}_{l}\,\&\,\mathrm{LoS/Reflection},\\ >0,\,n\in\boldsymbol{\phi}_{l}\,\&\,\mathrm{Diffraction},\end{cases}$ (5)
where $\boldsymbol{\phi}_{l}=[n_{l,s},n_{l,e}]$ denotes the VR of the $l$-th path with $n_{l,s}$ and $n_{l,e}$ indicating the start and end index [^1] For the diffraction paths, $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ is associated with the diffraction coefficient. As such, $\kappa_{l,n}$ can be set within $10<\kappa_{l,n}<$ 1, with the reference value chosen as the element with the highest received power. Thus, we assume $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ obeys a uniform distribution between 0 and 1. Additionally, we assume continuous VRs for simplicity. Notably, the proposed method is not restricted to continuous VRs and can be extended to scenarios involving disjoint VRs, where the visible antennas may be distributed across multiple disjoint regions for a specific path.. Assume the distance and direction of the $l$-th path to the reference antenna element as $r_{l}$ and $\vartheta_{l}$, as illustrated in Fig. 1. Then, $\tau_{l,n}$, for any $l\in\mathcal{L}$, can be rewritten as
$\tau_{l,n}=\tau_{l}+\frac{r_{l}-r_{l,n}}{c},$ (6)
where $\tau_{l}=r_{l}/c$; $r_{l,n}$ denotes the distance between the $l$-th scatterer and the $n$-th antenna element, which is defined as
$r_{l,n}=$ $\sqrt{(r_{l}\cos\vartheta_{l}-nd)^{2}+r_{l}^{2}\sin^{2}\vartheta_{l}}$ (7)
$\overset{(a)}{\approx}$ $r_{l}-nd\cos\vartheta_{l}+\frac{n^{2}d^{2}}{2r_{l}}{\sin^{2}\vartheta_{l}}.$
where $(a)$ is obtained by using the Fresnel approximation, i.e., $\sqrt{1+x}\approx 1+x/2-x^{2}/8$. In this Taylor expansion, the current point is set at $n=0$, and terms of third order and higher are neglected. Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}+n{\psi_{l}}/{f_{c}}-n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$. According to (6) and (7), the continuous time channel impulse response in (4) can be further expressed as
$h_{n}(t)=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\delta\left(t-\tau_{l,n}\right),$ (8)
where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain. Applying the continuous time Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (10), as shown in the top of this page.
$h_{n}(f)=\int_{-\infty}^{+\infty}h_{n}(t)\mathrm{e}^{-\mathrm{j}2\pi ft}\mathrm{d}t=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}\mathrm{e}^{-\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (10)
Further, Eq. (10) can be expressed in a more compact form as
$\mathbf{h}(f)=\sum_{l=1}^{L}\alpha_{l}a(f,\tau_{l})(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l}))\odot\boldsymbol{\theta}(f,\psi_{l},\varphi_{l}),$ (11)
where $a(f,\tau_{l})=\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}$; $\boldsymbol{\kappa}_{l}=[\kappa_{l,1},\kappa_{l,2},\cdots,\kappa_{l,N}]^{\mathrm{T}}$ denotes the visibility indicator vector with the $n$-th entry being $\kappa_{l,n}$; $\mathbf{b}(\psi_{l},\varphi_{l})$ and $\mathbf{a}(f,\tau_{l})\}$ denote the array response vector and the frequency response vector, respectively, with each of their $n$-th entries given by
$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)},$ (12)
$[\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (13)
Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K$. The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in Eq. (11) with $f$ replaced by $f_{k}$. Thus, the overall spatial-frequency channel matrix $\mathbf{H}$ can be reformulated as
$\mathbf{H}=\sum_{l=1}^{L}\alpha_{l}\left(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})\right)\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (14)
where $\mathbf{a}(\tau_{l})=[a(f_{1},\tau_{l}),\cdots,a(f_{K},\tau_{l})]^{\mathrm{T}}$, and $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})=[\boldsymbol{\theta}(f_{1},\psi_{l},\varphi_{l}),\cdots,\boldsymbol{\theta}(f_{K},\psi_{l},\varphi_{l})]$.
Remark 1.
Eq. (14) presents a comprehensive channel model for XL-MIMO systems, integrating spherical wavefront effects, SnS properties, as well as spatial and frequency-wideband characteristics. Specifically, the spatial-domain steering vector $\mathbf{b}(\psi_{l},\varphi_{l})$ is coupled with the SnS indicator vector $\boldsymbol{\kappa}_{l}$, the frequency response vector $\mathbf{a}(\tau_{l})$, and the frequency-dependent phase matrix $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$. However, due to the high dimensionality of spatial-frequency channels in mmWave XL-MIMO systems, direct estimation becomes computationally prohibitive. Thus, it is essential to leverage channel sparsity in a suitable transform domain.
III-B Angular-Delay Representation of SnS XL-MIMO Channel
Building on approaches used for traditional massive MIMO channels [6], the wideband XL-MIMO channels in (14) can be approximated as
$\mathbf{H}\approx\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}.$ (15)
where $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N\times N}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ denote the normalized $N$- and $K$-dimensions DFT matrices. $\mathbf{X}\in\mathbb{C}^{N\times K}$ denotes the angular-delay channel.
Lemma 1. The angular-delay channel $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{*}$ is a sparse matrix containing only $L$ non-zero blocks, with the non-zeros elements staying inside a square region, i.e.,
$\lim_{N,K\to\infty}[\mathbf{X}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{2},\\ 0,&(n,k)\notin\mathcal{A}_{2},\end{cases}$ (16)
where $\mathcal{A}_{2}=\cup_{l,p}\mathcal{A}_{l,2,p}$ with
$\mathcal{A}_{l,2,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (17)
$k$ $\left.=\mathrm{mod}(k_{l}+f_{s}\tau_{l},K),\forall(n_{l},k_{l})\in\mathcal{A}_{l,1}\right\},$
and $\mathrm{mod}(a,m)$ is the modulus of $a$ for $m$. In addition, $\epsilon_{l,p}=\psi_{l}-2n_{l,e}\varphi_{l}+\eta_{l}(p-1)$ with $\eta_{l}={2N\varphi_{l}}/{P}$ and $p\in\{p\in\mathbb{N}^{+}\mid\psi_{l}-2n_{l,e}\varphi_{l}\leq\eta_{l}p\leq\psi_{l}-2n_{l,s}\varphi_{l}\}$. $\mathcal{A}_{l,1}\triangleq\left\{(n_{l},k_{l})\in\mathbb{Z}^{2}\mid I_{l,e}\leq n_{l}\leq I_{l,s},J_{l,e}\leq k_{l}\leq J_{l,s}\right\}$ with $I_{l,e}=\min_{k}\lceil(i_{l,e}(k)+1)N/2\rceil$, $I_{l,s}=\max_{k}\lceil(i_{l,s}(k)+1)N/2\rceil$, $J_{l,e}=\min_{n}j_{l,e}(n)$, and $J_{l,s}\max_{n}j_{l,s}(n)$, $\forall n\in\boldsymbol{\phi}_{l}$, where
$i_{l,e}(k)$ $=2(a_{k}-b_{k}n_{l,e}),\quad i_{l,s}(k)=2(a_{k}-b_{k}n_{l,s}),$ (18)
$j_{l,e}(n)$ $=(t_{l,n}-{2}/K)K,\quad j_{l,s}(n)=(t_{l,n}+{2}/K)K,$ (19)
with $a_{k}={f_{k}\psi_{l}}/{f_{c}}$, $b_{k}={2f_{k}\varphi_{l}}/{f_{c}}$, and $t_{l,n}=(n{f_{s}\psi_{l}}-n^{2}{f_{s}\varphi_{l}})/(Kf_{c})$.
From (17), the non-zero region of $\mathbf{X}$ is synthesized from the shifted non-zero regions of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}$ across angular and delay domains for all $l$. The shift distance are jointly determined by $N\epsilon_{l,p}$ and $f_{s}\tau_{l}$. In addition, The index of significant angular-delay components $(n,k)$ within $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ form a parabolic shape, which arises because both $j_{l,e}(k)$ and $j_{l,s}(k)$ are quadratic functions of $n$.
Proof: Please see Appendix A.
Moreover, we analyze the effects of the spherical wavefront, SnS properties, and dual-wideband effects on the sparsity structure of $\mathbf{X}$ based on Lemma 1 from the perspectives of the angular and delay spread. 1) Angular Spread: The spherical wavefront introduces angular spread of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ in angular domain, ranging from $a_{k}-b_{k}n_{l,e}$ to $a_{k}-b_{k}n_{l,s}$ due to the inclusion of a quadratic term, $\varphi_{l}n^{2}$. Meanwhile, it affects the variations of $\epsilon_{l,p}$. The SnS properties influence the values of $n_{l,e}$ and $n_{l,s}$, thereby affecting the extent of the angular spread. Additionally, the frequency-wideband effect induces variations in $i_{l,e}(k)$ and $i_{l,s}(k)$ across subcarriers $k$, which further broadens the angular dispersion across the system bandwidth. 2) Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread. The SnS properties define the range of variation for $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively influence the angular and delay spreads, reshaping the sparse patterns of SnS dual-wideband channels in the angular-delay domain, i.e.,
  • Global Block Sparsity: The angular-delay channel contains $L$ non-zero block corresponding to $L$ propagation paths. The index range of each non-zero block is determined by $\mathcal{A}_{l,2}$.
  • Local Common-Delay Sparsity: Due to the spherical wavefront effect, within $\mathcal{A}_{2}$, different delay indices correspond to varying angular spreads, which occur in clusters and exhibit high correlation.
Remark 2.
The global block sparsity and local common-delay sparsity arise from the combined effects of the spherical wavefront, SnS properties, and dual-wideband characteristics. These features inherently embed inter-antenna and inter-subcarrier correlations into the angular-delay channel. Furthermore, (15) offers an alternative representation of the spatial-frequency channel, facilitating the estimation of SnS dual-wideband XL-MIMO channels. Once the angular-delay channel is reconstructed, key parameters such as AoAs, VRs, and the range of beam squint can be efficiently extracted.
To enhance intuitive understanding of the SnS XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$. Fig. 2 illustrates the block sparsity of the SnS XL-MIMO channel in the angular-delay domain. Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,e}(64)=513$, and $j_{l,s}(64)\approx 517$. Similarly, $i_{l,e}(1023)\approx 0.88$, $i_{l,s}(1023)\approx 0.72$, $j_{l,e}(959)=545$, and $j_{l,s}(959)\approx 549$. As depicted in Fig. 2(a) and Fig. 2(b), these values provide accurate boundaries for the ranges of angular and delay spread. Additionally, the figures show that both the angular and delay spectra shift to the right due to $\cos\vartheta_{l}>0$ and the monotonicity of $t_{n}$ with respect to $n$.
Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, and $J_{l,s}=549$. Fig. 2(c) illustrates the non-zero region of $\mathbf{F}^{\mathrm{A}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l,1}$, validating the effectiveness of Lemma 1. Additionally, we have $f_{s}\tau_{l}=100$, $\epsilon_{l,1}=\psi_{l}-2n_{l,s}\varphi_{l}\approx 0.4$, and $\epsilon_{l,P}=\psi_{l}-2n_{l,e}\varphi_{l}=0.32$, corresponding to $N\epsilon_{l,1}=410$ and $N\epsilon_{l,P}=328$, respectively. Consequently, Fig. 2(d) and Fig. 2(e) represent shifted versions of Fig. 2(c). The shift distances in the angular domain are $696$ and $614$, respectively, while the shift distance in the delay domain is $100$. Finally, by concentrating the shifted versions across all $\epsilon_{l,p}$, the sparse structure depicted in Fig. 2(f) is obtained. It can be observed that the angular range of $\mathcal{A}_{l,2}$ is broader than that of $\mathcal{A}_{l,1}$ due to the composite effect.
(a) Angular spread for $k=0$ and $K-1$.
(b) Delay spread for $n=0$ and $N_{\mathrm{R}}-1$.
(c) $\mathcal{A}_{l,1}$.
(d) $\mathcal{A}_{l,p,1}$ with $p=1$.
(e) $\mathcal{A}_{l,p,1}$ with $p=P$
(f) $\mathcal{A}_{l,2}$.
Figure 2: Illustration of block sparsity of SnS XL-MIMO channel in the angular-delay domain.
Remark 3.
Compared to conventional massive MIMO-OFDM channels, the proposed SnS dual-wideband XL-MIMO channel exhibits distinct behavior. Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular diffusion. Moreover, the quadratic phase variations introduced by the spherical wavefront reshape the trend of delay variations across antennas. Notably, Lemma 1 is applicable to conventional MIMO-OFDM channels. When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted. Thus, (18) simplifies to $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$. Consequently, the significant components are uniformly distributed within the non-zero region, as illustrated in [6].
IV SnS Dual-Wideband Channel Estimation for XL-MIMO Systems
In this section, leveraging the sparsity outlined in Lemma 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task. Subsequently, to account for the global block sparsity and local common-delay sparsity, we introduce a novel column-wise sparse prior model tailored to this structure.
IV-A Problem Formulation
Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a multiple measurement vectors (MMV) based sparse recovery problem. Utilizing the angular-delay representation in (15), the received signal model in (3) can be further written as
$\mathbf{Y}=\mathbf{W}\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}+{\mathbf{N}}=\boldsymbol{\Psi}\mathbf{X}\mathbf{F}_{\mathrm{D}}+{\mathbf{N}},$ (20)
where $\boldsymbol{\Psi}\triangleq\mathbf{W}\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{M\times N}$. Furthermore, utilizing the unitary property of $\mathbf{F}_{\mathrm{D}}$, (20) can be reformulated as
$\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}=\boldsymbol{\Psi}\mathbf{X}+\tilde{\mathbf{N}},$ (21)
where $\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\in\mathbb{C}^{M\times K}$ and $\tilde{\mathbf{N}}={\mathbf{N}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\in\mathbb{C}^{M\times K}$ denotes the equivalent received pilot signal and noise matrix.
In this paper, our primary objective is to develop an effective estimation algorithm for accurately reconstructing $\mathbf{X}$ or $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$. In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (21), including orthogonal matching pursuit (OMP) [28, 29] and optimization-based methods [30, 31]. However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional parameters, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available.
Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and sparse Bayesian learning (SBL) have shown superior recovery performance in channel estimation tasks [32, 33, 27, 34]. However, the effectiveness of these Bayesian methods relies heavily on the accuracy of the prior model. To address this, we propose a novel column-wise hierarchical sparse prior model, as detailed in the next section.
IV-B Column-Wise Hierarchical Sparse Prior Model
To capture the global and local common-delay sparsity characteristics of $\mathbf{X}$, the column-wise hierarchical prior model is characterized as
$p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})=p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S}),$ (22)
where the first layer is modeled as a complex Gaussian distribution with mean zero and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
$p(\mathbf{X}|\boldsymbol{\Gamma})=\prod_{n=1}^{N}\prod_{k=1}^{K}\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1}).$ (23)
In the second layer, each $\gamma_{n,k}$ is assigned a conditionally Bernoulli distribution given by
$p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})$ $=\prod_{n=1}^{N}\prod_{k=1}^{K}\delta\left(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k})\right),$ (24)
where $t(s_{n,k},\boldsymbol{\alpha}_{k})=\delta(1-s_{n,k})\alpha_{k}^{1}+\delta(1+s_{n,k})\alpha_{k}^{2}$ with $\boldsymbol{\alpha}_{k}=[\alpha_{k}^{1},\alpha_{k}^{2}]^{\mathrm{T}}$ and $\delta(\cdot)$ indicating the Dirac delta function. The precision parameter $\boldsymbol{\Omega}=[\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2},\cdots,\boldsymbol{\alpha}_{K}]^{\mathrm{T}}\in\mathbb{R}^{2\times K}$ of $\boldsymbol{\Gamma}$ is characterized by
$p(\boldsymbol{\Omega})=\prod_{i=1}^{2}\prod_{k=1}^{K}p(\alpha_{k}^{i})=\prod_{i=1}^{2}\prod_{k=1}^{K}\mathrm{Ga}(\alpha_{k}^{i};a_{i},b_{i}),$ (25)
In the third layer, to capture the global block sparsity, the variance state variable $\mathbf{S}$ is modeled as a MRF
$p(\mathbf{S})=\left(\prod_{n,k}\prod_{(n^{\prime},k^{\prime})\in\mathcal{D}_{n,k}}u(s_{n,k},s_{n^{\prime},k^{\prime}})\right)^{\frac{1}{2}}\prod_{n,k}v(s_{n,k}),$ (26)
where $\mathcal{D}_{n,k}$ denotes the element set of neighbor nodes of $s_{n,k}$ with $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$, and $v(s_{n,k})=\exp(-\eta s_{n,k})$; $\varpi$ and $\eta$ are the parameters of $p(\mathbf{S})$.
In the following, we discuss how prior models in (22)-(26) capture the structured sparsity in $\mathbf{X}$. 1) Variance-Driven Sparsity: The variance $\gamma_{n,k}^{-1}$ controls the sparsity of $x_{n,k}$, with values approaching zero effectively setting $x_{n,k}$ to zero. 2) Global Block Sparsity: The MRF models the variance states, encouraging block patterns while suppressing isolated coefficients that deviate from their neighbors. 3) Local Common-Delay Sparsity: As described in (24) and (25), the common-delay sparsity mechanism is introduced to ensure that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter. This enforces a unified sparsity structure among components associated with the same delay.
IV-C Bayesian Inference
Since the measurement matrix $\boldsymbol{\Psi}$ may be “bad” (e.g., rank-deficient, ill-conditioned, or having a non-zero mean) [32], the divergence issues might be arisen in the Bayesian inference. To address this, we first perform unitary transformations on the measurement matrix, i.e., $\boldsymbol{\Psi}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$. Consequently, the original signal model in (21) is rewritten as
$\mathbf{R}=\mathbf{A}\mathbf{X}+{\mathbf{W}}=\mathbf{Z}+{\boldsymbol{\Xi}},$ (27)
where $\mathbf{R}=\mathbf{U}^{\mathrm{H}}\tilde{\mathbf{Y}}$, $\mathbf{A}=\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$, $\mathbf{Z}=\mathbf{A}\mathbf{X}$, and $\boldsymbol{\Xi}=\mathbf{U}^{\mathrm{H}}{\tilde{\mathbf{N}}}$.
Based on the prior model provided in (22), the minimum mean square error (MMSE) estimator for the $(n,k)$-th entry of $\mathbf{X}$ can be expressed as
$\hat{x}_{n,k}=\int{x}_{n,k}p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\mathrm{d}\beta\mathrm{d}\mathbf{Z}\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\mathrm{d}\mathbf{\boldsymbol{\Omega}}\mathrm{d}\mathbf{S}\mathrm{d}\mathbf{X},$ (28)
where $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ denotes the joint posterior probability of $\beta$, $\mathbf{Z}$, $\mathbf{X}$, $\boldsymbol{\Gamma}$, $\boldsymbol{\Omega}$, and $\mathbf{S}$, satisfying
$p(\beta,\mathbf{X},\mathbf{Z},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\propto p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S},\mathbf{R})$ (29)
$=$ $p(\mathbf{R}|\mathbf{Z},\beta)p(\mathbf{Z}|\mathbf{X})p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S})p(\beta),$
where $p(\beta)\propto\beta^{-1}$ denotes the noise precise level; $p(\mathbf{R}|\mathbf{Z},\beta)$ and $p(\mathbf{Z}|\mathbf{X})$ are given by
$p(\mathbf{R}|\mathbf{Z},\beta)$ $=\frac{1}{(\pi\beta^{-1})^{MK}}\prod_{m}\prod_{k}p(r_{m,k}\mid z_{m,k}),$ (30)
$p(\mathbf{Z}|\mathbf{X})$ $=\delta(\mathbf{Z}-\mathbf{AX}),$ (31)
where $p(r_{m,k}|z_{m,k})=\mathcal{CN}(z_{m,k};r_{m,k},\beta^{-1})$; ${r}_{m,k}$ and ${z}_{m,k}$ denote the $(m,k)$-th elements of $\mathbf{R}$ and $\mathbf{Z}$. The dependencies of the random variables in the factorization (29) can be shown by a factor graph as depicted in Fig. 3, where the involved probability distributions are listed in Table I.
Due to the large number of antennas and subcarriers in XL-MIMO systems, solving problem (28) requires evaluating high-dimensional integrals, which is computationally prohibitive. Moreover, traditional SBL-based methods are infeasible in this context as they involve high-dimensional matrix inversions. In addition, the introduction of the common-delay sparsity mechanism and the MRF results in a more complex factor graph representation, making existing message passing algorithms [32, 35, 33, 27] less suitable. Specifically, [32, 35, 33] do not incorporate the common-delay sparsity mechanism or MRF-based variance state modeling, leading to a mismatch between these algorithms and the factor graph representation in Fig. 3. Additionally, [27] lacks the common-delay sparsity mechanism and uses a first-moment matching method to update the prior parameters, which neglects higher-order dependencies among variable nodes. To address these limitations, we propose a novel MMV-VMP algorithm utilizing the variational message passing and belief propagation in the next section.
Figure 3: Factor graph representation for the factorization (29).
TABLE I: Factor and Distribution in (29)
Factor Distribution Function
$f_{\beta}$ $p(\beta)$ $\beta^{-1}$
$f_{r_{m,k}}$ $p(r_{m,k}|z_{m,k},\beta)$ $\mathcal{CN}(r_{m,k};z_{m,k},\beta^{-1})$
$f_{z_{m,k}}$ $p(z_{m,k}|\mathbf{x}_{k})$ $\delta(z_{m,k}-\mathbf{A}_{m,:}\mathbf{x}_{k})$
$f_{x_{n,k}}$ $p(x_{n,k}|\gamma_{n,k})$ $\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1})$
$f_{\gamma_{n,k}}$ $p(\gamma_{n,k}|\boldsymbol{\alpha}_{k},s_{n,k})$ $\delta(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k}))$
V Proposed MMV-VMP Algorithm
The MMV-VMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (28). In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules as shown in Fig. 3.
V-A Measurement Module
Measurement module aims to obtain the likelihood estimation of $x_{n,k}$ based on the observation $\mathbf{R}$. Denote the belief of $z_{m,k}$ as $b(z_{m,k})\sim\mathcal{CN}(z_{m,k};\hat{z}_{m,k},\hat{\nu}^{z}_{m,k})$, which is defined in (38) with $\hat{z}_{m,k}$ and $\hat{\nu}^{z}_{m,k}$ indicating the $m$-th elements of $\hat{\mathbf{z}}_{k}$ and $\hat{\boldsymbol{\nu}}_{\mathbf{z}_{k}}$, respectively. In this manner, according to the variation message passing rule, the message from $f_{r_{m,k}}$ to $\beta$ is given by
$m_{f_{r_{m,k}}\rightarrow\beta}$ $\propto\exp\left\{\int\ln f_{r_{m,k}}b(z_{m,k})\mathrm{d}z_{m,k}\right\},$ (32)
$\overset{(a)}{\propto}\beta\exp\left\{-\beta\left(\left|r_{m,k}-\hat{z}_{m,k}\right|^{2}+\hat{\nu}^{z}_{m,k}\right)\right\},$
where $(a)$ is obtained by utilizing the first-order and second-order moment properties of Gaussian distribution.
Concentrating all messages that are input to $\beta$, the belief of $\beta$ is given by
$b(\beta)\propto f_{\beta}(\beta)\prod_{m=1}\prod_{k=1}m_{f_{r_{m,k}}\rightarrow\beta}(\beta),$ (33)
$\propto$ $\beta^{MK-1}\exp\left\{-\beta\sum_{m=1}\sum_{k=1}\left(\left|r_{m,k}-\hat{z}_{m,k}\right|^{2}+\hat{\nu}^{z}_{m,k}\right)\right\},$
According to (33), it is observed that the belief $b(\beta)$ obeys the Gamma distribution with shape parameters $MK$ and $\sum_{m=1}\sum_{k=1}\left(\left|r_{m,k}-\hat{z}_{m,k}\right|^{2}+\hat{\nu}^{z}_{m,k}\right)$. Thus, the approximate posterior mean $\hat{\beta}=\int\beta b(\beta)\mathrm{d}\beta$ is given by
$\hat{\beta}=\frac{MK}{\sum_{m=1}\sum_{k=1}\left(\left|r_{m,k}-\hat{z}_{m,k}\right|^{2}+\hat{\nu}^{z}_{m,k}\right)}.$ (34)
Similar to (32), the forward message passing from $f_{r_{m,k}}$ to $z_{m,k}$ can be given by
$m_{f_{r_{m,k}}\rightarrow z_{m,k}}$ $\propto\exp\left\{\int b(\beta)\ln f_{r_{m,k}}\mathrm{d}\beta\right\}$ (35)
$\propto\exp\left\{-\left|r_{m,k}-z_{m,k}\right|^{2}\int\beta b(\beta)\mathrm{d}\beta\right\}$
$\overset{(a)}{\propto}\mathcal{CN}(z_{m,k};r_{m,k},\hat{\beta}^{-1}),$
where $(a)$ is obtained by utilizing (34). Owing to the Gaussian form of the message in (35), we can obtain the following model $\mathbf{r}_{k}=\mathbf{z}_{k}+\boldsymbol{\xi}_{k}$, where $\boldsymbol{\xi}_{k}\sim\mathcal{CN}(\boldsymbol{\xi}_{k};0,\hat{\beta}^{-1}\mathbf{I}_{M\times M})$. This representation facilitates seamless integration with the forward passing of UAMP in the measurement module.
Denote the posterior mean and variance of $\mathbf{x}_{k}$ as $\hat{\mathbf{x}}_{k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$, which is defined in (55) and (56). According to the UAMP rules [33], define $\boldsymbol{\lambda}$ as $\left|\mathbf{A}\right|^{2}\mathbf{1}_{N}$. Thus, the message from $\mathbf{z}_{k}$ to $f_{\mathbf{r}_{k}}$ is given by $\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})$ with
$\boldsymbol{\nu}_{\mathbf{p}_{k}}=\boldsymbol{\lambda}\hat{\nu}_{\mathbf{x}_{k}},\quad\mathbf{p}_{k}=\mathbf{A}\hat{\mathbf{x}}_{k}-\boldsymbol{\nu}_{\mathbf{p}_{k}}\odot\boldsymbol{\mu}_{k},$ (36)
where $\odot$ denotes the element product; $\boldsymbol{\mu}_{k}$ is a intermediate vector, which is updated by
$\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}=\mathbf{1}_{M}./(\mathbf{1}_{M}\hat{\beta}^{-1}+\boldsymbol{\nu}_{\mathbf{p}_{k}}),\boldsymbol{\mu}_{k}=\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}\odot(\mathbf{r}_{k}-\mathbf{p}_{k}).$ (37)
In this case, utilizing the sum-product rule, the belief of $\mathbf{z}_{k}$ is given by
$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})$ (38)
$=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\mathcal{CN}(\mathbf{z}_{k};\mathbf{r}_{k},\hat{\beta}^{-1}\mathbf{I}_{M\times M})$
$\overset{(a)}{\propto}\mathcal{CN}(\mathbf{z}_{k};\hat{\mathbf{z}}_{k},\boldsymbol{\nu}_{\mathbf{z}_{k}}),$
where $(a)$ is obtained by the property of Gaussian distribution, i.e., $\mathcal{CN}(x;\mu_{1},\nu_{1})\mathcal{CN}(x;\mu_{2},\nu_{2})\propto\mathcal{CN}(x;\mu,\nu)$ with $\nu=(1/\nu_{1}+1/\nu_{2})^{-1}$ and $\mu=\nu(\mu_{1}/\nu_{1}+\mu_{2}/\nu_{2})$; In addition, $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}$ are respectively given by
${\boldsymbol{\nu}}_{\mathbf{z}_{k}}$ $=\boldsymbol{\nu}_{\mathbf{p}_{k}}./(\mathbf{1}+\hat{\beta}\boldsymbol{\nu}_{\mathbf{p}_{k}}),$ (39)
$\hat{\mathbf{z}}_{k}$ $=(\mathbf{p}_{k}+\hat{\beta}\boldsymbol{\nu}_{\mathbf{p}_{k}}\odot\mathbf{r}_{k})./(\mathbf{1}+\hat{\beta}\boldsymbol{\nu}_{\mathbf{p}_{k}}).$ (40)
Moreover, the output message from measurement module is given by
$\boldsymbol{\nu}_{\mathbf{q}_{k}}=\mathbf{1}./(\left|\mathbf{A}\right|^{2}\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}),\quad\mathbf{q}_{k}=\hat{\mathbf{x}}_{k}+\boldsymbol{\nu}_{\mathbf{q}_{k}}\odot(\mathbf{A}^{\mathrm{H}}\boldsymbol{\mu}_{k}).$ (41)
V-B Sparse Prior Module
The aim of the prior module is to update the distribution of ${x}_{n,k}$ according to extrinsic message $m_{x_{n,k}\rightarrow f_{x_{n,k}}}(x_{n,k})\propto\mathcal{CN}({x}_{n,k};{q}_{n,k},{\nu}^{q}_{n,k})$ and calculate the posterior estimation $\hat{{x}}_{n,k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$, where ${q}_{n,k}$ and ${\nu}^{q}_{n,k}$ denotes the $n$-th element of $\mathbf{q}_{k}$ and $\boldsymbol{\nu}_{\mathbf{q}_{k}}$, respectively.
V-B1 Forward Message Passing
Denote the belief of $x_{n,k}$ as $b(x_{n,k};\hat{x}_{n,k},\hat{\nu}_{\mathbf{x}_{k}})$, defined in (54). Utilizing the variational message passing rule, similar to (32), the message from $f_{x_{n,k}}$ to $\gamma_{n,k}$ is given by
$m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}\propto$ $\exp\left\{\int{b}(x_{n,k})\ln f_{x_{n,k}}\mathrm{d}x_{n,k}\right\}$ (42)
$\propto$ $\gamma_{n,k}\exp(-\gamma_{n,k}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})).$
Further, according to sum-product rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}\overset{(a)}{\propto}\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}b(\boldsymbol{\alpha}_{k})\mathrm{d}\gamma_{n,k}\mathrm{d}\boldsymbol{\alpha}_{k}$ (43)
$=\pi^{\mathrm{out}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{out}}_{n,k})\delta(1+s_{n,k}),$
where $(a)$ is obtained by utilizing the approximation $b(\boldsymbol{\alpha}_{k})=b(\alpha^{1}_{k})b(\alpha^{2}_{k})\approx\prod_{j\neq n}m_{\boldsymbol{\alpha}_{k}\rightarrow f_{\gamma_{j,k}}}$ and the expectation of Gamma distribution; $\pi^{\mathrm{out}}_{n}$ is defined as
$\pi^{\mathrm{out}}_{n,k}=\frac{\hat{a}_{k}^{1}(\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))}{\hat{a}_{k}^{1}(\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))+\hat{a}_{k}^{2}(\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})}.$
Utilizing the message $m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})$, we further derive the message update in the MRF with a 4-connect scheme, where $s_{n,k}^{{\text{l}}}\triangleq s_{n,k-1}$, $s_{n,k}^{\text{r}}\triangleq s_{n,k+1}$, $s_{n,k}^{{\text{t}}}\triangleq s_{n-1,k}$, and $s_{n,k}^{{\text{b}}}\triangleq s_{n,k+1}$ denote the left, right, top and bottom neighbors of $x_{n,k}$. The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions. Take $m_{n,k}^{\mathrm{l}}$ as an example, according to the sum-product rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (45), as shown in the top of next page.
$\lambda_{n,j}^{\mathrm{l}}=\frac{\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}}\mathrm{e}^{\varpi-\eta}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}\lambda^{w}_{s_{n,j}^{{\text{l}}}}+(1-\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}})\mathrm{e}^{\eta-\varpi}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}(1-\lambda^{w}_{s_{n,j}^{{\text{l}}}})}{(\mathrm{e}^{\varpi}+\mathrm{e}^{-\varpi})\left(\mathrm{e}^{-\eta}\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}\lambda^{w}_{s_{n,j}^{{\text{l}}}}+\mathrm{e}^{\eta}(1-\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}})\prod_{w\in\left\{\mathrm{l,t,b}\right\}}(1-\lambda^{w}_{s_{n,j}^{{\text{l}}}})\right)}.$ (45)
The other three messages can be obtained in a similar way.
V-B2 Backward Message Passing
With the messages of neighbors and $v(s_{n,k})$, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
$m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}$ $=\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}m_{n,k}^{\mathrm{l}}v(s_{n,k})$ (46)
$\propto\pi^{\mathrm{in}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{in}}_{n,k})\delta(1+s_{n,k})$
where $\pi^{\mathrm{in}}_{n,k}$ is defined as
$\frac{\mathrm{e}^{-\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}\lambda_{n,k}^{w}}{\mathrm{e}^{-\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}\lambda_{n,k}^{w}+\mathrm{e}^{\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}(1-\lambda_{n,k}^{w})}.$ (47)
According to sum-product rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto$ $\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}\mathrm{d}\gamma_{n,k}\mathrm{d}s_{n,k},$
$\overset{(a)}{\propto}$ $\alpha_{k}^{1}\exp(-\alpha_{k}^{1}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})),$ (48)
where $(a)$ is obtained by utilizing the property of delta function, i.e., $\int\delta(x-t)f(x)\mathrm{d}x=f(t)$. Similarly, $m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{2}}\propto\alpha_{k}^{2}\exp(-\alpha_{k}^{2}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$. As a result, the belief of $\alpha_{k}^{1}$ is given by
${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (49)
where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. According to the sum-product rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}=\pi^{\mathrm{in}}_{n,j}{b}(\alpha^{1}_{k})+(1-\pi^{\mathrm{in}}_{n,j}){b}(\alpha^{2}_{k}).$ (50)
Consequently, the belief of $\gamma_{n,k}$ is given by
$b(\gamma_{n,k})\propto m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})$ (51)
$\propto$ $\pi^{\mathrm{in}}_{n,j}\gamma_{n,k}^{\hat{a}_{k}^{1}}\exp(-\gamma_{n,k}(\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$
$+$ $(1-\pi^{\mathrm{in}}_{n,k})\gamma_{n,k}^{\hat{a}_{k}^{2}}\exp(-\gamma_{n,k}(\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})).$
As a result, the message from $f_{x_{n,k}}$ to $x_{n,k}$ is denoted as
$m_{f_{x_{n,k}}\rightarrow x_{n,k}}(x_{n,k})$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (52)
$\propto\exp\left\{-\int\left|x_{n,k}\right|^{2}\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}\right\}$
$\overset{(a)}{\propto}\mathcal{CN}(x_{n,k};0,\hat{\gamma}_{n,k}^{-1}),$
where $(a)$ is obtained by
$\hat{\gamma}_{n,j}$ $=\int\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}=\pi^{\mathrm{in}}_{n,k}\frac{\hat{a}_{k}^{1}+1}{\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}$ (53)
$+(1-\pi^{\mathrm{in}}_{n,k})\frac{\hat{a}_{k}^{2}+1}{\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}.$
Combining the message from measurement module as $\mathcal{CN}({x}_{n,k};{q}_{n,k},\nu_{n,k}^{q})$. Thus, the approximate posterior distribution of $x_{n,j}$ can be approximated as
$b(x_{n,k})$ $\propto\mathcal{CN}({x}_{n,k};{q}_{n,k},\nu_{n,k}^{q})\mathcal{CN}(x_{n,k};0,\hat{\gamma}_{n,k}^{-1})$ (54)
$\overset{(a)}{\propto}\mathcal{CN}(x_{n,k};\hat{x}_{n,k},\hat{\nu}_{\mathbf{x}_{k}}),$
where $(a)$ is obtained similar to (38), and the approximate posterior mean and variance of $x_{n,k}$ are respectively given by
$\hat{\nu}_{n,k}^{x}=\frac{\nu_{n,k}^{q}}{1+\nu_{n,k}^{q}\hat{\gamma}_{n,k}},\quad\hat{x}_{n,k}=\frac{q_{n,k}}{1+\nu_{n,k}^{q}\hat{\gamma}_{n,k}}.$ (55)
Performing the average operations to $\nu_{n,k}^{x}$, we further have
$\hat{\nu}_{\mathbf{x}_{k}}=\frac{1}{N}\sum_{n=1}^{N}\hat{\nu}_{n,k}^{x}.$ (56)
Algorithm 1 Proposed UAMP-SBL-MRF algorithm
1:received vector $\mathbf{R}$, measurement matrix $\mathbf{A}$.
2:Initialize: $\hat{\nu}_{\mathbf{x}_{k}}^{(0)}=1$, $\hat{\mathbf{x}}=\mathbf{0}$, $\hat{{\gamma}}_{n,k}={1},\hat{\beta}=1$, and $\boldsymbol{\mu}=\mathbf{0}$.
3:while the stopping criterion is not met do
4:/*Measurement module*/
5:     Update the $\boldsymbol{\nu}_{\mathbf{p}_{k}}$ and $\mathbf{p}_{k}$ according to (36);
6:     Update the $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}$ and $\boldsymbol{\mu}_{k}$ according to (37);
7:     Update the $\boldsymbol{\nu}_{\mathbf{z}_{k}}$ and $\mathbf{z}_{k}$ according to (39) and (40);
8:     Update $\hat{\beta}$ according to (34);
9:     Update the $\boldsymbol{\nu}_{\mathbf{q}_{k}}$ and $\mathbf{q}_{k}$ according to (41);
10:/*Prior Module*/
11:     Update the messages $m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}$ according to (43);
12:     Update the messages $m_{n}^{\mathrm{l}}$, $m_{n}^{\mathrm{r}}$, $m_{n}^{\mathrm{t}}$, and $m_{n}^{\mathrm{b}}$;
13:     Update the messages $m_{s_{n}\rightarrow f_{\gamma_{n}}}(s_{n})$ according to (47);
14:     Update the belief $b(\alpha_{k}^{1})$ and $b(\alpha_{k}^{2})$ according to (49);
15:     Update $\hat{\gamma}_{n}$ according to (53);
16:     Update $\hat{x}_{n,k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$ according to (55) and (56).
17:end while
18:$\hat{x}_{n,k}$.
The proposed MMV-VMP algorithm is summarized in Algorithm 1 and it can be terminated when it reached a maximum number of iteration or the difference between the estimates of two consecutive iterations is less than $10^{-5}$. In the following, we provide the computational complexity analysis for the proposed MMV-VMP algorithm. Examining the steps of Algorithm 1, it is evident that there is no matrix inversion involved. Thus, the most computationally intensive parts only involve matrix-vector products in lines 2 and 6, i.e., $\mathcal{O}(MN)$ per iteration. Consequently, the total complexity of the MMV-VMP algorithm is $\mathcal{O}(TKMN)$, where $T$ denotes the number of iterations.
TABLE II: Simulation Parameters
Notations Parameters
Number of BS antenna $N_{\mathrm{R}}$ 256
Number of RF chain $N_{\mathrm{RF}}$ 16
Carrier frequency $f_{c}$ 30GHz
Number of pilot carriers $K$ 64
System bandwidth $f_{s}$ 1.6GHz
Number of channel path $L$ 4
Angle of arrival $\vartheta_{l}$ $\mathcal{U}(-\pi/2,\pi/2)$
Distance between BS and UE or scatterers $r_{l}$ [5, 50]m
Proportion of visible antenna elements $\rho_{l}$ (0,1]
VI Simulation Results
In this section, we evaluate the performance of the proposed channel estimation scheme under various system setups. The simulation parameters are shown in Table II. In particular, we consider normalized mean square error (NMSE) as performance metrics, which is defined as $\mathrm{NMSE}\triangleq{\lVert\hat{\mathbf{H}}-\mathbf{H}\rVert^{2}_{\mathrm{F}}}/{\lVert\mathbf{H}\lVert^{2}_{\mathrm{F}}}$, where $\mathbf{H}$ and $\hat{\mathbf{H}}$ are the true channel and estimated channel, respectively. In addition, the SNR is defined in received side, which is given by $10\log_{10}\left(\lVert\mathbf{WH}\rVert^{2}_{\mathrm{F}}/\lVert\mathbf{N}\rVert^{2}_{\mathrm{F}}\right)$. Additionally, we compare the proposed MMV-VMP algorithm with the following baselines:
  • SOMP[28]: The simultaneous OMP algorithm designed for on-grid sparse recovery, which depends on the knowledge of the number of non-zeros components.
  • StdSBL[36]: The standard SBL algorithm, implemented within an expectation-maximization (EM) framework. The algorithm employs a two-layer Gaussian-Gamma hierarchical prior model, where the posterior estimates of $\mathbf{x}_{k}$ are updated in the E-step, and the hyperparameters of the prior model are updated in the M-step.
  • UAMP-SBL[33]: An improved version of the StdSBL algorithm that leverages the unitary AMP framework to perform the E-step and variational message passing to achieve the updates of prior parameters.
  • PC-SBL[37]: A variant of the StdSBL framework that incorporates a pattern-coupled Gaussian prior model to exploit the block sparsity inherent in signals.
  • VSP[27]: A variant of the StdSBL framework that employs a MRF-based hierarchical prior model to effectively capture the block sparsity of signals. Unlike the StdSBL and PC-SBL methods, the prior parameters in the VSP algorithm are updated using a moment-matching approach, providing a computationally efficient alternative.
VI-A Convergence and Complexity Performance
Figure 4: Convergence behavior of MMV-VMP.
Fig. 4 illustrates the convergence behavior of the proposed MMV-VMP algorithm by plotting the NMSE against the number of iterations under various pilot lengths and SNR conditions. The results clearly demonstrate a consistent monotonic decrease in NMSE across all iterations, indicating the algorithm’s stable and reliable convergence across different simulation scenarios. Based on the trade-off between estimation accuracy and computational complexity, it is observed that the NMSE stabilizes after approximately 20 iterations. Therefore, for subsequent simulations, the maximum number of iterations can be effectively limited to 20 without compromising performance.
Figure 5: Complexity evaluation of different algorithms.
Fig. 5 presents a computational complexity analysis by comparing the CPU runtime of various algorithms. Among these, SOMP exhibits the shortest runtime. This efficiency is attributed to its computational simplicity, involving only matrix-vector product between the residual and the measurement matrix, as well as a projection operation between the received signal and the low-dimensional basis matrix. In contrast, SBL-based methods such as StdSBL, PC-SBL, and VSP require significantly more computational time. This is primarily due to the matrix inversion operations inherent in their implementation, which increase computational complexity. In contrast, the proposed MMV-VMP algorithm demonstrates a shorter runtime compared to StdSBL, PC-SBL, and VSP due to avoidance of matrix inversion operations and stable convergence properties.
VI-B NMSE versus SNR and Pilot Symbol Number
Figure 6: NMSE versus pilot symbol number.
Fig. 6 illustrates the NMSE performance of various algorithms as a function of SNR for $P=8$. The results reveal that algorithms such as SOMP [29], StdSBL [36], and UAMP-SBL [33], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-VMP. Among block-sparsity prior-based methods, the PC-SBL algorithm leverages a pattern-coupled Gaussian prior model to effectively capture local common-delay sparsity. However, it does not consider global block sparsity. In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity. The proposed MMV-VMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both global block sparsity and local common-delay sparsity. This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
Figure 7: NMSE versus pilot symbol number.
Fig. 7 evaluates the estimation performance of various algorithms under different compression ratios by varying the number of pilot symbols, $P$, with the fixed SNR of 2 dB. As the number of pilot symbols varies from 6 to 12, the compression ratio $M/N$ ranges from 0.375 to 0.75. We can obtain two key observations: 1) Across most compression ratios, the proposed MMV-VMP algorithm outperforms baseline methods. This result highlights its robust sparse recovery capability under varying compression conditions. Moreover, the superior performance demonstrates that the proposed algorithm is a low-overhead solution, requiring fewer pilot symbols to achieve comparable or better performance than other algorithms. 2) As the pilot length increases, the rate of performance improvement gradually diminishes, eventually approaching saturation. This observation suggests that selecting a moderate value of $P$ is sufficient to strike a balance between estimation performance and pilot overhead.
VI-C NMSE versus Path Number
Figure 8: NMSE versus path number.
Fig. 8 illustrates the NMSE performance of various algorithms as a function of the number of propagation paths, with $P=8$ and $\mathrm{SNR}=2$dB. The SOMP algorithm exhibits significant sensitivity to prior knowledge of the number of paths, with its performance degrading sharply when the assumed and actual path numbers differ. In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios. Notably, the proposed MMV-VMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
VI-D Ablation Study to Evaluate the Effects of Prior Models
Figure 9: Ablation study of the column-wise hierarchical model.
To evaluate the effectiveness of the proposed column-wise prior model, the MMV-VMP algorithm is compared with two baselines: (1) a column-wise hierarchical prior without the MRF, labeled as “Column-Wise w.o. MRF,” and (2) a column-wise prior without the common-delay sparsity mechanism, labeled as “Column-Wise w.o. CD.” Fig. 9 reveals two key observations: 1) Both the MRF and the common-delay sparsity mechanism contribute to improved estimation performance. Notably, the improvement introduced by the MRF is more significant than that of the common-delay sparsity mechanism. 2) The simultaneous use of the MRF and the common-delay sparsity mechanism better captures the structured sparsity of the channel, leading to further enhancement in estimation accuracy.
To showcase the estimation performance of different prior more intuitively, Fig. 10 provides the results of a Monte Carlo simulation with $\mathrm{SNR}=2$ dB and $P=8$. The original spatial-delay channels are compared with the reconstructed channels obtained using “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and the proposed MMV-VMP algorithm. The figure clearly shows that the proposed MMV-VMP algorithm achieves the best visual quality, accurately reconstructing the spatial-delay channels and demonstrating its superior capability in sparse recovery. As expected, “Column-Wise w.o. MRF” provides the poorest reconstruction quality due to its limited ability to capture global block sparsity. Moreover, compared to “Column-Wise w.o. CD,” the reconstructed channels by MMV-VMP exhibit reduced power dispersion phenomenon near the non-zero regions, attributed to the incorporation of the common-delay sparsity mechanism.
(a) Original.
(b) Column-Wise w.o. MRF.
(c) Column-Wise w.o. CD.
(d) MMV-VMP.
Figure 10: Original and reconstructed channels by “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and MMV-VMP algorithm.
VII Conclusions
In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, spatially non-stationary (SnS) properties, and dual-wideband effects. We first rigorously quantified the angular and delay spread effects of the SnS dual-wideband XL-MIMO channel in the angular-delay domain. Building upon this, we highlighted the global and local common-delay sparsity inherent in the angular-delay channels. Leveraging this structured sparsity, we proposed a computationally efficient MMV-VMP algorithm, where variational message passing is employed to handle the complex factor graph arising from the column-wise hierarchical prior. Simulation results demonstrate that the proposed MMV-VMP algorithm consistently outperforms existing methods, offering robust performance with moderate computational complexity. Additionally, an ablation study was conducted to assess the effectiveness of the column-wise hierarchical prior, further validating the advantages of the proposed approach.
Appendix A Proof of Lemma 1
The proof of Theorem 1 is portioned as two part. We first prove that the angular-delay matrix of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ is block-sparse. Then, applying the shift property of 2D-IDFT, the block sparsity of $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}$ can be obtained. For notation simplification, we discard the subscript $l$, $\varphi_{l}$, and $\psi_{l}$ in the subsequent derivations. We begin by examining the angular-domain transformation of $\boldsymbol{\Theta}$ with $\cos\vartheta>0$. Mathematically, $\boldsymbol{\Theta}$ can be rewritten in a chirp signal form as
$[\boldsymbol{\Theta}]_{n,k}=\mathrm{e}^{-\mathrm{j}2\pi\left(a_{k}n-\frac{b_{k}}{2}n^{2}\right)},$ (A.1)
where $a_{k}={f_{k}\psi}/{f_{c}}$ and $b_{k}={2f_{k}\varphi}/{f_{c}}$ denote the starting spatial frequency and chirp rate, respectively. Utilizing the inherent chirp properties of Eq. (A.1), the IDFT of $\boldsymbol{\Theta}$ for the fixed $k$, i.e., $[\boldsymbol{\Xi}_{\mathrm{A}}]_{:,k}=\mathbf{F}^{\mathrm{H}}_{\mathrm{A}}\boldsymbol{\theta}(f_{k})\in\mathbb{C}^{N\times 1}$ exhibits a distinct rectangular shape [38], where $\mathbf{F}_{\mathrm{A}}$ denotes a $N$-dimension DFT matrix. The boundary spatial frequencies are given by
$i_{s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (A.2)
$i_{e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d},$ (A.3)
where $n_{s}$ and $n_{e}$ denote the start and ending index of VR $\boldsymbol{\phi}$, i.e., $\boldsymbol{\phi}\in[n_{s},n_{e}]$. In this case, due to $a_{k}>0$ and $b_{k}>0$, we have $i_{s}(k)>i_{e}(k)$. Consequently, $[\boldsymbol{\Xi}_{\mathrm{A}}]_{:,k}$ demonstrates block sparsity, with its prominent spatial frequency components concentrated from $i_{e}(k)$ to $i_{s}(k)$.
Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{e}^{\mathrm{min}}=\min_{k}i_{e}(k)$ to $i_{s}^{\mathrm{max}}=\max_{k}i_{s}(k)$. As a result, $\boldsymbol{\Xi}_{\mathrm{A}}$ is characterized as a row-wise block matrix
$\boldsymbol{\Xi}_{\mathrm{A}}=\left[\mathbf{0}_{K\times(I_{e}-1)}^{\mathrm{T}},\mathbf{U}_{K\times(I_{s}-I_{e}+1)}^{\mathrm{T}},\mathbf{0}_{K\times(N-I_{s})}^{\mathrm{T}}\right]^{\mathrm{T}},$ (A.4)
where $\mathbf{U}$ is a non-zero matrix; index $I_{e}$ and $I_{s}$ are given by $I_{e}=\lceil{(i_{e}^{\mathrm{min}}+1)N}/{2}\rceil$ and $I_{s}=\lceil{(i_{s}^{\mathrm{max}}+1)N}/{2}$.
Further, we examine the IDFT of $\boldsymbol{\Theta}$ in the delay domain, i.e., $\boldsymbol{\Xi}_{\mathrm{D}}=\boldsymbol{\Theta}\mathbf{F}_{\mathrm{D}}^{*}$, where $\mathbf{F}_{\mathrm{D}}$ denote a $K$-dimension DFT matrix. Specifically, we have
n,m $\propto\frac{1}{\sqrt{K}}\sum_{k=0}^{K-1}\mathrm{e}^{\mathrm{j}\frac{2\pi}{K}mk}\mathrm{e}^{-\mathrm{j}2\pi kt_{n}}$ (A.5)
$\propto\frac{1}{\sqrt{K}}\frac{\sin\left(\pi K(t_{n}-\frac{m}{K})\right)}{\sin\left(\pi(t_{n}-\frac{m}{K})\right)},$
where $t_{n}=n\frac{f_{s}\psi}{Kf_{c}}-n^{2}\frac{f_{s}\varphi}{Kf_{c}}$, which is a quadratic function about $n$. Further, we have
$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{n}-\frac{m}{K}\right).$ (A.6)
From (A.6), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{n}K$. However, due to the finite sampling size $K$ in practice, we consider the range from $j_{e}(n)=(t_{n}-{2}/{K})K$ to $j_{s}(n)=(t_{n}+{2}/{K})K$ to collect the delay indices near the peak effectively. Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
$\boldsymbol{\Xi}_{\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{e}-1)},\mathbf{L}_{N\times(J_{s}-J_{e}+1)},\mathbf{0}_{N\times(K-J_{s})}\right],$ (A.7)
where $J_{e}=\min_{n}j_{e}(n)$ to $J_{s}=\max_{n}j_{s}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
According to row-wise and column-wise block sparsity of (A.4) and (A.7), we assert that $\boldsymbol{\Xi}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ adheres to the following block-sparse structure, with its significant entries localized within the square region $\mathcal{A}_{1}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{e}\leq n\leq I_{s},J_{e}\leq k\leq J_{s}\right\}$, i.e.,
[…]
where $\mathbf{V}$ denotes a non-zero matrix.
Next, we examine the 2D-IDFT of channel $\mathbf{H}$, $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$. By leveraging the linearity of the 2D-IDFT, we can focus solely on analyzing the $l$-th path. Define $\mathbf{H}_{l}=\alpha_{l}\left(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})\right)\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}$, whose $(n,k)$-th element is given by
n,k $\propto\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}+f_{k}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k},$ (A.9)
$\overset{(a)}{\propto}\sum_{p=1}^{P}\mathrm{e}^{-\mathrm{j}2\pi\left(\frac{n}{N}N\epsilon_{l,p}+\frac{k}{K}f_{s}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k},$
where $(a)$ is obtained by decomposing the chirp signal $\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}$ into the superposition of $P$ complex exponential signals with frequencies $\epsilon_{l,p}=\psi_{l}-2n_{l,e}\varphi_{l}+\eta_{l}(p-1)$ with $\eta_{l}={2N\varphi_{l}}/{P}$ and $p\in\{p\in\mathbb{N}^{+}\mid\psi_{l}-2n_{l,e}\varphi_{l}\leq\eta_{l}p\leq\psi_{l}-2n_{l,s}\varphi_{l}\}$. Furthermore, we focus on $\mathbf{H}_{l,p}(n,k)=\mathrm{e}^{-\mathrm{j}2\pi\left(\frac{n}{N}N\epsilon_{l,p}+\frac{k}{K}f_{s}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k}$. Applying the shift property of 2D-IDFT, $\mathbf{X}_{l,p}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}_{l,p}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ should satisfy the block sparsity with its significant entries localized within the square region $\mathcal{A}_{l,2,p}$ with
$\mathcal{A}_{l,2,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (A.10)
$k$ $\left.=\mathrm{mod}(k_{l,1}+f_{s}\tau_{l},K),\forall(n_{l,1},k_{l,1})\in\mathcal{A}_{l,1}\right\}.$
Here, $\mathcal{A}_{l,1}$ is the non-zero region for $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$, and mod$(a,m)$ is the modulus of $a$ for $m$. Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{A}_{l,2}\triangleq\cup_{p}\mathcal{A}_{l,2,p}$. Thus, the proof is completed.

### v3

@@
 Abstract
-The deployment of extremely large antenna arrays (ELAAs) and operation at higher frequency bands in wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) systems introduce significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties.
-Combined with dual-wideband impacts, these effects fundamentally reshape the sparsity patterns of wideband XL-MIMO channels in the angular-delay domain, making existing sparsity-based channel estimation methods inadequate.
-To address these challenges, this paper revisits the channel estimation problem for wideband XL-MIMO systems, considering dual-wideband effects, spherical wavefront, and SnS properties.
-By leveraging the spatial-chirp property of near-field array responses, we quantitatively characterize the sparsity patterns of wideband XL-MIMO channels in the angular-delay domain, revealing global block sparsity and local common-delay sparsity.
-Building on this structured sparsity, we formulate the wideband XL-MIMO channel estimation problem as a multiple measurement vector (MMV)-based Bayesian inference task and propose a novel column-wise hierarchical prior model to effectively capture the sparsity characteristics.
-To enable efficient channel reconstruction, we develop an MMV-based variational message passing (MMV-VMP) algorithm, tailored to the complex factor graph induced by the hierarchical prior.
-Simulation results validate the proposed algorithm, demonstrating its convergence and superior performance compared to existing methods, thus establishing its effectiveness in addressing the challenges of wideband XL-MIMO channel estimation under complex near-field conditions.
+The deployment of extremely large antenna arrays (ELAAs) in wideband extremely large-scale multiple-input multiple-output (XL-MIMO) systems introduces significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties.
+When combined with the dual-wideband effects inherent to wideband systems, these phenomena fundamentally reshape the channel’s sparsity patterns in the angular-delay domain, rendering existing sparsity-based channel estimation methods inadequate.
+To address these challenges, this paper reconsiders the channel estimation problem for wideband XL-MIMO systems.
+By leveraging the spatial-chirp property of array responses, we quantitatively characterize the angular-delay domain sparsity of wideband XL-MIMO channels, revealing both global block sparsity and local common-delay sparsity.
+To effectively capture this structured sparsity, we propose a novel column-wise hierarchical prior model that incorporates precision parameter sharing mechanism and Markov random field (MRF) structure.
+Based on this prior model, the channel estimation task is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem.
+Tailored to the complex factor graph induced by this hierarchical prior, we develop an MMV-based hybrid message passing (MMV-HMP) algorithm.
+The algorithm performs message updates along the factor graph edges by selectively applying either the sum-product algorithm (SPA) or variational message passing (VMP) rules, depending on the factor-node structure and message tractability.
+Simulation results validate the effectiveness of the proposed column-wise hierarchical prior model through comprehensive ablation studies and demonstrate that the MMV-HMP algorithm consistently outperforms existing baselines, achieving robust estimation performance with moderate computational complexity.
 Index Terms:
@@
 I Introduction
-Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2].
-By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[3].
+Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2, 3].
+By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[4, 5].
 It also delivers substantial beamforming gains, effectively mitigating the severe path loss characteristic of millimeter-wave (mmWave) frequencies.
-Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [4].
+Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [6, 7].
 However, realizing these performance benefits critically depends on obtaining accurate channel state information (CSI), underscoring the importance of precise channel knowledge.
@@
 Considering the inherent spatial-domain sparsity induced by SnS properties, several Bayesian inference-based methods have been developed [20, 21] to mitigate the SnS effects.
-Additionally, exploiting the joint sparsity of XL-MIMO channels in both the spatial and angular domains, methods such as [8, 19] have proposed two-stage and joint VR detection and channel estimation algorithms, considering both spherical wavefront and SnS properties.
+Additionally, exploiting the joint sparsity of XL-MIMO channels in both the spatial and angular domains, methods such as [11, 22] have proposed two-stage and joint visibility region (VR) detection and channel estimation algorithms, considering both spherical wavefront and SnS properties.
 However, despite these advancements, these methods predominantly focus on narrowband systems without the considerations of dual-wideband effects.
@@
 More recently, methods addressing both the spherical wavefront effect and dual-wideband effects have been proposed.
-For example, [25] introduced a bilinear pattern detection (BPD)-based approach to recover wideband XL-MIMO channels, while [26] developed a hybrid message passing algorithm based on constrained Bethe free energy minimization.
+For example, [28] introduced a bilinear pattern detection (BPD)-based approach to recover wideband XL-MIMO channels, while [29] developed a message passing algorithm based on constrained Bethe free energy minimization.
 However, despite these advancements, none of these approaches explicitly incorporate SnS properties, highlighting a significant gap in wideband XL-MIMO channel estimation.
@@
 The main contributions of this paper are summarized as follows:
-  • To accurately characterize the SnS dual-wideband XL-MIMO channels, we derive a novel spatial-frequency domain model based on the measurement findings reported in [11], which comprehensively accounts for the spherical wavefront effect, SnS characteristics, and dual-wideband effects.
-By leveraging the spatial-chirp property of the array response, we quantitatively investigate the impact of the spherical wavefront effect, SnS properties, and dual-wideband effects on the angular and delay spreads.
-Furthermore, we reveal the sparsity patterns of SnS dual-wideband XL-MIMO channels in the angular-delay domain, incorporating both global block sparsity and local common-delay sparsity.
-  • Exploiting the sparsity of the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem.
-Recognizing the critical role of an accurate prior model in Bayesian inference, we propose a novel column-wise hierarchical prior model, where Markov random field (MRF) and common-delay sparsity mechanism is introduced to capture both the global block sparsity and the local common-delay sparsity, ensuring a more accurate and structured representation of the channel characteristics.
-  • To enable efficient Bayesian inference, we propose a MMV-based variational message passing (MMV-VMP) algorithm tailored for the column-wise hierarchical prior model.
-Unlike the variance states propagation algorithm in [27], which relies on first-moment matching methods under the mean-field approximation to update prior model parameters, the MMV-VMP algorithm employs structured variational inference combined with belief propagation.
-This approach effectively handles the higher-order dependencies and loops in the complex factor graph, ensuring precise and reliable message updates.
-  • Compared to existing methods, the proposed MMV-VMP algorithm achieves superior performance by jointly exploiting the global and common-delay sparsity of SnS dual-wideband channels in the angular-delay domain.
-Additionally, as the algorithm avoids matrix inversion operations, it maintains moderate computational complexity, making it well-suited for practical deployment in XL-MIMO systems.
+  • To accurately characterize the SnS dual-wideband XL-MIMO channels, we derive a spatial-frequency domain model that comprehensively captures the spherical wavefront propagation, SnS characteristics, and dual-wideband effects.
+By exploiting the spatial-chirp property of array responses, we quantitatively analyze the influence of spherical wavefront propagation, SnS properties, and dual-wideband effects on the angular and delay spreads.
+Furthermore, we reveal the resulting sparsity patterns in the angular-delay domain, which exhibit both global block sparsity and local common-delay sparsity.
+  • Exploiting the sparsity in the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation problem is formulated as a multiple measurement vector (MMV)-based sparse recovery task.
+Recognizing the importance of an accurate prior model in channel reconstruction, we introduce a novel column-wise hierarchical prior, which incorporates both a precision parameter sharing mechanism and a Markov random field (MRF) structure.
+This design enables the model to effectively capture global block sparsity and local common-delay sparsity, thereby providing a more structured and accurate representation of the underlying channel characteristics.
+  • To achieve efficient Bayesian inference on the intricate factor graph, resulting from the incorporation of both the precision parameter sharing mechanism and the MRF structure, we propose a MMV-based hybrid message passing (MMV-HMP) algorithm that jointly leverages belief propagation (BP) and variational message passing (VMP) rules.
+Thanks to this flexible message update strategy, the MMV-HMP algorithm can effectively manage the message updates for latent precision parameters with non-Gaussian distributions and accommodate the loops inherent in the complex factor graph, thereby enabling accurate and reliable inference.
+  • Compared to existing approaches, the proposed MMV-HMP algorithm achieves superior performance by jointly capturing the global block sparsity and local common-delay sparsity inherent in SnS dual-wideband XL-MIMO channels.
+This enables robust channel estimation while maintaining moderate computational complexity.
+In addition, the effectiveness of the proposed column-wise hierarchical prior model is validated through comprehensive ablation studies.
 I-C Organization and Notations
@@
 Section IV formulates the channel estimation problem as a sparse signal recovery task and proposes a column-wise hierarchical sparse prior model.
-Section V presents the MMV-VMP algorithm.
+Section V presents the MMV-HMP algorithm.
 Simulation results and conclusions are provided in Sections VI and VII, respectively.
@@
 II System Model
-Figure 1: Illustration of the XL-MIMO system model.
-Consider a mmWave XL-MIMO system, as shown in Fig.
-1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration to serve a single-antenna user.
+Consider a mmWave XL-MIMO OFDM system, as shown in Fig.
+1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration.
 To reduce hardware cost and energy consumption, the BS adopts a hybrid analog-digital architecture, where $N_{\mathrm{RF}}$ radio frequency (RF) chains are connected to the ELAA through phase shifters.
 The spacing between adjacent antennas is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$, with $c$ and $f_{c}$ indicating the speed of light and the central carrier frequency, respectively.
-In addition, to address frequency-selective fading, OFDM is utilized.
-Without loss of generality, assume that $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
-In this paper, we focus on the uplink channel estimation, where the BS combines the pilot signal using all RF chains associated with different beams.
+For the uplink channel estimation, we assume the $J$ users transmit mutual orthogonal pilot sequences to the BS [30, 26, 31], e.g., orthogonal time or frequency resources are utilized for different users to transmit pilot sequences.
+Therefore, channel estimation for each user is independent.
+Without loss of generality, we consider an arbitrary user and assume $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
+Figure 1: Illustration of the XL-MIMO system model.
+In the uplink channel estimation phase, the BS combines the pilot signal using all RF chains associated with different beams.
 Let the total number of received beams be denoted by $M$, with $M$ being an integer multiple of $N_{\mathrm{RF}}$.
@@
 Assume that there are $L$ propagation paths between the BS and user.
-Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l>1$ indicates reflection or diffraction path, as illustrated in Fig.
+Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l\geq 2$ indicates reflection or diffraction path, as illustrated in Fig.
 1.
 Denote $\tau_{l,n}$ as the time delay of the $l$-th path for the $n$-th antenna.
-Then the baseband channel impulse response of the $n$-th received antenna is given by
-$h_{n}(t)=\sum_{l=1}^{L}\tilde{\alpha}_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l,n}}\delta(t-\tau_{l,n}),$ (4)
+Then the time-invariant baseband channel impulse response of the $n$-th received antenna can be given by[32]
+$h_{n}(\tau)=\sum_{l=1}^{L}\tilde{\alpha}_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l,n}}\delta(\tau-\tau_{l,n}),$ (4)
 where $\tilde{\alpha}_{l}\in\mathbb{C}$ denotes the complex path gain.
@@
 Then, $\tau_{l,n}$, for any $l\in\mathcal{L}$, can be rewritten as
-$\tau_{l,n}=\tau_{l}+\frac{r_{l}-r_{l,n}}{c},$ (6)
+$\tau_{l,n}=\frac{r_{l,n}}{c}=\tau_{l}-\frac{r_{l}-r_{l,n}}{c},$ (6)
 where $\tau_{l}=r_{l}/c$; $r_{l,n}$ denotes the distance between the $l$-th scatterer and the $n$-th antenna element, which is defined as
@@
 In this Taylor expansion, the current point is set at $n=0$, and terms of third order and higher are neglected.
-Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}+n{\psi_{l}}/{f_{c}}-n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$.
-According to (6) and (7), the continuous time channel impulse response in (4) can be further expressed as
-$h_{n}(t)=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\delta\left(t-\tau_{l,n}\right),$ (8)
+Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}-n{\psi_{l}}/{f_{c}}+n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$.
+According to (6) and (7), the channel impulse response in (4) can be further expressed as
+$h_{n}(\tau)=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\delta\left(\tau-\tau_{l,n}\right),$ (8)
 where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain.
-Applying the continuous time Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (10), as shown in the top of this page.
-$h_{n}(f)=\int_{-\infty}^{+\infty}h_{n}(t)\mathrm{e}^{-\mathrm{j}2\pi ft}\mathrm{d}t=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}\mathrm{e}^{-\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (10)
+Applying the Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (9)[33, 34], as shown in the top of this page.
+$h_{n}(f)=\int_{-\infty}^{+\infty}h_{n}(\tau)\mathrm{e}^{-\mathrm{j}2\pi f\tau}\mathrm{d}\tau=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\mathrm{e}^{\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}.$ (9)
 Further, Eq. (9) can be expressed in a more compact form as
 $\mathbf{h}(f)=\sum_{l=1}^{L}\alpha_{l}a(f,\tau_{l})(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l}))\odot\boldsymbol{\theta}(f,\psi_{l},\varphi_{l}),$ (10)
-where $a(f,\tau_{l})=\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}$; $\boldsymbol{\kappa}_{l}=[\kappa_{l,1},\kappa_{l,2},\cdots,\kappa_{l,N}]^{\mathrm{T}}$ denotes the visibility indicator vector with the $n$-th entry being $\kappa_{l,n}$; $\mathbf{b}(\psi_{l},\varphi_{l})$ and $\mathbf{a}(f,\tau_{l})\}$ denote the array response vector and the frequency response vector, respectively, with each of their $n$-th entries given by
-$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)},$ (12)
-$[\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (13)
-Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K$.
+where $a(f,\tau_{l})=\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}$; $\boldsymbol{\kappa}_{l}=[\kappa_{l,1},\kappa_{l,2},\cdots,\kappa_{l,N}]^{\mathrm{T}}$ denotes the visibility indicator vector with the $n$-th entry being $\kappa_{l,n}$; $\mathbf{b}(\psi_{l},\varphi_{l})$ and $\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})$ denote the array response vector and the frequency-dependent phase-shift vector, respectively, with each of their $n$-th entries given by
+$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)},$ (11)
+$[\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (12)
+Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each pilot subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K-1$.
 The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in Eq.
@@
 $\mathbf{H}=\sum_{l=1}^{L}\alpha_{l}\left(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})\right)\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (13)
-where $\mathbf{a}(\tau_{l})=[a(f_{1},\tau_{l}),\cdots,a(f_{K},\tau_{l})]^{\mathrm{T}}$, and $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})=[\boldsymbol{\theta}(f_{1},\psi_{l},\varphi_{l}),\cdots,\boldsymbol{\theta}(f_{K},\psi_{l},\varphi_{l})]$.
+where $\mathbf{a}(\tau_{l})=[a(f_{1},\tau_{l}),\cdots,a(f_{K},\tau_{l})]^{\mathrm{T}}$, and $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})=[\boldsymbol{\theta}(f_{0},\psi_{l},\varphi_{l}),\cdots,\boldsymbol{\theta}(f_{K-1},\psi_{l},\varphi_{l})]$.
 Remark 1.
@@
 Building on approaches used for traditional massive MIMO channels [9], the wideband XL-MIMO channels in (13) can be approximated as
-$\mathbf{H}\approx\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}.$ (15)
-where $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N\times N}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ denote the normalized $N$- and $K$-dimensions DFT matrices.
+$\mathbf{H}=\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}},$ (14)
+where $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N\times N}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ denote the normalized $N$- and $K$-dimensions discrete Fourier transformation (DFT) matrices.
 $\mathbf{X}\in\mathbb{C}^{N\times K}$ denotes the angular-delay channel.
+From (11) and (12), it can be observed that both the frequency-dependent and frequency-independent array response vectors take the form of spatial chirp signals.
+Consequently, they can be reformulated as
+$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi\left({\psi_{l}}-n{\varphi_{l}}\right)n},$ (15)
+$[\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n},$ (16)
+where $a_{l,k}=\psi_{l}{f_{k}}/{f_{c}}$ and $b_{l,k}=2\varphi_{l}{f_{k}}/{f_{c}}$ denote the initial spatial frequency and the spatial chirp rate, respectively.
+By exploiting the chirp-like properties described in (15), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})$ can be approximated as [35]
+$i_{l,s}=\psi_{l}-2\varphi_{l}{n_{l,s}},\quad i_{l,e}=\psi_{l}-2\varphi_{l}{n_{l,e}}.$ (17)
+In this manner, the array response vector $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})$, which captures both the spherical wavefront effect and the SnS characteristics, can be approximated as a superposition of multiple far-field array response components, i.e.,
+$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\int_{i_{l,e}}^{i_{l,s}}c(\epsilon_{l})\mathbf{b}_{\mathrm{far}}(\epsilon_{l})\mathrm{d}\epsilon_{l},$ (18)
+where $c(\epsilon_{l})$ denotes the decomposition coefficient, and $[\mathbf{b}_{\mathrm{far}}(\epsilon_{l})]_{n}=\mathrm{e}^{\mathrm{j}2\pi\epsilon_{l}n}$ represents the far-field array response corresponding to spatial frequency $\epsilon_{l}\in[i_{l,e},i_{l,s}]$.
+Furthermore, the continuous decomposition in (18) can be discretized with a spatial frequency resolution of $2/N$ as
+$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\sum_{p=1}^{P}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p}),$ (19)
+where the number of sampling points is given by $P=\lceil({n_{l,s}}-{n_{l,e}})\varphi_{l}N\rceil$, and the discretized spatial frequency is defined as $\epsilon_{l,p}=\min(i_{l,s},i_{l,e})+{2(p-1)}/{N}$ for $p=\{1,2,\dots,P\}$.
+The coefficient $c_{l,p}$ denotes the corresponding weight associated with the $p$-th far-field array response component.
+By utilizing the decomposition in (19), the spatial-frequency channel described in (13) can be expressed as
+$\mathbf{H}=\sum_{l=1}^{L}\sum_{p=1}^{P}\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (20)
+Similar to the analysis in (17), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})$ can be approximated as
+$i_{l,s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (21)
+$i_{l,e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d}.$ (22)
+In addition, for notation simplification, we define
+$t_{l,n}=\frac{n^{2}{f_{s}\varphi_{l}}-n{f_{s}\psi_{l}}}{Kf_{c}}.$ (23)
 Lemma 1.
-The angular-delay channel $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{*}$ is a sparse matrix containing only $L$ non-zero blocks, with the non-zeros elements staying inside a square region, i.e.,
-$\lim_{N,K\to\infty}[\mathbf{X}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{2},\\ 0,&(n,k)\notin\mathcal{A}_{2},\end{cases}$ (16)
-where $\mathcal{A}_{2}=\cup_{l,p}\mathcal{A}_{l,2,p}$ with
-$\mathcal{A}_{l,2,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (17)
-$k$ $\left.=\mathrm{mod}(k_{l}+f_{s}\tau_{l},K),\forall(n_{l},k_{l})\in\mathcal{A}_{l,1}\right\},$
-and $\mathrm{mod}(a,m)$ is the modulus of $a$ for $m$.
-In addition, $\epsilon_{l,p}=\psi_{l}-2n_{l,e}\varphi_{l}+\eta_{l}(p-1)$ with $\eta_{l}={2N\varphi_{l}}/{P}$ and $p\in\{p\in\mathbb{N}^{+}\mid\psi_{l}-2n_{l,e}\varphi_{l}\leq\eta_{l}p\leq\psi_{l}-2n_{l,s}\varphi_{l}\}$.
-$\mathcal{A}_{l,1}\triangleq\left\{(n_{l},k_{l})\in\mathbb{Z}^{2}\mid I_{l,e}\leq n_{l}\leq I_{l,s},J_{l,e}\leq k_{l}\leq J_{l,s}\right\}$ with $I_{l,e}=\min_{k}\lceil(i_{l,e}(k)+1)N/2\rceil$, $I_{l,s}=\max_{k}\lceil(i_{l,s}(k)+1)N/2\rceil$, $J_{l,e}=\min_{n}j_{l,e}(n)$, and $J_{l,s}\max_{n}j_{l,s}(n)$, $\forall n\in\boldsymbol{\phi}_{l}$, where
-$i_{l,e}(k)$ $=2(a_{k}-b_{k}n_{l,e}),\quad i_{l,s}(k)=2(a_{k}-b_{k}n_{l,s}),$ (18)
-$j_{l,e}(n)$ $=(t_{l,n}-{2}/K)K,\quad j_{l,s}(n)=(t_{l,n}+{2}/K)K,$ (19)
-with $a_{k}={f_{k}\psi_{l}}/{f_{c}}$, $b_{k}={2f_{k}\varphi_{l}}/{f_{c}}$, and $t_{l,n}=(n{f_{s}\psi_{l}}-n^{2}{f_{s}\varphi_{l}})/(Kf_{c})$.
-From (17), the non-zero region of $\mathbf{X}$ is synthesized from the shifted non-zero regions of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}$ across angular and delay domains for all $l$.
-The shift distance are jointly determined by $N\epsilon_{l,p}$ and $f_{s}\tau_{l}$.
-In addition, The index of significant angular-delay components $(n,k)$ within $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ form a parabolic shape, which arises because both $j_{l,e}(k)$ and $j_{l,s}(k)$ are quadratic functions of $n$.
-Proof: Please see Appendix A.
-Moreover, we analyze the effects of the spherical wavefront, SnS properties, and dual-wideband effects on the sparsity structure of $\mathbf{X}$ based on Lemma 1 from the perspectives of the angular and delay spread.
-1) Angular Spread: The spherical wavefront introduces angular spread of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ in angular domain, ranging from $a_{k}-b_{k}n_{l,e}$ to $a_{k}-b_{k}n_{l,s}$ due to the inclusion of a quadratic term, $\varphi_{l}n^{2}$.
-Meanwhile, it affects the variations of $\epsilon_{l,p}$.
-The SnS properties influence the values of $n_{l,e}$ and $n_{l,s}$, thereby affecting the extent of the angular spread.
-Additionally, the frequency-wideband effect induces variations in $i_{l,e}(k)$ and $i_{l,s}(k)$ across subcarriers $k$, which further broadens the angular dispersion across the system bandwidth.
-2) Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread.
+By utilizing equations (20)-(22), it can be shown that the two-dimensional inverse discrete Fourier transform (2D-IDFT) of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ results in a sparse matrix, where the nonzero entries are confined within a square region.
+Within this region, the dominant components exhibit a clear column-wise clustered sparsity structure.
+Specifically, as $N,K\rightarrow\infty$, the support of the nonzero region converges to:
+$[\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{l},\\ 0,&(n,k)\notin\mathcal{A}_{l},\end{cases}$ (24)
+where the set $\mathcal{A}_{l}$ is defined as
+$\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\},$ (25)
+with
+$I_{l,e}$ $=\min_{k}\lceil(i_{l,e}(k)+1)N/2\rceil,$ (26)
+$I_{l,s}$ $=\max_{k}\lceil(i_{l,s}(k)+1)N/2\rceil,$
+$J_{l,e}$ $=\max_{n}(t_{l,n}+{2}/{K}+1/2)K,$
+$J_{l,s}$ $=\min_{n}(t_{l,n}-{2}/{K}+1/2)K,$
+indicating the boundaries.
+In addition, within the region $\mathcal{A}_{l}$, each column corresponds to the angular response associated with a specific propagation delay.
+The width of the significant spatial frequency components is primarily determined by the frequency-width effect.
+Assuming the starting spatial frequency of the $k$-th column is denoted by $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$, the corresponding angular spread can be approximated as
+$B_{l,k}=\frac{i_{l,k}f_{s}}{f_{c}}N.$ (27)
+As a result, the significant components of $\boldsymbol{\Xi}_{l}$ within $\mathcal{A}_{l}$ are confined to narrow bands along each column, thereby exhibiting a pronounced column-wise clustered sparsity pattern.
+To summarize, by leveraging (24)-(27), the sparse structure of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ can be accurately characterized.
+In particular, equation (24) delineates the overall boundaries of the nonzero region, while (27) provides a more precise description of the angular spread in each column, further refining the support of the dominant components.
+Proof.
+Please see Appendix A. ∎
+Theorem 1.
+Based on Lemma 1, the angular-delay domain channel matrix $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{*}$ exhibits a sparse structure, containing only $L$ nonzero blocks, each corresponding to an individual propagation path.
+For the $l$-th path, the support region of the nonzero components is denoted by $\mathcal{B}_{l}$, which can be expressed as $\mathcal{B}_{l}=\bigcup_{p}\mathcal{B}_{l,p}$, where $\mathcal{B}_{l,p}$ represents the shifted version of $\mathcal{A}_{l}$ in both the angular and delay domains, and is defined as
+$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (28)
+$k$ $\left.=\mathrm{mod}(k_{l}+f_{s}\tau_{l},K),\forall(n_{l},k_{l})\in\mathcal{A}_{l}\right\},$
+where $\mathrm{mod}(a,m)$ is the modulus of $a$ for $m$.
+According to Theorem 1, the nonzero region of $\mathbf{X}$ is formed by aggregating the shifted nonzero regions of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ across both the angular and delay domains for all paths $l$.
+The shift positions are jointly determined by the spatial shift $N\epsilon_{l,p}$ and the delay shift $f_{s}\tau_{l}$, corresponding to the spatial frequency and propagation delay of each path, respectively.
+Proof.
+Please see Appendix B. ∎
+Moreover, we analyze the effects of the spherical wavefront, SnS properties, and dual-wideband effects on the sparsity structure of $\mathbf{X}$ based on Lemma 1 and Theorem 1 from the perspectives of the angular and delay spread.
+  • Angular Spread: First, the spherical wavefront effect and the SnS characteristics determine the angular spread range at each subcarrier, characterized by $i_{l,s}(k)$ and $i_{l,e}(k)$.
+Then, the frequency-wideband effect governs the overall boundaries of the angular spread, denoted by $I_{l,e}$ and $I_{l,s}$.
+Moreover, the frequency-wideband effect also enables a more accurate description of the angular spread within each column, represented by $B_{l,k}$.
+  • Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread.
 The SnS properties define the range of variation for $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
-In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively influence the angular and delay spreads, reshaping the sparse patterns of SnS dual-wideband channels in the angular-delay domain, i.e.,
-  • Global Block Sparsity: The angular-delay channel contains $L$ non-zero block corresponding to $L$ propagation paths. The index range of each non-zero block is determined by $\mathcal{A}_{l,2}$.
-  • Local Common-Delay Sparsity: Due to the spherical wavefront effect, within $\mathcal{A}_{2}$, different delay indices correspond to varying angular spreads, which occur in clusters and exhibit high correlation.
+In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively govern the angular and delay spreads, thereby reshaping the sparsity structure of SnS dual-wideband channels in the angular-delay domain.
+Specifically,
+  • Global Block Sparsity: The angular-delay domain channel exhibits $L$ nonzero blocks, each corresponding to a distinct propagation path.
+The index range of each nonzero block is characterized by $\mathcal{B}_{l}$.
+  • Local Common-Delay Sparsity: Due to the frequency-wideband effect, the significant components within $\mathcal{B}_{l}$ are confined to narrow bands along each column, thereby exhibiting pronounced column-wise clustered structures with strong correlation.
 Remark 2.
@@
 Once the angular-delay channel is reconstructed, key parameters such as AoAs, VRs, and the range of beam squint can be efficiently extracted.
+(a) Angular spread for $k=0$ and $K-1$
+(b) Delay spread for $n=n_{s}$ and $n_{e}$
+(c) $\mathcal{A}_{l}$.
+(d) $\mathcal{B}_{l,1}$.
+(e) $\mathcal{B}_{l,P}$
+(f) $\mathcal{B}_{l}$.
+Figure 2: Illustration of block sparsity of SnS XL-MIMO channel in the angular-delay domain.
 To enhance intuitive understanding of the SnS XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$.
 Fig.
-2 illustrates the block sparsity of the SnS XL-MIMO channel in the angular-delay domain.
-Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,e}(64)=513$, and $j_{l,s}(64)\approx 517$.
-Similarly, $i_{l,e}(1023)\approx 0.88$, $i_{l,s}(1023)\approx 0.72$, $j_{l,e}(959)=545$, and $j_{l,s}(959)\approx 549$.
+2 illustrates the sparsity pattern of the SnS XL-MIMO channel in the angular-delay domain.
+Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,s}(n_{s})=513$, and $j_{l,e}(n_{s})\approx 517$.
+Similarly, $i_{l,s}(K-1)\approx 0.88$, $i_{l,e}(K-1)\approx 0.72$, $j_{l,s}(n_{e})=545$, and $j_{l,e}(n_{e})\approx 549$.
 As depicted in Fig.
@@
 Additionally, the figures show that both the angular and delay spectra shift to the right due to $\cos\vartheta_{l}>0$ and the monotonicity of $t_{n}$ with respect to $n$.
-Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, and $J_{l,s}=549$.
-Fig.
-2(c) illustrates the non-zero region of $\mathbf{F}^{\mathrm{A}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l,1}$, validating the effectiveness of Lemma 1.
+Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, $J_{l,s}=549$, and $B_{k,s}\approx 41$ with $k=J_{l,e}$ Fig.
+2(c) illustrates the non-zero region of $\mathbf{F}_{\mathrm{A}}^{{}_{\mathrm{H}}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l}$, validating the effectiveness of Lemma 1.
+In particular, for the $J_{l,e}$-th column, the width of angular spread also aligns the result of (27), which further validates the influence of frequency-wideband effect for local angular spread.
 Additionally, we have $f_{s}\tau_{l}=100$, $\epsilon_{l,1}=\psi_{l}-2n_{l,s}\varphi_{l}\approx 0.4$, and $\epsilon_{l,P}=\psi_{l}-2n_{l,e}\varphi_{l}=0.32$, corresponding to $N\epsilon_{l,1}=410$ and $N\epsilon_{l,P}=328$, respectively.
@@
 2(f) is obtained.
-It can be observed that the angular range of $\mathcal{A}_{l,2}$ is broader than that of $\mathcal{A}_{l,1}$ due to the composite effect.
-(a) Angular spread for $k=0$ and $K-1$.
-(b) Delay spread for $n=0$ and $N_{\mathrm{R}}-1$.
-(c) $\mathcal{A}_{l,1}$.
-(d) $\mathcal{A}_{l,p,1}$ with $p=1$.
-(e) $\mathcal{A}_{l,p,1}$ with $p=P$
-(f) $\mathcal{A}_{l,2}$.
-Figure 2: Illustration of block sparsity of SnS XL-MIMO channel in the angular-delay domain.
 Remark 3.
@@
 When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted.
-Thus, (18) simplifies to $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$.
+Thus, we have $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$.
 Consequently, the significant components are uniformly distributed within the non-zero region, as illustrated in [9].
-IV SnS Dual-Wideband Channel Estimation for XL-MIMO Systems
-In this section, leveraging the sparsity outlined in Lemma 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task.
+IV Problem Formulation and Sparse Prior Modeling
+In this section, leveraging the sparsity outlined in Theorem 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task.
 Subsequently, to account for the global block sparsity and local common-delay sparsity, we introduce a novel column-wise sparse prior model tailored to this structure.
 IV-A Problem Formulation
-Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a multiple measurement vectors (MMV) based sparse recovery problem.
+Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a MMV-based sparse recovery problem.
 Utilizing the angular-delay representation in (14), the received signal model in (3) can be further written as
-$\mathbf{Y}=\mathbf{W}\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}+{\mathbf{N}}=\boldsymbol{\Psi}\mathbf{X}\mathbf{F}_{\mathrm{D}}+{\mathbf{N}},$ (20)
+$\mathbf{Y}=\mathbf{W}\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}}+{\mathbf{N}}=\boldsymbol{\Psi}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}}+{\mathbf{N}},$ (29)
 where $\boldsymbol{\Psi}\triangleq\mathbf{W}\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{M\times N}$. Furthermore, utilizing the unitary property of $\mathbf{F}_{\mathrm{D}}$, (29) can be reformulated as
-$\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}=\boldsymbol{\Psi}\mathbf{X}+\tilde{\mathbf{N}},$ (21)
-where $\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\in\mathbb{C}^{M\times K}$ and $\tilde{\mathbf{N}}={\mathbf{N}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\in\mathbb{C}^{M\times K}$ denotes the equivalent received pilot signal and noise matrix.
+$\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{*}=\boldsymbol{\Psi}\mathbf{X}+\tilde{\mathbf{N}},$ (30)
+where $\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{*}\in\mathbb{C}^{M\times K}$ and $\tilde{\mathbf{N}}={\mathbf{N}}\mathbf{F}_{\mathrm{D}}^{*}\in\mathbb{C}^{M\times K}$ denotes the equivalent received pilot signal and noise matrix.
+Remark 4.
+While this work primarily focuses on the ULA configuration, the signal models in (29) and (30) are also applicable to the uniform planar array (UPA) configuration.
+Specifically, consider a BS equipped with a half-wavelength UPA comprising $N=N_{v}\times N_{h}$ antennas, where $N_{v}$ and $N_{h}$ denote the number of elements along the vertical and horizontal dimensions, respectively.
+In this case, the angular transformation matrix $\mathbf{F}_{\mathrm{A}}$ can be redefined as the Kronecker product $\mathbf{F}_{\mathrm{A}}=\mathbf{F}_{\mathrm{A},v}\otimes\mathbf{F}_{\mathrm{A},h}$, where $\mathbf{F}_{\mathrm{A},v}\in\mathbb{C}^{N_{v}\times N_{v}}$ and $\mathbf{F}_{\mathrm{A},h}\in\mathbb{C}^{N_{h}\times N_{h}}$ denote the vertical and horizontal DFT matrices, respectively.
+It is worth noting that the sparsity structure in the angular-delay domain may differ between ULA and UPA configurations, particularly with regard to local common-delay sparsity, due to the inherent differences in array geometry and two-dimensional angular resolution.
+Although a comprehensive analysis of the UPA case is beyond the scope of this paper, we acknowledge its importance and leave it as a promising direction for future work.
 In this paper, our primary objective is to develop an effective estimation algorithm for accurately reconstructing $\mathbf{X}$ or $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$.
-In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (21), including orthogonal matching pursuit (OMP) [28, 29] and optimization-based methods [30, 31].
+In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (30), including SOMP [30, 36] and optimization-based methods [37, 38].
 However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional parameters, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available.
-Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and sparse Bayesian learning (SBL) have shown superior recovery performance in channel estimation tasks [32, 33, 27, 34].
+Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and SBL have shown superior recovery performance in channel estimation tasks [39, 40, 41].
 However, the effectiveness of these Bayesian methods relies heavily on the accuracy of the prior model.
-To address this, we propose a novel column-wise hierarchical sparse prior model, as detailed in the next section.
 IV-B Column-Wise Hierarchical Sparse Prior Model
-To capture the global and local common-delay sparsity characteristics of $\mathbf{X}$, the column-wise hierarchical prior model is characterized as
-$p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})=p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S}),$ (22)
-where the first layer is modeled as a complex Gaussian distribution with mean zero and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
-$p(\mathbf{X}|\boldsymbol{\Gamma})=\prod_{n=1}^{N}\prod_{k=1}^{K}\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1}).$ (23)
-In the second layer, each $\gamma_{n,k}$ is assigned a conditionally Bernoulli distribution given by
+According to Lemma 1 and Theorem 1, the angular-delay domain channel exhibits prominent block sparsity and a column-wise clustered structure.
+However, existing sparse prior models fail to capture these characteristics fully [40, 39, 42, 41].
+For example, the two-layer hierarchical prior models in [40, 39] assign independent priors to the variance precision parameters of the channel coefficients, which limits their ability to model the correlation across coefficients.
+Although [42, 41] leverage MRF to promote global block sparsity, they still overlook local structural dependencies, especially those induced by column-wise clustering.
+As a result, these models are not well suited for the SnS wideband XL-MIMO channel estimation problem considered in this work.
+To this end, we propose a novel column-wise hierarchical sparse prior model, which incorporates both precision parameter sharing mechanism and MRF structure to promote structured sparsity.
+Specifically, the column-wise hierarchical prior is defined as
+$p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})=p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S}).$ (31)
+We next detail the distributions associated with each layer.
+1) Variance-Driven Sparsity: The first layer adopts a complex Gaussian prior with zero mean and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
+$p(\mathbf{X}|\boldsymbol{\Gamma})=\prod_{n=1}^{N}\prod_{k=1}^{K}\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1}),$ (32)
+where $\gamma_{n,k}$ denotes the variance precision parameter, with $\gamma_{n,k}^{-1}$ representing the variance of $x_{n,k}$.
+This precision-based modeling naturally induces sparsity: as the variance $\gamma_{n,k}^{-1}$ approaches zero, the corresponding $x_{n,k}$ is effectively pushed toward zero, thereby promoting a sparse solution.
+2) Precision Parameters Sharing Mechanisms: Due to the finite angular and delay resolutions, the angular-delay domain channel does not exhibit strict sparsity, as the coefficients near the boundaries of the sparse patterns often retain non-negligible energy.
+Enforcing hard zeros on these components may lead to degraded estimation performance.
+To mitigate this issue, in the second layer, we adopt a dual-precision parameter strategy, wherein each variance $\gamma_{n,k}$ is associated with two distinct precision parameters corresponding to its active and inactive states.
+In addition, to capture the local common-delay clustered sparsity, we further introduce the precision parameters sharing mechanism.
+Specifically, $\gamma_{n,k}$ is assigned a conditionally Bernoulli-Gaussian distribution given by
 $p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})$ $=\prod_{n=1}^{N}\prod_{k=1}^{K}\delta\left(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k})\right),$ (33)
@@
 The precision parameter $\boldsymbol{\Omega}=[\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2},\cdots,\boldsymbol{\alpha}_{K}]^{\mathrm{T}}\in\mathbb{R}^{2\times K}$ of $\boldsymbol{\Gamma}$ is characterized by
-$p(\boldsymbol{\Omega})=\prod_{i=1}^{2}\prod_{k=1}^{K}p(\alpha_{k}^{i})=\prod_{i=1}^{2}\prod_{k=1}^{K}\mathrm{Ga}(\alpha_{k}^{i};a_{i},b_{i}),$ (25)
-In the third layer, to capture the global block sparsity, the variance state variable $\mathbf{S}$ is modeled as a MRF
+$p(\boldsymbol{\Omega})=\prod_{i=1}^{2}\prod_{k=1}^{K}p(\alpha_{k}^{i})=\prod_{i=1}^{2}\prod_{k=1}^{K}\mathrm{Ga}(\alpha_{k}^{i};a_{i},b_{i}).$ (34)
+As described in (33) and (34), this common-delay sparsity mechanism ensures that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter.
+This coupling enforces a consistent sparsity structure among components associated with the same delay.
+3) MRF-Based Variance State Modeling: In the third layer, to capture the global block sparsity, the variance state variable $\mathbf{S}$ is modeled as a MRF
 $p(\mathbf{S})=\left(\prod_{n,k}\prod_{(n^{\prime},k^{\prime})\in\mathcal{D}_{n,k}}u(s_{n,k},s_{n^{\prime},k^{\prime}})\right)^{\frac{1}{2}}\prod_{n,k}v(s_{n,k}),$ (35)
-where $\mathcal{D}_{n,k}$ denotes the element set of neighbor nodes of $s_{n,k}$ with $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$, and $v(s_{n,k})=\exp(-\eta s_{n,k})$; $\varpi$ and $\eta$ are the parameters of $p(\mathbf{S})$.
-In the following, we discuss how prior models in (22)-(26) capture the structured sparsity in $\mathbf{X}$.
-1) Variance-Driven Sparsity: The variance $\gamma_{n,k}^{-1}$ controls the sparsity of $x_{n,k}$, with values approaching zero effectively setting $x_{n,k}$ to zero.
-2) Global Block Sparsity: The MRF models the variance states, encouraging block patterns while suppressing isolated coefficients that deviate from their neighbors.
-3) Local Common-Delay Sparsity: As described in (24) and (25), the common-delay sparsity mechanism is introduced to ensure that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter.
-This enforces a unified sparsity structure among components associated with the same delay.
-IV-C Bayesian Inference
+where $\mathcal{D}{n,k}$ denotes the set of neighboring nodes of $s_{n,k}$; $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$ and $v(s_{n,k})=\exp(-\eta s_{n,k})$ represent the pairwise potential and self-potential functions, respectively, with $\varpi$ and $\eta$ being the model parameters associated with $p(\mathbf{S})$.
+By leveraging the formulation in (35), the prior encourages block patterns in the support structure and suppresses isolated coefficients that deviate from their neighboring states.
+V Proposed MMV-HMP Algorithm
+Based on the proposed column-wise hierarchical sparse prior model, this section first formulates the sparse signal recovery problem as a MMV-based Bayesian inference task.
+Then, to effectively perform the inference, we propose an MMV-HMP algorithm.
+V-A Bayesian Inference
 Since the measurement matrix $\boldsymbol{\Psi}$ may be “bad” (e.g., rank-deficient, ill-conditioned, or having a non-zero mean) [39], the divergence issues might be arisen in the Bayesian inference.
@@
 where $\mathbf{R}=\mathbf{U}^{\mathrm{H}}\tilde{\mathbf{Y}}$, $\mathbf{A}=\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$, $\mathbf{Z}=\mathbf{A}\mathbf{X}$, and $\boldsymbol{\Xi}=\mathbf{U}^{\mathrm{H}}{\tilde{\mathbf{N}}}$.
-Based on the prior model provided in (22), the minimum mean square error (MMSE) estimator for the $(n,k)$-th entry of $\mathbf{X}$ can be expressed as
-$\hat{x}_{n,k}=\int{x}_{n,k}p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\mathrm{d}\beta\mathrm{d}\mathbf{Z}\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\mathrm{d}\mathbf{\boldsymbol{\Omega}}\mathrm{d}\mathbf{S}\mathrm{d}\mathbf{X},$ (28)
-where $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ denotes the joint posterior probability of $\beta$, $\mathbf{Z}$, $\mathbf{X}$, $\boldsymbol{\Gamma}$, $\boldsymbol{\Omega}$, and $\mathbf{S}$, satisfying
+Based on the prior model provided in (31), the maximum a posterior (MAP) estimator for the $(n,k)$-th entry of $\mathbf{X}$ can be expressed as
+$\hat{x}_{n,k}=\int{x}_{n,k}p(x_{n,k}|\mathbf{R})\mathrm{d}x_{n,k},$ (37)
+where the posterior distribution $p(x_{n,k}|\mathbf{R})$ is given by
+$\int p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\,\mathrm{d}\beta\,\mathrm{d}\mathbf{Z}\,\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\,\mathrm{d}\mathbf{\boldsymbol{\Omega}}\,\mathrm{d}\mathbf{S}\,\mathrm{d}\mathbf{X}_{\backslash n,k}$ (38)
+with $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ indicating the joint posterior probability and satisfying
 $p(\beta,\mathbf{X},\mathbf{Z},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\propto p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S},\mathbf{R})$ (39)
@@
 where $p(r_{m,k}|z_{m,k})=\mathcal{CN}(z_{m,k};r_{m,k},\beta^{-1})$; ${r}_{m,k}$ and ${z}_{m,k}$ denote the $(m,k)$-th elements of $\mathbf{R}$ and $\mathbf{Z}$.
-The dependencies of the random variables in the factorization (29) can be shown by a factor graph as depicted in Fig.
-3, where the involved probability distributions are listed in Table I.
 Due to the large number of antennas and subcarriers in XL-MIMO systems, solving problem (37) requires evaluating high-dimensional integrals, which is computationally prohibitive.
 Moreover, traditional SBL-based methods are infeasible in this context as they involve high-dimensional matrix inversions.
-In addition, the introduction of the common-delay sparsity mechanism and the MRF results in a more complex factor graph representation, making existing message passing algorithms [32, 35, 33, 27] less suitable.
-Specifically, [32, 35, 33] do not incorporate the common-delay sparsity mechanism or MRF-based variance state modeling, leading to a mismatch between these algorithms and the factor graph representation in Fig.
-3.
-Additionally, [27] lacks the common-delay sparsity mechanism and uses a first-moment matching method to update the prior parameters, which neglects higher-order dependencies among variable nodes.
-To address these limitations, we propose a novel MMV-VMP algorithm utilizing the variational message passing and belief propagation in the next section.
+Consequently, this highlights the pressing need for novel channel estimation techniques capable of addressing the challenges posed by the SnS dual-wideband channel estimation in XL-MIMO systems.
+Recently, message passing-based techniques have been widely adopted for solving MAP estimation problems, owing to their computational efficiency.
+Motivated by these advances, we propose a MMV-HMP algorithm to efficiently solve problem (37).
+V-B Factor Graph Representation
+The dependencies among the random variables in the factorization (39) are illustrated using a factor graph, as shown in Fig.
+3, with the corresponding probability distributions summarized in Table I.
+It is evident that the factor graph in Fig.
+3 is significantly more intricate than those considered in [39, 43, 40, 42, 41], primarily due to the incorporation of precision parameter sharing mechanism and MRF structure.
+Specifically, the works in [39, 43, 40] adopt independent priors on the variances of coefficients, thereby promoting element-wise sparsity.
+However, such models fail to capture the angular or delay-domain correlations among the coefficients.
+Furthermore, the model in [42] does not include a precision parameter sharing mechanism, and thus overlooks higher-order dependencies among variable nodes.
+Due to these structural distinctions, existing message passing algorithms developed in [39, 43, 40, 42, 41] are not directly applicable to our setting.
+Consequently, the message update equations at the variable nodes must be carefully reformulated to accommodate the enhanced complexity of our factor graph.
 Figure 3: Factor graph representation for the factorization (39).
@@
 $f_{\gamma_{n,k}}$ $p(\gamma_{n,k}|\boldsymbol{\alpha}_{k},s_{n,k})$ $\delta(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k}))$
-V Proposed MMV-VMP Algorithm
-The MMV-VMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (28).
+Moreover, in the proposed sparse prior model, the latent precision parameters $\beta$ and $\alpha^{i}_{k}$ follow non-Gaussian distributions.
+Meanwhile, the shared precision parameter simultaneously influences the message updates in both the first and third layers, resulting in inter-layer dependencies, which significantly increase the complexity of inference and render the exact computation of the node beliefs computationally intractable under a pure belief propagation (BP) framework.
+To overcome these challenges, we propose a novel MMV-HMP algorithm jointly leveraging VMP and sum-produce algorithm (SPA) rules.
+Specifically, tailored to the intractable structure of the factor graph, we reformulate the message update equations along the edges, selectively applying sum-product or variational message passing rules to different edges to enable more flexible and efficient message updates.
+The MMV-HMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (37).
 In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules as shown in Fig.
 3.
-V-A Measurement Module
+V-C Measurement Module
 Measurement module aims to obtain the likelihood estimation of $x_{n,k}$ based on the observation $\mathbf{R}$.
 Denote the belief of $z_{m,k}$ as $b(z_{m,k})\sim\mathcal{CN}(z_{m,k};\hat{z}_{m,k},\hat{\nu}^{z}_{m,k})$, which is defined in (48) with $\hat{z}_{m,k}$ and $\hat{\nu}^{z}_{m,k}$ indicating the $m$-th elements of $\hat{\mathbf{z}}_{k}$ and $\hat{\boldsymbol{\nu}}_{\mathbf{z}_{k}}$, respectively.
-In this manner, according to the variation message passing rule, the message from $f_{r_{m,k}}$ to $\beta$ is given by
+In this manner, according to the VMP rule [44, 45], the message from $f_{r_{m,k}}$ to $\beta$ is given by
 $m_{f_{r_{m,k}}\rightarrow\beta}$ $\propto\exp\left\{\int\ln f_{r_{m,k}}b(z_{m,k})\mathrm{d}z_{m,k}\right\},$ (42)
@@
 $\hat{\beta}=\frac{MK}{\sum_{m=1}\sum_{k=1}\left(\left|r_{m,k}-\hat{z}_{m,k}\right|^{2}+\hat{\nu}^{z}_{m,k}\right)}.$ (44)
-Similar to (32), the forward message passing from $f_{r_{m,k}}$ to $z_{m,k}$ can be given by
+Similar to (42), utilizing VMP rule, the forward message passing from $f_{r_{m,k}}$ to $z_{m,k}$ can be given by
 $m_{f_{r_{m,k}}\rightarrow z_{m,k}}\propto$ $\exp\left\{\int b(\beta)\ln f_{r_{m,k}}\mathrm{d}\beta\right\}$ (45)
@@
 $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}=\mathbf{1}_{M}./(\mathbf{1}_{M}\hat{\beta}^{-1}+\boldsymbol{\nu}_{\mathbf{p}_{k}}),\boldsymbol{\mu}_{k}=\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}\odot(\mathbf{r}_{k}-\mathbf{p}_{k}).$ (47)
-In this case, utilizing the sum-product rule, the belief of $\mathbf{z}_{k}$ is given by
-$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})$ (38)
+In this case, utilizing SPA rule, the belief of $\mathbf{z}_{k}$ is given by
+$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}$ (48)
 $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\mathcal{CN}(\mathbf{z}_{k};\mathbf{r}_{k},\hat{\beta}^{-1}\mathbf{I}_{M\times M})$
@@
 $\boldsymbol{\nu}_{\mathbf{q}_{k}}=\mathbf{1}./(\left|\mathbf{A}\right|^{2}\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}),\quad\mathbf{q}_{k}=\hat{\mathbf{x}}_{k}+\boldsymbol{\nu}_{\mathbf{q}_{k}}\odot(\mathbf{A}^{\mathrm{H}}\boldsymbol{\mu}_{k}).$ (51)
-V-B Sparse Prior Module
+V-D Sparse Prior Module
 The aim of the prior module is to update the distribution of ${x}_{n,k}$ according to extrinsic message $m_{x_{n,k}\rightarrow f_{x_{n,k}}}(x_{n,k})\propto\mathcal{CN}({x}_{n,k};{q}_{n,k},{\nu}^{q}_{n,k})$ and calculate the posterior estimation $\hat{{x}}_{n,k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$, where ${q}_{n,k}$ and ${\nu}^{q}_{n,k}$ denotes the $n$-th element of $\mathbf{q}_{k}$ and $\boldsymbol{\nu}_{\mathbf{q}_{k}}$, respectively.
-V-B1 Forward Message Passing
+V-D1 Forward Message Passing
 Denote the belief of $x_{n,k}$ as $b(x_{n,k};\hat{x}_{n,k},\hat{\nu}_{\mathbf{x}_{k}})$, defined in (63).
-Utilizing the variational message passing rule, similar to (32), the message from $f_{x_{n,k}}$ to $\gamma_{n,k}$ is given by
+Utilizing VMP rule, similar to (42), the message from $f_{x_{n,k}}$ to $\gamma_{n,k}$ is given by
 $m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}\propto$ $\exp\left\{\int{b}(x_{n,k})\ln f_{x_{n,k}}\mathrm{d}x_{n,k}\right\}$ (52)
 $\propto$ $\gamma_{n,k}\exp(-\gamma_{n,k}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})).$
-Further, according to sum-product rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
-$m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}\overset{(a)}{\propto}\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}b(\boldsymbol{\alpha}_{k})\mathrm{d}\gamma_{n,k}\mathrm{d}\boldsymbol{\alpha}_{k}$ (43)
+Furthermore, according to SPA rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
+$m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})\overset{(a)}{\propto}\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}b(\boldsymbol{\alpha}_{k})\mathrm{d}\gamma_{n,k}\mathrm{d}\boldsymbol{\alpha}_{k}$ (53)
 $=\pi^{\mathrm{out}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{out}}_{n,k})\delta(1+s_{n,k}),$
@@
 The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions.
-Take $m_{n,k}^{\mathrm{l}}$ as an example, according to the sum-product rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (45), as shown in the top of next page.
+Take $m_{n,k}^{\mathrm{l}}$ as an example, according to SPA rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (54), as shown in the top of this page.
 $\lambda_{n,j}^{\mathrm{l}}=\frac{\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}}\mathrm{e}^{\varpi-\eta}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}\lambda^{w}_{s_{n,j}^{{\text{l}}}}+(1-\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}})\mathrm{e}^{\eta-\varpi}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}(1-\lambda^{w}_{s_{n,j}^{{\text{l}}}})}{(\mathrm{e}^{\varpi}+\mathrm{e}^{-\varpi})\left(\mathrm{e}^{-\eta}\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}\lambda^{w}_{s_{n,j}^{{\text{l}}}}+\mathrm{e}^{\eta}(1-\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}})\prod_{w\in\left\{\mathrm{l,t,b}\right\}}(1-\lambda^{w}_{s_{n,j}^{{\text{l}}}})\right)}.$ (54)
-The other three messages can be obtained in a similar way.
-V-B2 Backward Message Passing
-With the messages of neighbors and $v(s_{n,k})$, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
+The other messages can be obtained in a similar way.
+V-D2 Backward Message Passing
+With the messages of neighbors and $v(s_{n,k})$ and SPA rule, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
 $m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}=\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}m_{n,k}^{\mathrm{l}}v(s_{n,k})$ (55)
-$\propto\pi^{\mathrm{in}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{in}}_{n,k})\delta(1+s_{n,k})$
+$\propto$ $\pi^{\mathrm{in}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{in}}_{n,k})\delta(1+s_{n,k}),$
 where $\pi^{\mathrm{in}}_{n,k}$ is defined as
 $\frac{\mathrm{e}^{-\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}\lambda_{n,k}^{w}}{\mathrm{e}^{-\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}\lambda_{n,k}^{w}+\mathrm{e}^{\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}(1-\lambda_{n,k}^{w})}.$ (56)
-According to sum-product rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
+According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
 $m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto$ $\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}\mathrm{d}\gamma_{n,k}\mathrm{d}s_{n,k},$
@@
 As a result, the belief of $\alpha_{k}^{1}$ is given by
-${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (49)
+${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (58)
 where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$.
 Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$.
-According to the sum-product rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
+According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
 $m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}=\pi^{\mathrm{in}}_{n,j}{b}(\alpha^{1}_{k})+(1-\pi^{\mathrm{in}}_{n,j}){b}(\alpha^{2}_{k}).$ (59)
@@
 $+$ $(1-\pi^{\mathrm{in}}_{n,k})\gamma_{n,k}^{\hat{a}_{k}^{2}}\exp(-\gamma_{n,k}(\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})).$
-As a result, the message from $f_{x_{n,k}}$ to $x_{n,k}$ is denoted as
-$m_{f_{x_{n,k}}\rightarrow x_{n,k}}(x_{n,k})$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (52)
+As a result, utilizing VMP rule, the message from $f_{x_{n,k}}$ to $x_{n,k}$ is denoted as
+$m_{f_{x_{n,k}}\rightarrow x_{n,k}}$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (61)
 $\propto\exp\left\{-\int\left|x_{n,k}\right|^{2}\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}\right\}$
@@
 where $(a)$ is obtained by
-$\hat{\gamma}_{n,j}$ $=\int\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}=\pi^{\mathrm{in}}_{n,k}\frac{\hat{a}_{k}^{1}+1}{\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}$ (53)
+$\hat{\gamma}_{n,k}$ $=\int\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}=\pi^{\mathrm{in}}_{n,k}\frac{\hat{a}_{k}^{1}+1}{\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}$ (62)
 $+(1-\pi^{\mathrm{in}}_{n,k})\frac{\hat{a}_{k}^{2}+1}{\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}.$
@@
 $\hat{\nu}_{\mathbf{x}_{k}}=\frac{1}{N}\sum_{n=1}^{N}\hat{\nu}_{n,k}^{x}.$ (65)
-Algorithm 1 Proposed UAMP-SBL-MRF algorithm
+Algorithm 1 Proposed MMV-HMP algorithm
 1:received vector $\mathbf{R}$, measurement matrix $\mathbf{A}$.
-2:Initialize: $\hat{\nu}_{\mathbf{x}_{k}}^{(0)}=1$, $\hat{\mathbf{x}}=\mathbf{0}$, $\hat{{\gamma}}_{n,k}={1},\hat{\beta}=1$, and $\boldsymbol{\mu}=\mathbf{0}$.
+2:Initialize: $\hat{\nu}_{\mathbf{x}_{k}}^{(0)}=1$, $\hat{\mathbf{x}}_{k}=\mathbf{0}$, $\hat{{\gamma}}_{n,k}={1},\hat{\beta}=1$, and $\boldsymbol{\mu}_{k}=\mathbf{0}$.
 3:while the stopping criterion is not met do
@@
 10:/*Prior Module*/
-11:     Update the messages $m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}$ according to (43);
-12:     Update the messages $m_{n}^{\mathrm{l}}$, $m_{n}^{\mathrm{r}}$, $m_{n}^{\mathrm{t}}$, and $m_{n}^{\mathrm{b}}$;
-13:     Update the messages $m_{s_{n}\rightarrow f_{\gamma_{n}}}(s_{n})$ according to (47);
+11:     Update the messages $m_{f_{x_{n,k}}\rightarrow s_{n,k}}$ according to (53);
+12:     Update the messages $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$;
+13:     Update the messages $m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}$ according to (56);
 14:     Update the belief $b(\alpha_{k}^{1})$ and $b(\alpha_{k}^{2})$ according to (58);
-15:     Update $\hat{\gamma}_{n}$ according to (53);
+15:     Update $\hat{\gamma}_{n,k}$ according to (62);
 16:     Update $\hat{x}_{n,k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$ according to (64) and (65).
@@
 18:$\hat{x}_{n,k}$.
-The proposed MMV-VMP algorithm is summarized in Algorithm 1 and it can be terminated when it reached a maximum number of iteration or the difference between the estimates of two consecutive iterations is less than $10^{-5}$.
-In the following, we provide the computational complexity analysis for the proposed MMV-VMP algorithm.
+The proposed MMV-HMP algorithm is summarized in Algorithm 1 and it can be terminated when it reached a maximum number of iteration or the difference between the estimates of two consecutive iterations is less than $10^{-5}$.
+In the following, we provide the computational complexity analysis for the proposed MMV-HMP algorithm.
 Examining the steps of Algorithm 1, it is evident that there is no matrix inversion involved.
 Thus, the most computationally intensive parts only involve matrix-vector products in lines 2 and 6, i.e., $\mathcal{O}(MN)$ per iteration.
-Consequently, the total complexity of the MMV-VMP algorithm is $\mathcal{O}(TKMN)$, where $T$ denotes the number of iterations.
+Consequently, the total complexity of the MMV-HMP algorithm is $\mathcal{O}(TKMN)$, where $T$ denotes the number of iterations.
+VI Simulation Results
 TABLE II: Simulation Parameters
@@
 Proportion of visible antenna elements $\rho_{l}$ (0,1]
-VI Simulation Results
 In this section, we evaluate the performance of the proposed channel estimation scheme under various system setups.
@@
 In addition, the SNR is defined in received side, which is given by $10\log_{10}\left(\lVert\mathbf{WH}\rVert^{2}_{\mathrm{F}}/\lVert\mathbf{N}\rVert^{2}_{\mathrm{F}}\right)$.
-Additionally, we compare the proposed MMV-VMP algorithm with the following baselines:
+Additionally, we compare the proposed MMV-HMP algorithm with the following baselines:
   • SOMP[30]: The simultaneous OMP algorithm designed for on-grid sparse recovery, which depends on the knowledge of the number of non-zeros components.
@@
 Unlike the StdSBL and PC-SBL methods, the prior parameters in the VSP algorithm are updated using a moment-matching approach, providing a computationally efficient alternative.
+Figure 4: Convergence behavior of MMV-HMP.
+Figure 5: Complexity evaluation of different algorithms.
 VI-A Convergence and Complexity Performance
-Figure 4: Convergence behavior of MMV-VMP.
 Fig.
-4 illustrates the convergence behavior of the proposed MMV-VMP algorithm by plotting the NMSE against the number of iterations under various pilot lengths and SNR conditions.
+4 illustrates the convergence behavior of the proposed MMV-HMP algorithm by plotting the NMSE against the number of iterations under various pilot lengths and SNR conditions.
 The results clearly demonstrate a consistent monotonic decrease in NMSE across all iterations, indicating the algorithm’s stable and reliable convergence across different simulation scenarios.
@@
 Therefore, for subsequent simulations, the maximum number of iterations can be effectively limited to 20 without compromising performance.
-Figure 5: Complexity evaluation of different algorithms.
 Fig.
-5 presents a computational complexity analysis by comparing the CPU runtime of various algorithms.
+5 presents a computational complexity analysis by comparing the central processing unit (CPU) runtime of various algorithms.
 Among these, SOMP exhibits the shortest runtime.
@@
 This is primarily due to the matrix inversion operations inherent in their implementation, which increase computational complexity.
-In contrast, the proposed MMV-VMP algorithm demonstrates a shorter runtime compared to StdSBL, PC-SBL, and VSP due to avoidance of matrix inversion operations and stable convergence properties.
+As anticipated, the proposed MMV-HMP algorithm demonstrates a shorter runtime compared to StdSBL, PC-SBL, and VSP due to avoidance of matrix inversion operations and stable convergence properties.
 VI-B NMSE versus SNR and Pilot Symbol Number
-Figure 6: NMSE versus pilot symbol number.
 Fig.
 6 illustrates the NMSE performance of various algorithms as a function of SNR for $P=8$.
-The results reveal that algorithms such as SOMP [29], StdSBL [36], and UAMP-SBL [33], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-VMP.
+The results reveal that algorithms such as SOMP [36], StdSBL [46], and UAMP-SBL [40], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-HMP.
 Among block-sparsity prior-based methods, the PC-SBL algorithm leverages a pattern-coupled Gaussian prior model to effectively capture local common-delay sparsity.
@@
 In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity.
-The proposed MMV-VMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both global block sparsity and local common-delay sparsity.
+In contrast, the proposed MMV-HMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both MRF structure and precision parameter sharing mechanism.
 This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
+Figure 6: NMSE versus SNR.
 Figure 7: NMSE versus pilot symbol number.
@@
 As the number of pilot symbols varies from 6 to 12, the compression ratio $M/N$ ranges from 0.375 to 0.75.
-We can obtain two key observations: 1) Across most compression ratios, the proposed MMV-VMP algorithm outperforms baseline methods.
+We can obtain two key observations: 1) Across most compression ratios, the proposed MMV-HMP algorithm outperforms baseline methods.
 This result highlights its robust sparse recovery capability under varying compression conditions.
@@
 In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios.
-Notably, the proposed MMV-VMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
-VI-D Ablation Study to Evaluate the Effects of Prior Models
+Notably, the proposed MMV-HMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
 Figure 9: Ablation study of the column-wise hierarchical model.
-To evaluate the effectiveness of the proposed column-wise prior model, the MMV-VMP algorithm is compared with two baselines: (1) a column-wise hierarchical prior without the MRF, labeled as “Column-Wise w.o.
-MRF,” and (2) a column-wise prior without the common-delay sparsity mechanism, labeled as “Column-Wise w.o.
-CD.” Fig.
-9 reveals two key observations: 1) Both the MRF and the common-delay sparsity mechanism contribute to improved estimation performance.
-Notably, the improvement introduced by the MRF is more significant than that of the common-delay sparsity mechanism.
-2) The simultaneous use of the MRF and the common-delay sparsity mechanism better captures the structured sparsity of the channel, leading to further enhancement in estimation accuracy.
-To showcase the estimation performance of different prior more intuitively, Fig.
-10 provides the results of a Monte Carlo simulation with $\mathrm{SNR}=2$ dB and $P=8$.
-The original spatial-delay channels are compared with the reconstructed channels obtained using “Column-Wise w.o.
-MRF,” “Column-Wise w.o.
-CD,” and the proposed MMV-VMP algorithm.
-The figure clearly shows that the proposed MMV-VMP algorithm achieves the best visual quality, accurately reconstructing the spatial-delay channels and demonstrating its superior capability in sparse recovery.
-As expected, “Column-Wise w.o.
-MRF” provides the poorest reconstruction quality due to its limited ability to capture global block sparsity.
-Moreover, compared to “Column-Wise w.o.
-CD,” the reconstructed channels by MMV-VMP exhibit reduced power dispersion phenomenon near the non-zero regions, attributed to the incorporation of the common-delay sparsity mechanism.
 (a) Original.
@@
 (c) Column-Wise w.o. CD.
-(d) MMV-VMP.
-Figure 10: Original and reconstructed channels by “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and MMV-VMP algorithm.
+(d) MMV-HMP.
+Figure 10: Original and reconstructed channels by “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and MMV-HMP algorithm.
+VI-D Ablation Study to Evaluate the Effects of Prior Models
+To evaluate the effectiveness of the proposed column-wise hierarchical sparse prior model, we conduct an ablation study.
+Specifically, the proposed MMV-HMP algorithm is compared against the following two baselines:
+  • Column-Wise w.o. MRF: This baseline removes the MRF structure from the proposed column-wise hierarchical sparse prior model, thereby ignoring coefficient correlations.
+  • Column-Wise w.o.
+CD: This baseline eliminates the precision parameter sharing strategy in the proposed prior model, thus modeling each coefficient’s precision independently without promoting local clustering.
+Fig.
+9 investigates the NMSE performance versus SNR for different prior models.
+As expected, the estimation performance of the two baselines exhibits significant degradation, as they fail to fully capture both the global and local structured sparsity.
+This validates the necessity of jointly employing the MRF structure and the precision parameter sharing mechanism.
+Moreover, it is observed that the performance degradation of the “Column-Wise w.o.
+MRF” model is substantially greater than that of the “Column-Wise w.o.
+CD” model.
+This indicates that the improvement introduced by incorporating the MRF is more critical than the gain achieved by the precision parameter sharing mechanism.
+To intuitively demonstrate the estimation performance of different priors, Fig.
+10 presents the results of a Monte Carlo simulation with $\mathrm{SNR}=2$ dB and $P=8$.
+The original spatial-delay channels are compared with the reconstructed channels obtained using the baseline methods and the proposed MMV-HMP algorithm.
+The figure clearly shows that the proposed MMV-HMP algorithm achieves the best reconstruction quality, accurately recovering the spatial-delay channels and demonstrating superior sparse recovery capability.
+As expected, “Column-Wise w.o.
+MRF” yields the poorest reconstruction, with significant residual noise, mainly due to its limited ability to capture the global block sparsity.
+Furthermore, compared to “Column-Wise w.o.
+CD,” the channels reconstructed by MMV-HMP exhibit noticeably reduced power dispersion near the non-zero regions.
+This improvement is attributed to the incorporation of the precision parameter sharing mechanism.
 VII Conclusions
-In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, spatially non-stationary (SnS) properties, and dual-wideband effects.
-We first rigorously quantified the angular and delay spread effects of the SnS dual-wideband XL-MIMO channel in the angular-delay domain.
-Building upon this, we highlighted the global and local common-delay sparsity inherent in the angular-delay channels.
-Leveraging this structured sparsity, we proposed a computationally efficient MMV-VMP algorithm, where variational message passing is employed to handle the complex factor graph arising from the column-wise hierarchical prior.
-Simulation results demonstrate that the proposed MMV-VMP algorithm consistently outperforms existing methods, offering robust performance with moderate computational complexity.
-Additionally, an ablation study was conducted to assess the effectiveness of the column-wise hierarchical prior, further validating the advantages of the proposed approach.
+In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, SnS properties, and dual-wideband effects.
+We began by rigorously quantifying the angular and delay spread properties of SnS dual-wideband channels in the angular-delay domain and revealed their inherent global block sparsity and local common-delay sparsity.
+To exploit this structured sparsity, we proposed a computationally efficient MMV-HMP algorithm.
+The algorithm performs message updates along the factor graph by selectively applying either the SPA or VMP rules, depending on the structure and tractability of each factor node.
+Simulation results demonstrate that the proposed MMV-HMP algorithm consistently outperforms existing baseline methods, achieving robust estimation performance with moderate computational complexity.
+Furthermore, an ablation study confirms the effectiveness of the proposed column-wise hierarchical prior, validating its contribution to improved estimation accuracy.
 Appendix A Proof of Lemma 1
-The proof of Theorem 1 is portioned as two part.
-We first prove that the angular-delay matrix of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ is block-sparse.
-Then, applying the shift property of 2D-IDFT, the block sparsity of $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}$ can be obtained.
-For notation simplification, we discard the subscript $l$, $\varphi_{l}$, and $\psi_{l}$ in the subsequent derivations.
-We begin by examining the angular-domain transformation of $\boldsymbol{\Theta}$ with $\cos\vartheta>0$.
-Mathematically, $\boldsymbol{\Theta}$ can be rewritten in a chirp signal form as
-$[\boldsymbol{\Theta}]_{n,k}=\mathrm{e}^{-\mathrm{j}2\pi\left(a_{k}n-\frac{b_{k}}{2}n^{2}\right)},$ (A.1)
-where $a_{k}={f_{k}\psi}/{f_{c}}$ and $b_{k}={2f_{k}\varphi}/{f_{c}}$ denote the starting spatial frequency and chirp rate, respectively.
-Utilizing the inherent chirp properties of Eq.
-(A.1), the IDFT of $\boldsymbol{\Theta}$ for the fixed $k$, i.e., $[\boldsymbol{\Xi}_{\mathrm{A}}]_{:,k}=\mathbf{F}^{\mathrm{H}}_{\mathrm{A}}\boldsymbol{\theta}(f_{k})\in\mathbb{C}^{N\times 1}$ exhibits a distinct rectangular shape [38], where $\mathbf{F}_{\mathrm{A}}$ denotes a $N$-dimension DFT matrix.
-The boundary spatial frequencies are given by
-$i_{s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (A.2)
-$i_{e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d},$ (A.3)
-where $n_{s}$ and $n_{e}$ denote the start and ending index of VR $\boldsymbol{\phi}$, i.e., $\boldsymbol{\phi}\in[n_{s},n_{e}]$.
-In this case, due to $a_{k}>0$ and $b_{k}>0$, we have $i_{s}(k)>i_{e}(k)$.
-Consequently, $[\boldsymbol{\Xi}_{\mathrm{A}}]_{:,k}$ demonstrates block sparsity, with its prominent spatial frequency components concentrated from $i_{e}(k)$ to $i_{s}(k)$.
-Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{e}^{\mathrm{min}}=\min_{k}i_{e}(k)$ to $i_{s}^{\mathrm{max}}=\max_{k}i_{s}(k)$.
+The proof of Lemma 1 is divided into two parts.
+First, we demonstrate that the angular-frequency representation $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ exhibits row-wise sparsity.
+Next, we show that the spatial-delay representation $\boldsymbol{\Theta}(\varphi_{,}\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ exhibits column-wise sparsity.
+Without loss of generality, we assume that $\cos\vartheta>0$.
+We begin by examining the angular transformation of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$.
+For a fixed $k$, the angular transformation of $\boldsymbol{\theta}(f_{k},\varphi_{l},\psi_{l})$ is given by
+$[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{m,k}=\sum_{n=1}^{N}\mathrm{e}^{\mathrm{j}\frac{2\pi}{N}mn}\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n}.$ (A.1)
+According to (21) and (22), it can be obtained that $[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{:,k}$ demonstrates block sparsity with its prominent spatial frequency components concentrated from $i_{l,e}(k)$ to $i_{l,s}(k)$.
+Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{l,e}^{\mathrm{min}}=\min_{k}i_{l,e}(k)$ to $i_{l,s}^{\mathrm{max}}=\max_{k}i_{l,s}(k)$.
 As a result, $\boldsymbol{\Xi}_{\mathrm{A}}$ is characterized as a row-wise block matrix
-$\boldsymbol{\Xi}_{\mathrm{A}}=\left[\mathbf{0}_{K\times(I_{e}-1)}^{\mathrm{T}},\mathbf{U}_{K\times(I_{s}-I_{e}+1)}^{\mathrm{T}},\mathbf{0}_{K\times(N-I_{s})}^{\mathrm{T}}\right]^{\mathrm{T}},$ (A.4)
-where $\mathbf{U}$ is a non-zero matrix; index $I_{e}$ and $I_{s}$ are given by $I_{e}=\lceil{(i_{e}^{\mathrm{min}}+1)N}/{2}\rceil$ and $I_{s}=\lceil{(i_{s}^{\mathrm{max}}+1)N}/{2}$.
-Further, we examine the IDFT of $\boldsymbol{\Theta}$ in the delay domain, i.e., $\boldsymbol{\Xi}_{\mathrm{D}}=\boldsymbol{\Theta}\mathbf{F}_{\mathrm{D}}^{*}$, where $\mathbf{F}_{\mathrm{D}}$ denote a $K$-dimension DFT matrix.
+$\boldsymbol{\Xi}_{l,\mathrm{A}}=\left[\mathbf{0}_{K\times(I_{l,e}-1)}^{\mathrm{T}},\mathbf{U}_{K\times(I_{l,s}-I_{l,e}+1)}^{\mathrm{T}},\mathbf{0}_{K\times(N-I_{l,s})}^{\mathrm{T}}\right]^{\mathrm{T}},$ (A.2)
+where $\mathbf{U}$ is a non-zero matrix; index $I_{l,e}$ and $I_{l,s}$ are given by $I_{l,e}=\lceil{(i_{l,e}^{\mathrm{min}}+1)N}/{2}\rceil$ and $I_{l,s}=\lceil{(i_{l,s}^{\mathrm{max}}+1)N}/{2}\rceil$.
+Further, we examine the IDFT of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ in the delay domain, i.e., $\boldsymbol{\Xi}_{l,\mathrm{D}}=\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$, where $\mathbf{F}_{\mathrm{D}}$ denote a $K$-dimension DFT matrix.
 Specifically, we have
-n,m $\propto\frac{1}{\sqrt{K}}\sum_{k=0}^{K-1}\mathrm{e}^{\mathrm{j}\frac{2\pi}{K}mk}\mathrm{e}^{-\mathrm{j}2\pi kt_{n}}$ (A.5)
-$\propto\frac{1}{\sqrt{K}}\frac{\sin\left(\pi K(t_{n}-\frac{m}{K})\right)}{\sin\left(\pi(t_{n}-\frac{m}{K})\right)},$
-where $t_{n}=n\frac{f_{s}\psi}{Kf_{c}}-n^{2}\frac{f_{s}\varphi}{Kf_{c}}$, which is a quadratic function about $n$. Further, we have
-$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{n}-\frac{m}{K}\right).$ (A.6)
-From (A.6), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{n}K$.
-However, due to the finite sampling size $K$ in practice, we consider the range from $j_{e}(n)=(t_{n}-{2}/{K})K$ to $j_{s}(n)=(t_{n}+{2}/{K})K$ to collect the delay indices near the peak effectively.
+n,m $\propto\frac{1}{\sqrt{K}}\sum_{k=0}^{K-1}\mathrm{e}^{\mathrm{j}\frac{2\pi}{K}mk}\mathrm{e}^{-\mathrm{j}2\pi kt_{l,n}}$ (A.3)
+$\propto\frac{1}{\sqrt{K}}\frac{\sin\left(\pi K(t_{l,n}-\frac{m}{K})\right)}{\sin\left(\pi(t_{l,n}-\frac{m}{K})\right)},$
+where $t_{l,n}=n^{2}\frac{f_{s}\varphi_{l}}{Kf_{c}}-n\frac{f_{s}\psi_{l}}{Kf_{c}}$, which is a quadratic function about $n$. Further, we have
+$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{l,n}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{l,n}-\frac{m}{K}\right).$ (A.4)
+From (A.4), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{l,n}K$.
+However, due to the finite sampling size $K$ in practice, we consider the range from $J_{l,e}(n)=(t_{n}-{2}/{K}+1/2)K$ to $J_{l,s}(n)=(t_{n}+{2}/{K}+1/2)K$ to collect the delay indices near the peak effectively.
 Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
-$\boldsymbol{\Xi}_{\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{e}-1)},\mathbf{L}_{N\times(J_{s}-J_{e}+1)},\mathbf{0}_{N\times(K-J_{s})}\right],$ (A.7)
-where $J_{e}=\min_{n}j_{e}(n)$ to $J_{s}=\max_{n}j_{s}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
-According to row-wise and column-wise block sparsity of (A.4) and (A.7), we assert that $\boldsymbol{\Xi}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ adheres to the following block-sparse structure, with its significant entries localized within the square region $\mathcal{A}_{1}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{e}\leq n\leq I_{s},J_{e}\leq k\leq J_{s}\right\}$, i.e.,
+$\boldsymbol{\Xi}_{l,\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{l,e}-1)},\mathbf{L}_{N\times(J_{l,s}-J_{l,e}+1)},\mathbf{0}_{N\times(K-J_{l,s})}\right],$ (A.5)
+where $J_{l,s}=\min_{n}J_{l,s}(n)$ to $J_{l,e}=\max_{n}J_{l,e}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
+According to row-wise and column-wise block sparsity of (A.2) and (A.5), we assert that $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ adheres to the following block-sparse structure, with its significant entries localized within the square region $\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\}$, i.e.,
 […]
 where $\mathbf{V}$ denotes a non-zero matrix.
-Next, we examine the 2D-IDFT of channel $\mathbf{H}$, $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$.
+Moreover, based on the definition $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$, it follows that, for a fixed column index $k$, $\boldsymbol{\Xi}_{l}$ characterizes the angular distribution corresponding to a specific propagation delay, which is determined by the spatial frequencies across different subcarriers.
+Accordingly, within the region $\mathcal{A}_{l}$, each column captures the angular response for a given delay, and the width of significant spatial frequency components is approximately determined by the system width and the start spatial frequency, i.e., $i_{l,k}N{f_{s}}/{f_{c}}<I_{l,s}-I_{l,e}$, where $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$ denotes the starting spatial frequency of the $k$-th column.
+As a result, the nonzero components of $\boldsymbol{\Xi}_{l}$ are confined within a narrow band in each column, leading to a distinct column-wise clustered sparsity pattern.
+Appendix B A Proof of Theorem 1
+Now, we examine the 2D-IDFT of channel $\mathbf{H}$, i.e., $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$.
 By leveraging the linearity of the 2D-IDFT, we can focus solely on analyzing the $l$-th path.
-Define $\mathbf{H}_{l}=\alpha_{l}\left(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})\right)\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}$, whose $(n,k)$-th element is given by
-n,k $\propto\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}+f_{k}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k},$ (A.9)
-$\overset{(a)}{\propto}\sum_{p=1}^{P}\mathrm{e}^{-\mathrm{j}2\pi\left(\frac{n}{N}N\epsilon_{l,p}+\frac{k}{K}f_{s}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k},$
-where $(a)$ is obtained by decomposing the chirp signal $\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}$ into the superposition of $P$ complex exponential signals with frequencies $\epsilon_{l,p}=\psi_{l}-2n_{l,e}\varphi_{l}+\eta_{l}(p-1)$ with $\eta_{l}={2N\varphi_{l}}/{P}$ and $p\in\{p\in\mathbb{N}^{+}\mid\psi_{l}-2n_{l,e}\varphi_{l}\leq\eta_{l}p\leq\psi_{l}-2n_{l,s}\varphi_{l}\}$.
-Furthermore, we focus on $\mathbf{H}_{l,p}(n,k)=\mathrm{e}^{-\mathrm{j}2\pi\left(\frac{n}{N}N\epsilon_{l,p}+\frac{k}{K}f_{s}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k}$.
-Applying the shift property of 2D-IDFT, $\mathbf{X}_{l,p}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}_{l,p}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ should satisfy the block sparsity with its significant entries localized within the square region $\mathcal{A}_{l,2,p}$ with
-$\mathcal{A}_{l,2,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (A.10)
-$k$ $\left.=\mathrm{mod}(k_{l,1}+f_{s}\tau_{l},K),\forall(n_{l,1},k_{l,1})\in\mathcal{A}_{l,1}\right\}.$
-Here, $\mathcal{A}_{l,1}$ is the non-zero region for $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$, and mod$(a,m)$ is the modulus of $a$ for $m$.
-Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{A}_{l,2}\triangleq\cup_{p}\mathcal{A}_{l,2,p}$.
+according to (20), we further define
+$\mathbf{H}_{l,p}=\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (B.1)
+Applying the shift property of 2D-IDFT, $\mathbf{X}_{l,p}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}_{l,p}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ should satisfy the block sparsity with its significant entries localized within the square region $\mathcal{B}_{l,p}$ with
+$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (B.2)
+$k$ $\left.=\mathrm{mod}(k_{l,1}+f_{s}\tau_{l},K),\forall(n_{l,1},k_{l,1})\in\mathcal{A}_{l}\right\}.$
+Here, $\mathcal{A}_{l}$ is the non-zero region for $\boldsymbol{\Xi}_{l}$ given by (25), and mod$(a,m)$ is the modulus of $a$ for $m$.
+Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{B}_{l}\triangleq\cup_{p}\mathcal{B}_{l,p}$.
 Thus, the proof is completed.

### v4

@@
 Abstract
-The deployment of extremely large antenna arrays (ELAAs) in wideband extremely large-scale multiple-input multiple-output (XL-MIMO) systems introduces significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties.
-When combined with the dual-wideband effects inherent to wideband systems, these phenomena fundamentally reshape the channel’s sparsity patterns in the angular-delay domain, rendering existing sparsity-based channel estimation methods inadequate.
+The deployment of extremely large antenna arrays (ELAAs) in extremely large-scale multiple-input multiple-output (XL-MIMO) systems introduces significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties.
+When combined with the dual-wideband effects inherent to wideband systems, these phenomena fundamentally alter the channel’s sparsity patterns in the angular-delay domain, rendering existing estimation methods insufficient.
 To address these challenges, this paper reconsiders the channel estimation problem for wideband XL-MIMO systems.
-By leveraging the spatial-chirp property of array responses, we quantitatively characterize the angular-delay domain sparsity of wideband XL-MIMO channels, revealing both global block sparsity and local common-delay sparsity.
-To effectively capture this structured sparsity, we propose a novel column-wise hierarchical prior model that incorporates precision parameter sharing mechanism and Markov random field (MRF) structure.
-Based on this prior model, the channel estimation task is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem.
-Tailored to the complex factor graph induced by this hierarchical prior, we develop an MMV-based hybrid message passing (MMV-HMP) algorithm.
-The algorithm performs message updates along the factor graph edges by selectively applying either the sum-product algorithm (SPA) or variational message passing (VMP) rules, depending on the factor-node structure and message tractability.
-Simulation results validate the effectiveness of the proposed column-wise hierarchical prior model through comprehensive ablation studies and demonstrate that the MMV-HMP algorithm consistently outperforms existing baselines, achieving robust estimation performance with moderate computational complexity.
+Leveraging the spatial-chirp property of array responses, we first quantitatively characterize the angular-delay domain sparsity of wideband XL-MIMO channels, revealing both global block sparsity and local common-delay sparsity.
+To effectively capture this structured sparsity, we then propose a novel column-wise hierarchical prior model that integrates a precision sharing mechanism and a Markov random field (MRF) structure.
+Building on this prior model, the channel estimation task is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem.
+Tailored to the complex factor graph induced by this hierarchical prior, we develop a MMV-based hybrid message passing (MMV-HMP) algorithm.
+This algorithm performs message updates along the edges of the factor graph, and selectively applies either the variational message passing (VMP) or sum-product (SP) rules, depending on the factor-node structure and message tractability.
+Simulation results validate the effectiveness of the proposed column-wise hierarchical prior model through ablation studies and demonstrate that the MMV-HMP algorithm, while maintaining moderate computational complexity, consistently outperforms existing baselines which fail to capture the structured sparsity of wideband XL-MIMO channels.
 Index Terms:
-Wideband XL-MIMO systems, dual-wideband effects, near-field effects, channel estimation, Bayesian inference.
+Wideband XL-MIMO systems, dual-wideband effects, near-field effects, channel estimation, message passing.
 I Introduction
-Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2, 3].
-By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[4, 5].
-It also delivers substantial beamforming gains, effectively mitigating the severe path loss characteristic of millimeter-wave (mmWave) frequencies.
+Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2].
+By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[3, 4, 5].
+It also delivers substantial beamforming gains, effectively compensating the severe path loss of millimeter-wave (mmWave) frequencies.
 Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [6, 7].
-However, realizing these performance benefits critically depends on obtaining accurate channel state information (CSI), underscoring the importance of precise channel knowledge.
+However, realizing these performance benefits critically depends on the accurate channel state information (CSI).
 Typically, wideband XL-MIMO systems adopt orthogonal frequency division multiplexing (OFDM) to combat the frequency-selective fading of wireless channels [8].
@@
 Specifically, the large system bandwidth shortens the OFDM symbol duration, potentially causing the propagation delay difference across the antenna array to exceed the symbol duration, leading to the spatial wideband effect [9].
-Concurrently, the varying central frequencies of subcarriers introduce the frequency-wideband effect, where phase shifts differ across subcarrier channels, further complicating the channel estimation process.
-Moreover, with the deployment of ELAAs and operation at higher frequency bands, near-field effects, such as the spherical wavefront effect [10, 11, 12, 13] and spatially non-stationary (SnS) properties [14, 15], become prominent, exacerbating the dual-wideband effects.
-For instance, the sparsity structure of SnS dual-wideband XL-MIMO channels in the angular-delay domain undergoes significant changes.
-Specifically, the curvature of the spherical wavefront renders a single spatial frequency insufficient to characterize a near-field propagation path, requiring multiple spatial frequencies instead.
-This spatial frequency spread, combined with the frequency-wideband effect, amplifies angular spread compared to conventional massive MIMO systems [9, 16].
-Additionally, the spherical wavefront effect disrupts the linear path delay variations across adjacent antennas, while SnS properties influence both spatial frequency and delay spreads.
-These combined factors reshape the sparse patterns of SnS dual-wideband XL-MIMO channels in the angular-delay domain, significantly complicating channel estimation in wideband XL-MIMO systems.
+Concurrently, the varying central frequencies of subcarriers introduce the frequency-wideband effect, where phase shifts differ across subcarrier channels, further complicating the acquisition of CSI.
+Moreover, with the deployment of ELAAs and operation at higher frequency bands, near-field effects, such as the spherical wavefront effect [10, 11, 12] and spatially non-stationary (SnS) properties [13, 14], become prominent, exacerbating the dual-wideband effects.
+For instance, the sparsity structure of XL-MIMO channels in the angular-delay domain alters significantly.
+Specifically, the curvature of the spherical wavefront means a single spatial frequency is no longer sufficient to characterize a propagation path; instead, multiple spatial frequencies are required.
+This spatial frequency spread, combined with the frequency-wideband effect, results in a more pronounced angular spread compared to that of conventional massive MIMO systems [9, 15].
+Additionally, the spherical wavefront effect disrupts the linear variations of path delay across antennas, while SnS properties influence both spatial frequency and delay spreads.
+These combined factors reshape the sparse patterns of XL-MIMO channels in the angular-delay domain, significantly complicating channel estimation in wideband XL-MIMO systems.
 I-A Related Works
@@
 Specifically, a well-established channel model for wideband XL-MIMO systems is lacking, and the sparsity patterns of these channels in the angular-delay domain have yet to be fully explored.
-Additionally, with the increasing number of antennas and subcarriers in XL-MIMO systems, there is a pressing need for computationally efficient channel estimation algorithms to handle the growing complexity of these systems.
+Additionally, with the increasing number of antennas and subcarriers in XL-MIMO systems, there is a pressing need for computationally efficient channel estimation algorithms to handle the growing complexity.
 To address these issues, this paper investigates channel estimation for XL-MIMO systems by incorporating the spherical wavefront effect, SnS properties, and dual-wideband effects.
@@
 Furthermore, we reveal the resulting sparsity patterns in the angular-delay domain, which exhibit both global block sparsity and local common-delay sparsity.
-  • Exploiting the sparsity in the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation problem is formulated as a multiple measurement vector (MMV)-based sparse recovery task.
-Recognizing the importance of an accurate prior model in channel reconstruction, we introduce a novel column-wise hierarchical prior, which incorporates both a precision parameter sharing mechanism and a Markov random field (MRF) structure.
-This design enables the model to effectively capture global block sparsity and local common-delay sparsity, thereby providing a more structured and accurate representation of the underlying channel characteristics.
-  • To achieve efficient Bayesian inference on the intricate factor graph, resulting from the incorporation of both the precision parameter sharing mechanism and the MRF structure, we propose a MMV-based hybrid message passing (MMV-HMP) algorithm that jointly leverages belief propagation (BP) and variational message passing (VMP) rules.
-Thanks to this flexible message update strategy, the MMV-HMP algorithm can effectively manage the message updates for latent precision parameters with non-Gaussian distributions and accommodate the loops inherent in the complex factor graph, thereby enabling accurate and reliable inference.
-  • Compared to existing approaches, the proposed MMV-HMP algorithm achieves superior performance by jointly capturing the global block sparsity and local common-delay sparsity inherent in SnS dual-wideband XL-MIMO channels.
-This enables robust channel estimation while maintaining moderate computational complexity.
-In addition, the effectiveness of the proposed column-wise hierarchical prior model is validated through comprehensive ablation studies.
+  • Exploiting the sparsity in the angular-delay domain, the channel estimation problem is formulated as a multiple measurement vector (MMV)-based Bayesian inference task.
+Recognizing the importance of an accurate prior model in channel reconstruction, we devise a novel column-wise hierarchical prior, which incorporates both a precision sharing mechanism and a Markov random field (MRF) structure.
+This design enables the model to effectively capture global block sparsity and local common-delay sparsity, thereby providing a more structured representation of the underlying channel characteristics.
+  • To achieve efficient Bayesian inference, resulting from the hierarchical prior, we propose a MMV-based hybrid message passing (MMV-HMP) algorithm that jointly leverages sum-product (SP) and variational message passing (VMP) rules.
+Thanks to this flexible message update strategy, the MMV-HMP algorithm can effectively manage the message updates for latent precision parameters with non-Gaussian distributions and accommodate the loops inherent in the complex factor graph, thereby enabling accurate inference.
+  • Compared to existing approaches, the proposed MMV-HMP algorithm achieves superior performance by jointly capturing the global block sparsity and local common-delay sparsity while maintaining moderate computational complexity.
+In addition, comprehensive ablation studies validate the effectiveness of the proposed column-wise hierarchical prior model, confirming its ability to capture the channel’s structured sparsity.
 I-C Organization and Notations
 The rest of the paper is organized as follows: Section II introduces the mmWave XL-MIMO system model.
-Section III develops a spatial-frequency channel model, incorporating spherical wavefront effects, SnS properties, and dual-wideband effects, and analyzes the angular-delay characteristics of SnS dual-wideband XL-MIMO channels.
-Section IV formulates the channel estimation problem as a sparse signal recovery task and proposes a column-wise hierarchical sparse prior model.
+Section III develops a spatial-frequency channel model, incorporating near-field and dual-wideband effects, and analyzes the angular-delay characteristics of XL-MIMO channels.
+Section IV formulates the channel estimation problem and proposes a column-wise hierarchical sparse prior model.
 Section V presents the MMV-HMP algorithm.
 Simulation results and conclusions are provided in Sections VI and VII, respectively.
-Notations: lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\odot$ denotes the Hadamard product; $\mathbf{I}_{N}$ denotes a $N\times N$ identity matrix; $\mathbb{C}^{M\times N}$ denotes a $M\times N$ complex matrix.
+Notations: lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\otimes$ denotes the Kronecker product; $\mathbf{I}_{N}$ denotes a $N\times N$ identity matrix; $\mathbb{C}^{M\times N}$ denotes a $M\times N$ complex matrix.
 $\mathbf{1}_{N}$ denotes a $N\times 1$ all-one vector.
 In addition, a random variable $x\in\mathbb{C}$ drawn from the complex Gaussian distribution with mean $x_{0}$ and variance $v$ is characterized by the probability density function $\mathcal{CN}(x;m,v)=\exp\{-{\left|x-m\right|^{2}}/{v}\}/{\pi v}$; a random variable $\gamma\in\mathbb{R}$ from Gamma distribution with mean $a/b$ and variance $a/b^{2}$ is characterized by the probability density function $\mathcal{G}a(\gamma;a,b)\propto\gamma^{a-1}\exp(-\gamma b)$.
+Figure 1: Illustration of the XL-MIMO system model.
 II System Model
@@
 Therefore, channel estimation for each user is independent.
-Without loss of generality, we consider an arbitrary user and assume $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
-Figure 1: Illustration of the XL-MIMO system model.
+Without loss of generality, we consider an arbitrary user and assume $K$ subcarriers are uniformly selected from the available subcarriers to carry pilots.
 In the uplink channel estimation phase, the BS combines the pilot signal using all RF chains associated with different beams.
@@
 Assume that there are $L$ propagation paths between the BS and user.
-Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l\geq 2$ indicates reflection or diffraction path, as illustrated in Fig.
+Denote $\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l\geq 2$ indicates reflection or diffraction path, as illustrated in Fig.
 1.
@@
 where $\boldsymbol{\phi}_{l}=[n_{l,s},n_{l,e}]$ denotes the VR of the $l$-th path with $n_{l,s}$ and $n_{l,e}$ indicating the start and end index [^1] For the diffraction paths, $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ is associated with the diffraction coefficient.
-As such, $\kappa_{l,n}$ can be set within $10<\kappa_{l,n}<$ 1, with the reference value chosen as the element with the highest received power.
+As such, $\kappa_{l,n}$ can be set within $0<\kappa_{l,n}<1$, with the reference value chosen as the element with the highest received power.
 Thus, we assume $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ obeys a uniform distribution between 0 and 1.
@@
 $r_{l,n}=$ $\sqrt{(r_{l}\cos\vartheta_{l}-nd)^{2}+r_{l}^{2}\sin^{2}\vartheta_{l}}$ (7)
-$\overset{(a)}{\approx}$ $r_{l}-nd\cos\vartheta_{l}+\frac{n^{2}d^{2}}{2r_{l}}{\sin^{2}\vartheta_{l}}.$
-where $(a)$ is obtained by using the Fresnel approximation, i.e., $\sqrt{1+x}\approx 1+x/2-x^{2}/8$.
-In this Taylor expansion, the current point is set at $n=0$, and terms of third order and higher are neglected.
+$\overset{(a)}{\approx}$ $r_{l}-nd\cos\vartheta_{l}+\frac{n^{2}d^{2}}{2r_{l}}{\sin^{2}\vartheta_{l}},$
+where $(a)$ is obtained by applying the second-order Taylor approximation with the current point set at $n=0$ and terms of third order and higher neglected.
+Notably, this approximation maintains high accuracy when the distance between the BS and the user or scatterer exceeds the Fraunhofer distance [29].
+Additionally, unlike the Euclidean-domain model in [32], the approximation enables the array response vector to align with a spatial chirp form, simplifying subsequent analysis.
 Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}-n{\psi_{l}}/{f_{c}}+n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$.
@@
 where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain.
-Applying the Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (9)[33, 34], as shown in the top of this page.
+Applying the Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (9)[33, 34], as shown in the top of the next page.
 $h_{n}(f)=\int_{-\infty}^{+\infty}h_{n}(\tau)\mathrm{e}^{-\mathrm{j}2\pi f\tau}\mathrm{d}\tau=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\mathrm{e}^{\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}.$ (9)
-Further, Eq. (9) can be expressed in a more compact form as
+Further, (9) can be expressed in a more compact form as
 $\mathbf{h}(f)=\sum_{l=1}^{L}\alpha_{l}a(f,\tau_{l})(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l}))\odot\boldsymbol{\theta}(f,\psi_{l},\varphi_{l}),$ (10)
@@
 Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each pilot subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K-1$.
-The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in Eq.
-(10) with $f$ replaced by $f_{k}$.
+The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in (10) with $f$ replaced by $f_{k}$.
 Thus, the overall spatial-frequency channel matrix $\mathbf{H}$ can be reformulated as
@@
 Eq.
-(13) presents a comprehensive channel model for XL-MIMO systems, integrating spherical wavefront effects, SnS properties, as well as spatial and frequency-wideband characteristics.
+(13) presents an approximated channel model for XL-MIMO systems, integrating spherical wavefront effects, SnS properties, as well as spatial and frequency-wideband characteristics.
 Specifically, the spatial-domain steering vector $\mathbf{b}(\psi_{l},\varphi_{l})$ is coupled with the SnS indicator vector $\boldsymbol{\kappa}_{l}$, the frequency response vector $\mathbf{a}(\tau_{l})$, and the frequency-dependent phase matrix $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$.
-However, due to the high dimensionality of spatial-frequency channels in mmWave XL-MIMO systems, direct estimation becomes computationally prohibitive.
-Thus, it is essential to leverage channel sparsity in a suitable transform domain.
 III-B Angular-Delay Representation of SnS XL-MIMO Channel
-Building on approaches used for traditional massive MIMO channels [9], the wideband XL-MIMO channels in (13) can be approximated as
+To enable efficient channel estimation, leveraging channel sparsity is critical.
+Regarding the spatial domain, the Euclidean and polar domains are theoretically preferable transformation choices where each propagation path in these domains may correspond to a single line spectrum, yielding a highly sparse representation[21].
+However, these domains rely on joint two-dimensional sampling, which leads to transformation matrices with drastically high dimensionality, and consequently, heightened computational complexity.
+As an alternative, we propose leveraging angular-delay sparsity for estimation [9]. Thus, the spatial-frequency channels in (13) can be approximated as
 $\mathbf{H}=\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}},$ (14)
@@
 By exploiting the chirp-like properties described in (15), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})$ can be approximated as [35]
-$i_{l,s}=\psi_{l}-2\varphi_{l}{n_{l,s}},\quad i_{l,e}=\psi_{l}-2\varphi_{l}{n_{l,e}}.$ (17)
+$\mathcal{I}_{l,s}=\psi_{l}-2\varphi_{l}{n_{l,s}},\quad\mathcal{I}_{l,e}=\psi_{l}-2\varphi_{l}{n_{l,e}}.$ (17)
 In this manner, the array response vector $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})$, which captures both the spherical wavefront effect and the SnS characteristics, can be approximated as a superposition of multiple far-field array response components, i.e.,
-$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\int_{i_{l,e}}^{i_{l,s}}c(\epsilon_{l})\mathbf{b}_{\mathrm{far}}(\epsilon_{l})\mathrm{d}\epsilon_{l},$ (18)
-where $c(\epsilon_{l})$ denotes the decomposition coefficient, and $[\mathbf{b}_{\mathrm{far}}(\epsilon_{l})]_{n}=\mathrm{e}^{\mathrm{j}2\pi\epsilon_{l}n}$ represents the far-field array response corresponding to spatial frequency $\epsilon_{l}\in[i_{l,e},i_{l,s}]$.
+$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\int_{\mathcal{I}_{l,e}}^{\mathcal{I}_{l,s}}c(\epsilon_{l})\mathbf{b}_{\mathrm{far}}(\epsilon_{l})\mathrm{d}\epsilon_{l},$ (18)
+where $c(\epsilon_{l})$ denotes the decomposition coefficient, and $\mathbf{b}_{\mathrm{far}}(\epsilon_{l})]_{n}=\mathrm{e}^{\mathrm{j}2\pi\epsilon_{l}n}$ represents the far-field array response corresponding to spatial frequency $\epsilon_{l}\in[\mathcal{I}_{l,e},\mathcal{I}_{l,s}]$.
 Furthermore, the continuous decomposition in (18) can be discretized with a spatial frequency resolution of $2/N$ as
 $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\sum_{p=1}^{P}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p}),$ (19)
-where the number of sampling points is given by $P=\lceil({n_{l,s}}-{n_{l,e}})\varphi_{l}N\rceil$, and the discretized spatial frequency is defined as $\epsilon_{l,p}=\min(i_{l,s},i_{l,e})+{2(p-1)}/{N}$ for $p=\{1,2,\dots,P\}$.
+where the number of sampling points is given by $P=\lceil({n_{l,s}}-{n_{l,e}})\varphi_{l}N\rceil$, and the discretized spatial frequency is defined as $\epsilon_{l,p}=\min(\mathcal{I}_{l,s},\mathcal{I}_{l,e})+{2(p-1)}/{N}$ for $p=\{1,2,\dots,P\}$.
 The coefficient $c_{l,p}$ denotes the corresponding weight associated with the $p$-th far-field array response component.
@@
 $\mathbf{H}=\sum_{l=1}^{L}\sum_{p=1}^{P}\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (20)
-Similar to the analysis in (17), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})$ can be approximated as
-$i_{l,s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (21)
-$i_{l,e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d}.$ (22)
+Similar to the analysis in (17), the start and end spatial frequencies of the frequency-dependent phase-shift vector $\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})$ can be approximated as
+$i_{l,s}(k)$ $=(a_{l,k}-b_{l,k}n_{l,s}){\lambda_{\mathrm{c}}}/{d},$ (21)
+$i_{l,e}(k)$ $={(a_{l,k}-b_{l,k}n_{l,e})}\lambda_{\mathrm{c}}/{d}.$ (22)
 In addition, for notation simplification, we define
@@
 Lemma 1.
-By utilizing equations (20)-(22), it can be shown that the two-dimensional inverse discrete Fourier transform (2D-IDFT) of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ results in a sparse matrix, where the nonzero entries are confined within a square region.
+By utilizing (20)-(22), it can be shown that the two-dimensional inverse discrete Fourier transform (2D-IDFT) of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ results in a sparse matrix, where the nonzero entries are confined within a square region.
 Within this region, the dominant components exhibit a clear column-wise clustered sparsity structure.
 Specifically, as $N,K\rightarrow\infty$, the support of the nonzero region converges to:
-$[\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{l},\\ 0,&(n,k)\notin\mathcal{A}_{l},\end{cases}$ (24)
+$\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{l},\\ 0,&(n,k)\notin\mathcal{A}_{l},\end{cases}$ (24)
 where the set $\mathcal{A}_{l}$ is defined as
-$\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\},$ (25)
+$\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,s}\leq k\leq J_{l,e}\right\},$ (25)
 with
@@
 $I_{l,s}$ $=\max_{k}\lceil(i_{l,s}(k)+1)N/2\rceil,$
-$J_{l,e}$ $=\max_{n}(t_{l,n}+{2}/{K}+1/2)K,$
-$J_{l,s}$ $=\min_{n}(t_{l,n}-{2}/{K}+1/2)K,$
-indicating the boundaries.
-In addition, within the region $\mathcal{A}_{l}$, each column corresponds to the angular response associated with a specific propagation delay.
-The width of the significant spatial frequency components is primarily determined by the frequency-width effect.
-Assuming the starting spatial frequency of the $k$-th column is denoted by $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$, the corresponding angular spread can be approximated as
-$B_{l,k}=\frac{i_{l,k}f_{s}}{f_{c}}N.$ (27)
-As a result, the significant components of $\boldsymbol{\Xi}_{l}$ within $\mathcal{A}_{l}$ are confined to narrow bands along each column, thereby exhibiting a pronounced column-wise clustered sparsity pattern.
-To summarize, by leveraging (24)-(27), the sparse structure of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ can be accurately characterized.
-In particular, equation (24) delineates the overall boundaries of the nonzero region, while (27) provides a more precise description of the angular spread in each column, further refining the support of the dominant components.
+$J_{l,s}$ $=\min_{n}j_{l,s}(n),\,J_{l,e}=\max_{n}j_{l,e}(n),$
+indicating the boundaries with $j_{l,s}(n)=(t_{n}-{2}/{K}+1/2)K$ and $j_{l,e}(n)=(t_{n}+{2}/{K}+1/2)K$.
 Proof.
 Please see Appendix A. ∎
+Within the region $\mathcal{A}_{l}$, each column corresponds to the angular response associated with a specific propagation delay.
+The width of the significant spatial frequency components is primarily determined by the frequency-width effect.
+Assuming the starting spatial frequency of the $k$-th column is denoted by $i_{l,k}\in[\min_{k}i_{l,e}(k),\max_{k}i_{l,s}(k)]$, the corresponding angular spread can be approximated as
+$B_{l,k}=\frac{i_{l,k}f_{s}}{f_{c}}N.$ (27)
+As a result, the significant components of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ within $\mathcal{A}_{l}$ are confined to narrow bands along each column, thereby exhibiting a pronounced column-wise clustered sparsity pattern.
+To summarize, by leveraging (24)-(27), the sparse structure of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ can be accurately characterized.
+In particular, (24) determines the overall boundaries of the nonzero region, while (27) provides a more precise description of the angular spread in each column, further refining the support of the dominant components.
 Theorem 1.
@@
 For the $l$-th path, the support region of the nonzero components is denoted by $\mathcal{B}_{l}$, which can be expressed as $\mathcal{B}_{l}=\bigcup_{p}\mathcal{B}_{l,p}$, where $\mathcal{B}_{l,p}$ represents the shifted version of $\mathcal{A}_{l}$ in both the angular and delay domains, and is defined as
-$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (28)
+$\mathcal{B}_{l,p}$ $=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right$. (28)
 $k$ $\left.=\mathrm{mod}(k_{l}+f_{s}\tau_{l},K),\forall(n_{l},k_{l})\in\mathcal{A}_{l}\right\},$
 where $\mathrm{mod}(a,m)$ is the modulus of $a$ for $m$.
+Proof.
+Please see Appendix B. ∎
 According to Theorem 1, the nonzero region of $\mathbf{X}$ is formed by aggregating the shifted nonzero regions of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ across both the angular and delay domains for all paths $l$.
 The shift positions are jointly determined by the spatial shift $N\epsilon_{l,p}$ and the delay shift $f_{s}\tau_{l}$, corresponding to the spatial frequency and propagation delay of each path, respectively.
-Proof.
-Please see Appendix B. ∎
 Moreover, we analyze the effects of the spherical wavefront, SnS properties, and dual-wideband effects on the sparsity structure of $\mathbf{X}$ based on Lemma 1 and Theorem 1 from the perspectives of the angular and delay spread.
   • Angular Spread: First, the spherical wavefront effect and the SnS characteristics determine the angular spread range at each subcarrier, characterized by $i_{l,s}(k)$ and $i_{l,e}(k)$.
-Then, the frequency-wideband effect governs the overall boundaries of the angular spread, denoted by $I_{l,e}$ and $I_{l,s}$.
-Moreover, the frequency-wideband effect also enables a more accurate description of the angular spread within each column, represented by $B_{l,k}$.
-  • Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread.
-The SnS properties define the range of variation for $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
-In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively govern the angular and delay spreads, thereby reshaping the sparsity structure of SnS dual-wideband channels in the angular-delay domain.
+Then, the frequency-wideband effect governs the overall boundaries of the angular spread, denoted by $I_{l,e}$ and $I_{l,s}$, and enables a more accurate description of the angular spread within each column, represented by $B_{l,k}$.
+  • Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, which in turn leads to delay spread.
+The SnS properties define the range of variation for the antenna index $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
+To summarize, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively govern the angular and delay spreads, thereby reshaping the sparsity structure of XL-MIMO channels in the angular-delay domain.
 Specifically,
@@
 Once the angular-delay channel is reconstructed, key parameters such as AoAs, VRs, and the range of beam squint can be efficiently extracted.
-(a) Angular spread for $k=0$ and $K-1$
+(a) Angular spread for $k=0$ and K1K-1
 (b) Delay spread for $n=n_{s}$ and $n_{e}$
@@
 Figure 2: Illustration of block sparsity of SnS XL-MIMO channel in the angular-delay domain.
-To enhance intuitive understanding of the SnS XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$.
+To enhance intuitive understanding of the XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$.
 Fig.
@@
 Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,s}(n_{s})=513$, and $j_{l,e}(n_{s})\approx 517$.
-Similarly, $i_{l,s}(K-1)\approx 0.88$, $i_{l,e}(K-1)\approx 0.72$, $j_{l,s}(n_{e})=545$, and $j_{l,e}(n_{e})\approx 549$.
+Similarly, $i_{l,e}(K-1)\approx 0.72$, $i_{l,s}(K-1)\approx 0.88$, $j_{l,s}(n_{e})=545$, and $j_{l,e}(n_{e})\approx 549$.
 As depicted in Fig.
@@
 2(b), these values provide accurate boundaries for the ranges of angular and delay spread.
-Additionally, the figures show that both the angular and delay spectra shift to the right due to $\cos\vartheta_{l}>0$ and the monotonicity of $t_{n}$ with respect to $n$.
-Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, $J_{l,s}=549$, and $B_{k,s}\approx 41$ with $k=J_{l,e}$ Fig.
+Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,s}=513$, $J_{l,e}=549$, and $B_{k,s}\approx 41$ with $k=J_{l,e}$.
+Fig.
 2(c) illustrates the non-zero region of $\mathbf{F}_{\mathrm{A}}^{{}_{\mathrm{H}}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l}$, validating the effectiveness of Lemma 1.
-In particular, for the $J_{l,e}$-th column, the width of angular spread also aligns the result of (27), which further validates the influence of frequency-wideband effect for local angular spread.
+In particular, for the $J_{l,s}$-th column, the width of angular spread also aligns the result of (27), which further validates the influence of frequency-wideband effect for local angular spread.
 Additionally, we have $f_{s}\tau_{l}=100$, $\epsilon_{l,1}=\psi_{l}-2n_{l,s}\varphi_{l}\approx 0.4$, and $\epsilon_{l,P}=\psi_{l}-2n_{l,e}\varphi_{l}=0.32$, corresponding to $N\epsilon_{l,1}=410$ and $N\epsilon_{l,P}=328$, respectively.
@@
 Remark 3.
-Compared to conventional massive MIMO-OFDM channels, the proposed SnS dual-wideband XL-MIMO channel exhibits distinct behavior.
-Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular diffusion.
+Compared to massive MIMO-OFDM channels, the wideband XL-MIMO channel exhibits distinct behavior.
+Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular spread.
 Moreover, the quadratic phase variations introduced by the spherical wavefront reshape the trend of delay variations across antennas.
-Notably, Lemma 1 is applicable to conventional MIMO-OFDM channels.
+Notably, Lemma 1 and Theorem 1 are applicable to massive MIMO-OFDM channels.
 When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted.
 Thus, we have $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$.
-Consequently, the significant components are uniformly distributed within the non-zero region, as illustrated in [9].
+Consequently, the significant components are uniformly distributed in the non-zero region [9].
 IV Problem Formulation and Sparse Prior Modeling
-In this section, leveraging the sparsity outlined in Theorem 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task.
-Subsequently, to account for the global block sparsity and local common-delay sparsity, we introduce a novel column-wise sparse prior model tailored to this structure.
+In this section, leveraging the sparsity outlined in Theorem 1, we first formulate the XL-MIMO channel estimation problem as a sparse recovery task.
+Subsequently, to capture the global block sparsity and local common-delay sparsity, we introduce a column-wise sparse prior model.
 IV-A Problem Formulation
-Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a MMV-based sparse recovery problem.
+Motivated by the sparsity of XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a MMV-based sparse recovery problem.
 Utilizing the angular-delay representation in (14), the received signal model in (3) can be further written as
@@
 Remark 4.
-While this work primarily focuses on the ULA configuration, the signal models in (29) and (30) are also applicable to the uniform planar array (UPA) configuration.
-Specifically, consider a BS equipped with a half-wavelength UPA comprising $N=N_{v}\times N_{h}$ antennas, where $N_{v}$ and $N_{h}$ denote the number of elements along the vertical and horizontal dimensions, respectively.
-In this case, the angular transformation matrix $\mathbf{F}_{\mathrm{A}}$ can be redefined as the Kronecker product $\mathbf{F}_{\mathrm{A}}=\mathbf{F}_{\mathrm{A},v}\otimes\mathbf{F}_{\mathrm{A},h}$, where $\mathbf{F}_{\mathrm{A},v}\in\mathbb{C}^{N_{v}\times N_{v}}$ and $\mathbf{F}_{\mathrm{A},h}\in\mathbb{C}^{N_{h}\times N_{h}}$ denote the vertical and horizontal DFT matrices, respectively.
-It is worth noting that the sparsity structure in the angular-delay domain may differ between ULA and UPA configurations, particularly with regard to local common-delay sparsity, due to the inherent differences in array geometry and two-dimensional angular resolution.
-Although a comprehensive analysis of the UPA case is beyond the scope of this paper, we acknowledge its importance and leave it as a promising direction for future work.
-In this paper, our primary objective is to develop an effective estimation algorithm for accurately reconstructing $\mathbf{X}$ or $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$.
-In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (30), including SOMP [30, 36] and optimization-based methods [37, 38].
-However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional parameters, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available.
+While this work primarily focuses on the ULA configuration, the signal model in (30) is also applicable to the uniform planar array (UPA) configuration.
+Specifically, consider a half-wavelength UPA comprising $N=N_{v}\times N_{h}$ antennas, where $N_{v}$ and $N_{h}$ denote the number of elements along the vertical and horizontal dimensions, respectively.
+In this case, the angular transformation matrix $\mathbf{F}_{\mathrm{A}}$ can be redefined as $\mathbf{F}_{\mathrm{A}}=\mathbf{F}_{\mathrm{A},v}\otimes\mathbf{F}_{\mathrm{A},h}$, where $\mathbf{F}_{\mathrm{A},v}\in\mathbb{C}^{N_{v}\times N_{v}}$ and $\mathbf{F}_{\mathrm{A},h}\in\mathbb{C}^{N_{h}\times N_{h}}$ denote the vertical and horizontal DFT matrices, respectively.
+This paper aims to develop an effective estimation algorithm for reconstructing $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$.
+So far, several sparse signal recovery algorithms have been employed to address the problem in (30), including SOMP [29, 36] and optimization-based methods [37, 38].
+However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional knowledge, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available.
 Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and SBL have shown superior recovery performance in channel estimation tasks [39, 40, 41].
@@
 IV-B Column-Wise Hierarchical Sparse Prior Model
-According to Lemma 1 and Theorem 1, the angular-delay domain channel exhibits prominent block sparsity and a column-wise clustered structure.
+According to Lemma 1 and Theorem 1, the angular-delay domain channel exhibits prominent block sparsity and a column-wise clustered sparsity structure.
 However, existing sparse prior models fail to capture these characteristics fully [40, 39, 42, 41].
-For example, the two-layer hierarchical prior models in [40, 39] assign independent priors to the variance precision parameters of the channel coefficients, which limits their ability to model the correlation across coefficients.
+For example, the two-layer hierarchical prior models in [40, 39] assign independent priors to the precision parameters of the channel coefficients, which limits their ability to model the correlation across coefficients.
 Although [42, 41] leverage MRF to promote global block sparsity, they still overlook local structural dependencies, especially those induced by column-wise clustering.
-As a result, these models are not well suited for the SnS wideband XL-MIMO channel estimation problem considered in this work.
-To this end, we propose a novel column-wise hierarchical sparse prior model, which incorporates both precision parameter sharing mechanism and MRF structure to promote structured sparsity.
+As a result, these models are not well suited for the XL-MIMO channel estimation problem considered in this work.
+To this end, we propose a novel column-wise hierarchical sparse prior model, which incorporates both precision sharing mechanism and MRF structure to promote structured sparsity.
 Specifically, the column-wise hierarchical prior is defined as
@@
 We next detail the distributions associated with each layer.
-1) Variance-Driven Sparsity: The first layer adopts a complex Gaussian prior with zero mean and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
+1) Variance-Driven Sparsity: In the first layer, we adopt a complex Gaussian prior with mean zero and variance $\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$ for each coefficient, i.e.,
 $p(\mathbf{X}|\boldsymbol{\Gamma})=\prod_{n=1}^{N}\prod_{k=1}^{K}\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1}),$ (32)
-where $\gamma_{n,k}$ denotes the variance precision parameter, with $\gamma_{n,k}^{-1}$ representing the variance of $x_{n,k}$.
-This precision-based modeling naturally induces sparsity: as the variance $\gamma_{n,k}^{-1}$ approaches zero, the corresponding $x_{n,k}$ is effectively pushed toward zero, thereby promoting a sparse solution.
-2) Precision Parameters Sharing Mechanisms: Due to the finite angular and delay resolutions, the angular-delay domain channel does not exhibit strict sparsity, as the coefficients near the boundaries of the sparse patterns often retain non-negligible energy.
-Enforcing hard zeros on these components may lead to degraded estimation performance.
-To mitigate this issue, in the second layer, we adopt a dual-precision parameter strategy, wherein each variance $\gamma_{n,k}$ is associated with two distinct precision parameters corresponding to its active and inactive states.
-In addition, to capture the local common-delay clustered sparsity, we further introduce the precision parameters sharing mechanism.
-Specifically, $\gamma_{n,k}$ is assigned a conditionally Bernoulli-Gaussian distribution given by
+where $\gamma_{n,k}^{-1}$ denotes the variance corresponding to coefficient $x_{n,k}$.
+This variance-driven modeling naturally induces sparsity.
+Specifically, as the variance approaches zero, the corresponding $x_{n,k}$ is effectively pushed toward zero, thereby promoting a sparse solution.
+2) Precision Sharing Mechanism: Considering the sparsity of angular-delay channel, we adopt a dual-precision strategy, where each $\gamma_{n,k}$ is associated with two distinct precision parameters corresponding to the active and inactive states of channel coefficients.
+Building on this, to capture the local common-delay clustered sparsity, we further introduce the precision sharing mechanism, where $\gamma_{n,k}$ is assigned a conditionally Bernoulli-Gamma distribution as
 $p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})$ $=\prod_{n=1}^{N}\prod_{k=1}^{K}\delta\left(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k})\right),$ (33)
@@
 $p(\boldsymbol{\Omega})=\prod_{i=1}^{2}\prod_{k=1}^{K}p(\alpha_{k}^{i})=\prod_{i=1}^{2}\prod_{k=1}^{K}\mathrm{Ga}(\alpha_{k}^{i};a_{i},b_{i}).$ (34)
-As described in (33) and (34), this common-delay sparsity mechanism ensures that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter.
-This coupling enforces a consistent sparsity structure among components associated with the same delay.
+As described in (33) and (34), this common-delay sparsity mechanism ensures that, for a fixed $k$, all $x_{n,k}$ sharing the same state are governed by a common precision parameter.
+This coupling enforces a consistent clustered sparsity structure among components associated with the same delay, which aligns with the local common-delay sparsity.
 3) MRF-Based Variance State Modeling: In the third layer, to capture the global block sparsity, the variance state variable $\mathbf{S}$ is modeled as a MRF
 $p(\mathbf{S})=\left(\prod_{n,k}\prod_{(n^{\prime},k^{\prime})\in\mathcal{D}_{n,k}}u(s_{n,k},s_{n^{\prime},k^{\prime}})\right)^{\frac{1}{2}}\prod_{n,k}v(s_{n,k}),$ (35)
-where $\mathcal{D}{n,k}$ denotes the set of neighboring nodes of $s_{n,k}$; $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$ and $v(s_{n,k})=\exp(-\eta s_{n,k})$ represent the pairwise potential and self-potential functions, respectively, with $\varpi$ and $\eta$ being the model parameters associated with $p(\mathbf{S})$.
+where $\mathcal{D}_{n,k}$ denotes the set of neighboring nodes of $s_{n,k}$; $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$ and $v(s_{n,k})=\exp(-\eta s_{n,k})$ represent the pairwise potential and self-potential functions, respectively, with $\varpi$ and $\eta$ being the model parameters associated with $p(\mathbf{S})$.
 By leveraging the formulation in (35), the prior encourages block patterns in the support structure and suppresses isolated coefficients that deviate from their neighboring states.
@@
 Since the measurement matrix $\boldsymbol{\Psi}$ may be “bad” (e.g., rank-deficient, ill-conditioned, or having a non-zero mean) [39], the divergence issues might be arisen in the Bayesian inference.
-To address this, we first perform unitary transformations on the measurement matrix, i.e., $\boldsymbol{\Psi}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$.
-Consequently, the original signal model in (30) is rewritten as
+To address this, we first perform unitary transformations on the received signal.
+Let the singular value decomposition (SVD) of the measurement matrix be denoted as $\boldsymbol{\Psi}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$, where $\mathbf{U}$ and $\mathbf{V}$ are two unitary matrices.
+Performing a unitary transformation with $\mathbf{U}^{\mathrm{H}}$ on the received signal model in (30) yields the following model
 $\mathbf{R}=\mathbf{A}\mathbf{X}+{\mathbf{W}}=\mathbf{Z}+{\boldsymbol{\Xi}},$ (36)
@@
 Based on the prior model provided in (31), the maximum a posterior (MAP) estimator for the $(n,k)$-th entry of $\mathbf{X}$ can be expressed as
-$\hat{x}_{n,k}=\int{x}_{n,k}p(x_{n,k}|\mathbf{R})\mathrm{d}x_{n,k},$ (37)
-where the posterior distribution $p(x_{n,k}|\mathbf{R})$ is given by
-$\int p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\,\mathrm{d}\beta\,\mathrm{d}\mathbf{Z}\,\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\,\mathrm{d}\mathbf{\boldsymbol{\Omega}}\,\mathrm{d}\mathbf{S}\,\mathrm{d}\mathbf{X}_{\backslash n,k}$ (38)
-with $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ indicating the joint posterior probability and satisfying
-$p(\beta,\mathbf{X},\mathbf{Z},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\propto p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S},\mathbf{R})$ (39)
-$=$ $p(\mathbf{R}|\mathbf{Z},\beta)p(\mathbf{Z}|\mathbf{X})p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S})p(\beta),$
-where $p(\beta)\propto\beta^{-1}$ denotes the noise precise level; $p(\mathbf{R}|\mathbf{Z},\beta)$ and $p(\mathbf{Z}|\mathbf{X})$ are given by
-$p(\mathbf{R}|\mathbf{Z},\beta)$ $=\frac{1}{(\pi\beta^{-1})^{MK}}\prod_{m}\prod_{k}p(r_{m,k}\mid z_{m,k}),$ (40)
+$\hat{x}_{n,k}=\int{x}_{n,k}p(\boldsymbol{\Theta}|\mathbf{R})\mathrm{d}\beta\mathrm{d}\mathbf{Z}\mathrm{d}\mathbf{X}\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\mathrm{d}\mathbf{\boldsymbol{\Omega}}\mathrm{d}\mathbf{S},$ (37)
+where $\boldsymbol{\Theta}\triangleq\{\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}\}$, and $p(\boldsymbol{\Theta}|\mathbf{R})$ denotes the joint posterior probability, which is defined as
+$p(\boldsymbol{\Theta}|\mathbf{R})\propto p(\mathbf{R}|\mathbf{Z},\beta)p(\mathbf{Z}|\mathbf{X})p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})p(\beta),$ (38)
+where $p(\beta)\propto\beta^{-1}$ denotes the prior distribution of noise precision; the conditional distributions $p(\mathbf{R}|\mathbf{Z},\beta)$ and $p(\mathbf{Z}|\mathbf{X})$ are respectively given by
+$p(\mathbf{R}|\mathbf{Z},\beta)$ $=\prod_{m}\prod_{k}p(r_{m,k}\mid z_{m,k}),$ (39)
 $p(\mathbf{Z}|\mathbf{X})$ $=\delta(\mathbf{Z}-\mathbf{AX}),$ (40)
@@
 V-B Factor Graph Representation
-The dependencies among the random variables in the factorization (39) are illustrated using a factor graph, as shown in Fig.
+TABLE I: Factor and Distribution in (38)
+Factor
+Distribution
+Function
+$f_{\beta}$ $p(\beta)$
+$\beta^{-1}$
+$f_{r_{m,k}}$ $p(r_{m,k}|z_{m,k},\beta)$
+$\mathcal{CN}(r_{m,k};z_{m,k},\beta^{-1})$
+$f_{z_{m,k}}$ $p(z_{m,k}|\mathbf{x}_{k})$
+$\delta(z_{m,k}-\mathbf{A}_{m,:}\mathbf{x}_{k})$
+$f_{x_{n,k}}$ $p(x_{n,k}|\gamma_{n,k})$ $\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1})$
+$f_{\gamma_{n,k}}$ $p(\gamma_{n,k}|\boldsymbol{\alpha}_{k},s_{n,k})$ $\delta(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k}))$
+The dependencies among the random variables in the factorization (38) are illustrated in Fig.
 3, with the corresponding probability distributions summarized in Table I.
@@
 Figure 3: Factor graph representation for the factorization (38).
-TABLE I: Factor and Distribution in (39)
-Factor Distribution Function
-$f_{\beta}$ $p(\beta)$ $\beta^{-1}$
-$f_{r_{m,k}}$ $p(r_{m,k}|z_{m,k},\beta)$ $\mathcal{CN}(r_{m,k};z_{m,k},\beta^{-1})$
-$f_{z_{m,k}}$ $p(z_{m,k}|\mathbf{x}_{k})$ $\delta(z_{m,k}-\mathbf{A}_{m,:}\mathbf{x}_{k})$
-$f_{x_{n,k}}$ $p(x_{n,k}|\gamma_{n,k})$ $\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1})$
-$f_{\gamma_{n,k}}$ $p(\gamma_{n,k}|\boldsymbol{\alpha}_{k},s_{n,k})$ $\delta(\gamma_{n,k}-t(s_{n,k},\boldsymbol{\alpha}_{k}))$
 Moreover, in the proposed sparse prior model, the latent precision parameters $\beta$ and $\alpha^{i}_{k}$ follow non-Gaussian distributions.
-Meanwhile, the shared precision parameter simultaneously influences the message updates in both the first and third layers, resulting in inter-layer dependencies, which significantly increase the complexity of inference and render the exact computation of the node beliefs computationally intractable under a pure belief propagation (BP) framework.
-To overcome these challenges, we propose a novel MMV-HMP algorithm jointly leveraging VMP and sum-produce algorithm (SPA) rules.
-Specifically, tailored to the intractable structure of the factor graph, we reformulate the message update equations along the edges, selectively applying sum-product or variational message passing rules to different edges to enable more flexible and efficient message updates.
+Meanwhile, the shared precision parameter simultaneously influences the message updates in both the first and third layers, resulting in inter-layer dependencies, which significantly increase the complexity of inference and render the exact computation of the node beliefs computationally intractable under a pure SP framework.
+To overcome these challenges, we propose a novel MMV-HMP algorithm jointly leveraging VMP and SP rules.
+Specifically, tailored to the intractable structure of the factor graph, we reformulate the message update equations along the edges, selectively applying SP or VMP rules to different edges to enable more flexible and efficient message updates.
 The MMV-HMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (37).
-In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules as shown in Fig.
-3.
+In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules.
 V-C Measurement Module
@@
 In this manner, according to the VMP rule [44, 45], the message from $f_{r_{m,k}}$ to $\beta$ is given by
-$m_{f_{r_{m,k}}\rightarrow\beta}$ $\propto\exp\left\{\int\ln f_{r_{m,k}}b(z_{m,k})\mathrm{d}z_{m,k}\right\},$ (42)
+$m_{f_{r_{m,k}}\rightarrow\beta}(\beta)$ $\propto\exp\left\{\int\ln f_{r_{m,k}}b(z_{m,k})\mathrm{d}z_{m,k}\right\},$ (41)
 $\overset{(a)}{\propto}\beta\exp\left\{-\beta\left(\left|r_{m,k}-\hat{z}_{m,k}\right|^{2}+\hat{\nu}^{z}_{m,k}\right)\right\},$
@@
 Similar to (41), utilizing VMP rule, the forward message passing from $f_{r_{m,k}}$ to $z_{m,k}$ can be given by
-$m_{f_{r_{m,k}}\rightarrow z_{m,k}}\propto$ $\exp\left\{\int b(\beta)\ln f_{r_{m,k}}\mathrm{d}\beta\right\}$ (45)
+$m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})\propto\exp\left\{\int b(\beta)\ln f_{r_{m,k}}\mathrm{d}\beta\right\}$ (44)
 $\propto$ $\exp\left\{-\left|r_{m,k}-z_{m,k}\right|^{2}\int\beta b(\beta)\mathrm{d}\beta\right\}$
@@
 where $(a)$ is obtained by utilizing (43).
-Owing to the Gaussian form of the message in (45), we can obtain the following model $\mathbf{r}_{k}=\mathbf{z}_{k}+\boldsymbol{\xi}_{k}$, where $\boldsymbol{\xi}_{k}\sim\mathcal{CN}(\boldsymbol{\xi}_{k};0,\hat{\beta}^{-1}\mathbf{I}_{M\times M})$.
-This representation facilitates seamless integration with the forward passing of UAMP in the measurement module.
+Owing to the Gaussian form of the message in (44), we can obtain the following model $\mathbf{r}_{k}=\mathbf{z}_{k}+\boldsymbol{\xi}_{k}$, where $\boldsymbol{\xi}_{k}\sim\mathcal{CN}(\boldsymbol{\xi}_{k};0,\hat{\beta}^{-1}\mathbf{I}_{M})$, $\mathbf{r}_{k}=[r_{1,k},r_{2,k},\cdots,r_{M,k}]^{\mathrm{T}}$, and $\mathbf{z}_{k}=[z_{1,k},z_{2,k},\cdots,z_{M,k}]^{\mathrm{T}}$.
+This representation enables seamless integration with the measurement module of unitary AMP (UAMP) algorithm in [40].
 Denote the posterior mean and variance of $\mathbf{x}_{k}$ as $\hat{\mathbf{x}}_{k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$, which is defined in (63) and (64).
-According to the UAMP rules [40], define $\boldsymbol{\lambda}$ as $\left|\mathbf{A}\right|^{2}\mathbf{1}_{N}$.
+According to the UAMP algorithm [40], define $\boldsymbol{\lambda}$ as $\left|\mathbf{A}\right|^{2}\mathbf{1}_{N}$.
 Thus, the message from $\mathbf{z}_{k}$ to $f_{\mathbf{r}_{k}}$ is given by $\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})$ with
@@
 $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}=\mathbf{1}_{M}./(\mathbf{1}_{M}\hat{\beta}^{-1}+\boldsymbol{\nu}_{\mathbf{p}_{k}}),\boldsymbol{\mu}_{k}=\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}\odot(\mathbf{r}_{k}-\mathbf{p}_{k}).$ (46)
-In this case, utilizing SPA rule, the belief of $\mathbf{z}_{k}$ is given by
-$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}$ (48)
-$=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\mathcal{CN}(\mathbf{z}_{k};\mathbf{r}_{k},\hat{\beta}^{-1}\mathbf{I}_{M\times M})$
-$\overset{(a)}{\propto}\mathcal{CN}(\mathbf{z}_{k};\hat{\mathbf{z}}_{k},\boldsymbol{\nu}_{\mathbf{z}_{k}}),$
-where $(a)$ is obtained by the property of Gaussian distribution, i.e., $\mathcal{CN}(x;\mu_{1},\nu_{1})\mathcal{CN}(x;\mu_{2},\nu_{2})\propto\mathcal{CN}(x;\mu,\nu)$ with $\nu=(1/\nu_{1}+1/\nu_{2})^{-1}$ and $\mu=\nu(\mu_{1}/\nu_{1}+\mu_{2}/\nu_{2})$; In addition, $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}$ are respectively given by
+In this case, utilizing SP rule, the belief of $\mathbf{z}_{k}$ is given by
+$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})$ (47)
+$=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\mathcal{CN}(\mathbf{z}_{k};\mathbf{r}_{k},\hat{\beta}^{-1}\mathbf{I}_{M})$
+${\propto}\mathcal{CN}(\mathbf{z}_{k};\hat{\mathbf{z}}_{k},\boldsymbol{\nu}_{\mathbf{z}_{k}}),$
+where $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}$ are respectively given by
 ${\boldsymbol{\nu}}_{\mathbf{z}_{k}}$ $=\boldsymbol{\nu}_{\mathbf{p}_{k}}./(\mathbf{1}+\hat{\beta}\boldsymbol{\nu}_{\mathbf{p}_{k}}),$ (48)
@@
 Utilizing VMP rule, similar to (41), the message from $f_{x_{n,k}}$ to $\gamma_{n,k}$ is given by
-$m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}\propto$ $\exp\left\{\int{b}(x_{n,k})\ln f_{x_{n,k}}\mathrm{d}x_{n,k}\right\}$ (52)
+$m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})\propto$ $\exp\left\{\int{b}(x_{n,k})\ln f_{x_{n,k}}\mathrm{d}x_{n,k}\right\}$ (51)
 $\propto$ $\gamma_{n,k}\exp(-\gamma_{n,k}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})).$
-Furthermore, according to SPA rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
+Furthermore, according to SP rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
 $m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})\overset{(a)}{\propto}\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}b(\boldsymbol{\alpha}_{k})\mathrm{d}\gamma_{n,k}\mathrm{d}\boldsymbol{\alpha}_{k}$ (52)
 $=\pi^{\mathrm{out}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{out}}_{n,k})\delta(1+s_{n,k}),$
-where $(a)$ is obtained by utilizing the approximation $b(\boldsymbol{\alpha}_{k})=b(\alpha^{1}_{k})b(\alpha^{2}_{k})\approx\prod_{j\neq n}m_{\boldsymbol{\alpha}_{k}\rightarrow f_{\gamma_{j,k}}}$ and the expectation of Gamma distribution; $\pi^{\mathrm{out}}_{n}$ is defined as
+where $(a)$ is obtained by utilizing the approximation $b(\boldsymbol{\alpha}_{k})=b(\alpha^{1}_{k})b(\alpha^{2}_{k})\approx\prod_{j\neq n}m_{\boldsymbol{\alpha}_{k}\rightarrow f_{\gamma_{j,k}}}(\gamma_{j,k})$ and the expectation of Gamma distribution; $\pi^{\mathrm{out}}_{n}$ is defined as
 $\pi^{\mathrm{out}}_{n,k}=\frac{\hat{a}_{k}^{1}(\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))}{\hat{a}_{k}^{1}(\hat{b}_{k}^{2}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))+\hat{a}_{k}^{2}(\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})}.$
@@
 The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions.
-Take $m_{n,k}^{\mathrm{l}}$ as an example, according to SPA rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (54), as shown in the top of this page.
+Take $m_{n,k}^{\mathrm{l}}$ as an example, according to SP rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (53), as shown in the top of this page.
 $\lambda_{n,j}^{\mathrm{l}}=\frac{\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}}\mathrm{e}^{\varpi-\eta}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}\lambda^{w}_{s_{n,j}^{{\text{l}}}}+(1-\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}})\mathrm{e}^{\eta-\varpi}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}(1-\lambda^{w}_{s_{n,j}^{{\text{l}}}})}{(\mathrm{e}^{\varpi}+\mathrm{e}^{-\varpi})\left(\mathrm{e}^{-\eta}\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}}\prod_{w\in\left\{\mathrm{l,t,b}\right\}}\lambda^{w}_{s_{n,j}^{{\text{l}}}}+\mathrm{e}^{\eta}(1-\pi^{\mathrm{out}}_{s_{n,j}^{{\text{l}}}})\prod_{w\in\left\{\mathrm{l,t,b}\right\}}(1-\lambda^{w}_{s_{n,j}^{{\text{l}}}})\right)}.$ (53)
@@
 V-D2 Backward Message Passing
-With the messages of neighbors and $v(s_{n,k})$ and SPA rule, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
+With the messages of neighbors and $v(s_{n,k})$ and SP rule, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
 $m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}=\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}m_{n,k}^{\mathrm{l}}v(s_{n,k})$ (54)
@@
 $\frac{\mathrm{e}^{-\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}\lambda_{n,k}^{w}}{\mathrm{e}^{-\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}\lambda_{n,k}^{w}+\mathrm{e}^{\eta}\prod_{w\in\left\{\mathrm{l,r,t,b}\right\}}(1-\lambda_{n,k}^{w})}.$ (55)
-According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
-$m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto$ $\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}\mathrm{d}\gamma_{n,k}\mathrm{d}s_{n,k},$
+Using SP rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
+$m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto$ $\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}\mathrm{d}\gamma_{n,k}\mathrm{d}s_{n,k},$
 $\overset{(a)}{\propto}$ $\alpha_{k}^{1}\exp(-\alpha_{k}^{1}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}})),$ (56)
-where $(a)$ is obtained by utilizing the property of delta function, i.e., $\int\delta(x-t)f(x)\mathrm{d}x=f(t)$.
-Similarly, $m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{2}}\propto\alpha_{k}^{2}\exp(-\alpha_{k}^{2}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$.
+where $(a)$ is obtained by utilizing the property of delta function.
+Similarly, $m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{2}}(\alpha_{k}^{2})\propto\alpha_{k}^{2}\exp(-\alpha_{k}^{2}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$.
 As a result, the belief of $\alpha_{k}^{1}$ is given by
-${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (58)
+${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (57)
 where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$.
 Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$.
-According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
-$m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}=\pi^{\mathrm{in}}_{n,j}{b}(\alpha^{1}_{k})+(1-\pi^{\mathrm{in}}_{n,j}){b}(\alpha^{2}_{k}).$ (59)
+According to SP rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
+$m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})=\pi^{\mathrm{in}}_{n,j}{b}(\alpha^{1}_{k})+(1-\pi^{\mathrm{in}}_{n,j}){b}(\alpha^{2}_{k}).$ (58)
 Consequently, the belief of $\gamma_{n,k}$ is given by
-$b(\gamma_{n,k})\propto m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})$ (60)
+$b(\gamma_{n,k})\propto m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})$ (59)
 $\propto$ $\pi^{\mathrm{in}}_{n,j}\gamma_{n,k}^{\hat{a}_{k}^{1}}\exp(-\gamma_{n,k}(\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$
@@
 As a result, utilizing VMP rule, the message from $f_{x_{n,k}}$ to $x_{n,k}$ is denoted as
-$m_{f_{x_{n,k}}\rightarrow x_{n,k}}$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (61)
+$m_{f_{x_{n,k}}\rightarrow x_{n,k}}(x_{n,k})\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (60)
 $\propto\exp\left\{-\int\left|x_{n,k}\right|^{2}\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}\right\}$
@@
 17:end while
-18:$\hat{x}_{n,k}$.
+[…]
 The proposed MMV-HMP algorithm is summarized in Algorithm 1 and it can be terminated when it reached a maximum number of iteration or the difference between the estimates of two consecutive iterations is less than $10^{-5}$.
@@
 TABLE II: Simulation Parameters
-Notations Parameters
-Number of BS antenna $N_{\mathrm{R}}$ 256
-Number of RF chain $N_{\mathrm{RF}}$ 16
-Carrier frequency $f_{c}$ 30GHz
-Number of pilot carriers $K$ 64
-System bandwidth $f_{s}$ 1.6GHz
-Number of channel path $L$ 4
-Angle of arrival $\vartheta_{l}$ $\mathcal{U}(-\pi/2,\pi/2)$
-Distance between BS and UE or scatterers $r_{l}$ [5, 50]m
-Proportion of visible antenna elements $\rho_{l}$ (0,1]
+Notations
+Parameters
+Number of BS antenna $N_{\mathrm{R}}$
+256
+Number of RF chain $N_{\mathrm{RF}}$
+16
+Carrier frequency $f_{c}$
+30GHz
+Number of pilot carriers $K$
+64
+System bandwidth $f_{s}$
+1.6GHz
+Number of channel path $L$
+4
+Angle of arrival $\vartheta_{l}$
+$\mathcal{U}(-\pi/2,\pi/2)$
+Distance between BS and UE or scatterers $r_{l}$
+[5, 50]m
+Proportion of visible antenna elements $\rho_{l}$
+(0,1]
 In this section, we evaluate the performance of the proposed channel estimation scheme under various system setups.
@@
 The algorithm employs a two-layer Gaussian-Gamma hierarchical prior model, where the posterior estimates of $\mathbf{x}_{k}$ are updated in the E-step, and the hyperparameters of the prior model are updated in the M-step.
-  • UAMP-SBL[40]: An improved version of the StdSBL algorithm that leverages the unitary AMP framework to perform the E-step and variational message passing to achieve the updates of prior parameters.
+  • UAMP-SBL[40]: An improved version of the StdSBL algorithm that leverages the UAMP framework to perform the E-step and variational message passing to achieve the updates of prior parameters.
   • PC-SBL[47]: A variant of the StdSBL framework that incorporates a pattern-coupled Gaussian prior model to exploit the block sparsity inherent in signals.
@@
 In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity.
-In contrast, the proposed MMV-HMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both MRF structure and precision parameter sharing mechanism.
+Overall, the performance degradation of these SBL-based baselines stems from the mismatch between their prior models and the structured sparsity of XL-MIMO channels.
+In contrast, the proposed MMV-HMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both MRF structure and precision sharing mechanism.
 This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
@@
 In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios.
-Notably, the proposed MMV-HMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
+Notably, the proposed MMV-HMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic scattering environments.
 Figure 9: Ablation study of the column-wise hierarchical model.
@@
 To exploit this structured sparsity, we proposed a computationally efficient MMV-HMP algorithm.
-The algorithm performs message updates along the factor graph by selectively applying either the SPA or VMP rules, depending on the structure and tractability of each factor node.
-Simulation results demonstrate that the proposed MMV-HMP algorithm consistently outperforms existing baseline methods, achieving robust estimation performance with moderate computational complexity.
-Furthermore, an ablation study confirms the effectiveness of the proposed column-wise hierarchical prior, validating its contribution to improved estimation accuracy.
+Simulation results demonstrate the superiority of the MMV-HMP algorithm in both computational complexity and estimation performance.
+Furthermore, an ablation study confirms the effectiveness of the proposed column-wise hierarchical prior, validating its key contribution to enhanced estimation accuracy.
 Appendix A Proof of Lemma 1
@@
 Next, we show that the spatial-delay representation $\boldsymbol{\Theta}(\varphi_{,}\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ exhibits column-wise sparsity.
-Without loss of generality, we assume that $\cos\vartheta>0$.
-We begin by examining the angular transformation of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$.
 For a fixed $k$, the angular transformation of $\boldsymbol{\theta}(f_{k},\varphi_{l},\psi_{l})$ is given by
-$[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{m,k}=\sum_{n=1}^{N}\mathrm{e}^{\mathrm{j}\frac{2\pi}{N}mn}\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n}.$ (A.1)
-According to (21) and (22), it can be obtained that $[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{:,k}$ demonstrates block sparsity with its prominent spatial frequency components concentrated from $i_{l,e}(k)$ to $i_{l,s}(k)$.
+$\boldsymbol{\Xi}_{l,\mathrm{A}}]_{m,k}=\sum_{n=1}^{N}\mathrm{e}^{\mathrm{j}\frac{2\pi}{N}mn}\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n}.$ (A.1)
+According to (21) and (22), it can be obtained that $\boldsymbol{\Xi}_{l,\mathrm{A}}]_{:,k}$ demonstrates block sparsity with its prominent spatial frequency components concentrated from $i_{l,e}(k)$ to $i_{l,s}(k)$.
 Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{l,e}^{\mathrm{min}}=\min_{k}i_{l,e}(k)$ to $i_{l,s}^{\mathrm{max}}=\max_{k}i_{l,s}(k)$.
@@
 where $t_{l,n}=n^{2}\frac{f_{s}\varphi_{l}}{Kf_{c}}-n\frac{f_{s}\psi_{l}}{Kf_{c}}$, which is a quadratic function about $n$. Further, we have
-$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{l,n}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{l,n}-\frac{m}{K}\right).$ (A.4)
-From (A.4), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{l,n}K$.
-However, due to the finite sampling size $K$ in practice, we consider the range from $J_{l,e}(n)=(t_{n}-{2}/{K}+1/2)K$ to $J_{l,s}(n)=(t_{n}+{2}/{K}+1/2)K$ to collect the delay indices near the peak effectively.
+$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{l,D}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{l,n}-\frac{m}{K}\right).$ (A.4)
+From (A.4), when $K\to\infty$, $\boldsymbol{\Xi}_{l,D}]_{n,:}$ can be approximated as a delta function centered at $m=t_{l,n}K$.
+However, due to the finite sampling size $K$ in practice, we consider the range from $j_{l,s}(n)=(t_{n}-{2}/{K}+1/2)K$ to $j_{l,e}(n)=(t_{n}+{2}/{K}+1/2)K$ to collect the delay indices near the peak effectively.
 Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
-$\boldsymbol{\Xi}_{l,\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{l,e}-1)},\mathbf{L}_{N\times(J_{l,s}-J_{l,e}+1)},\mathbf{0}_{N\times(K-J_{l,s})}\right],$ (A.5)
-where $J_{l,s}=\min_{n}J_{l,s}(n)$ to $J_{l,e}=\max_{n}J_{l,e}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
+$\boldsymbol{\Xi}_{l,\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{l,s}-1)},\mathbf{L}_{N\times(J_{l,e}-J_{l,s}+1)},\mathbf{0}_{N\times(K-J_{l,e})}\right],$ (A.5)
+where $J_{l,s}=\min_{n}j_{l,s}(n)$ and $J_{l,e}=\max_{n}j_{l,e}(n)$ with $n\in\boldsymbol{\phi}_{l}$, and $\mathbf{L}$ is a non-zero matrix.
 According to row-wise and column-wise block sparsity of (A.2) and (A.5), we assert that $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ adheres to the following block-sparse structure, with its significant entries localized within the square region $\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\}$, i.e.,
@@
 where $\mathbf{V}$ denotes a non-zero matrix.
-Moreover, based on the definition $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$, it follows that, for a fixed column index $k$, $\boldsymbol{\Xi}_{l}$ characterizes the angular distribution corresponding to a specific propagation delay, which is determined by the spatial frequencies across different subcarriers.
-Accordingly, within the region $\mathcal{A}_{l}$, each column captures the angular response for a given delay, and the width of significant spatial frequency components is approximately determined by the system width and the start spatial frequency, i.e., $i_{l,k}N{f_{s}}/{f_{c}}<I_{l,s}-I_{l,e}$, where $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$ denotes the starting spatial frequency of the $k$-th column.
+Moreover, for a fixed column index $k$, $\boldsymbol{\Xi}_{l}$ characterizes the angular distribution corresponding to a specific propagation delay, which is determined by the spatial frequencies across different subcarriers.
+Accordingly, within the region $\mathcal{A}_{l}$, each column captures the angular response for a given delay, and the width of significant spatial frequency components is approximately determined by the system bandwidth and the start spatial frequency, i.e., $i_{l,k}N{f_{s}}/{f_{c}}$, where $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$ denotes the starting spatial frequency of the $k$-th column.
 As a result, the nonzero components of $\boldsymbol{\Xi}_{l}$ are confined within a narrow band in each column, leading to a distinct column-wise clustered sparsity pattern.
@@
 According to (20), we further define
-$\mathbf{H}_{l,p}=\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (B.1)
+$\mathbf{H}_{l,p}=\alpha_{l}c_{l,p}\left(\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\right)\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (B.1)
 Applying the shift property of 2D-IDFT, $\mathbf{X}_{l,p}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}_{l,p}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ should satisfy the block sparsity with its significant entries localized within the square region $\mathcal{B}_{l,p}$ with
-$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (B.2)
+$\mathcal{B}_{l,p}$ $=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right$. (B.2)
 $k$ $\left.=\mathrm{mod}(k_{l,1}+f_{s}\tau_{l},K),\forall(n_{l,1},k_{l,1})\in\mathcal{A}_{l}\right\}.$
-Here, $\mathcal{A}_{l}$ is the non-zero region for $\boldsymbol{\Xi}_{l}$ given by (25), and mod$(a,m)$ is the modulus of $a$ for $m$.
+Here, $\mathcal{A}_{l}$ is the non-zero region for $\boldsymbol{\Xi}_{l}$ given by (25), and mod(a,m)(a,m) is the modulus of $a$ for $m$.
 Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{B}_{l}\triangleq\cup_{p}\mathcal{B}_{l,p}$.